### v1

# A Nonparametric Bayesian Local-Global Model for Enhanced Adverse Event Signal Detection in Spontaneous Reporting System Data

Abstract
Spontaneous reporting system databases are key resources for post-marketing surveillance, providing real-world evidence (RWE) on the adverse events (AEs) of regulated drugs or other medical products. Various statistical methods have been proposed for AE signal detection in these databases, flagging drug-specific AEs with disproportionately high observed counts compared to expected counts under independence. However, signal detection remains challenging for rare AEs or newer drugs, which receive small observed and expected counts and thus suffer from reduced statistical power. Principled information sharing on signal strengths across drugs/AEs is crucial in such cases to enhance signal detection. However, existing methods typically ignore complex between-drug associations on AE signal strengths, limiting their ability to detect signals. We propose novel local-global mixture Dirichlet process (DP) prior-based nonparametric Bayesian models to capture these associations, enabling principled information sharing between drugs while balancing flexibility and shrinkage for each drug, thereby enhancing statistical power. We develop efficient Markov chain Monte Carlo algorithms for implementation and employ a false discovery rate (FDR)-controlled, false negative rate (FNR)-optimized hypothesis testing framework for AE signal detection. Extensive simulations demonstrate our methods’ superior sensitivity—often surpassing existing approaches by a twofold or greater margin—while strictly controlling the FDR. An application to FDA FAERS data on statin drugs further highlights our methods’ effectiveness in real-world AE signal detection. Software implementing our methods is provided as supplementary material.
1 Introduction
Adverse events (AEs) related to drugs and other medical products have a profound impact on public health and have received heightened attention over the past decades. While clinical trials remain the gold standard for assessing medical product safety before approval, they have limitations due to factors such as small sample sizes, ethical considerations, restricted timelines, etc., and the critical need for postmarket monitoring is increasingly recognized. For example, severe AEs such as congenital malformations from Thalidomide and liver injury from the Ketoconazole (Lavertu et al.,, 2021) were identified only after these drugs were approved for the U.S. market. To complement clinical trial findings with real-world evidence (RWE), continuous post-marketing surveillance—a cornerstone of contemporary pharmacovigilance—has proven essential. Accordingly, several Spontaneous Reporting System (SRS) databases have been established globally, including VigiBase (World Health Organization), and FAERS and VAERS (U.S. Food and Drug Administration). These databases collect and manage AE data for regulated medical products, such as drugs, medical devices, vaccines, and therapeutic biologics. For exposition, we focus only on drugs hereinafter, though analogous considerations apply to other medical products monitored in SRS databases.
The SRS databases serve as key resources for assessing real-world evidence in postmarket drug safety by cataloging a large number of case safety reports linking drugs to relevant AEs. For example, the FAERS database has compiled over $29$ million reports by 2024. These databases are inherently observational and suffer from selection bias, measurement error, confounding, and missing data, among other factors (Markatou and Ball,, 2014), precluding causal inference. Instead, the focus is on detecting signals: identifying drug-AE pairs that appear disproportionately more often than their null expected counts, derived under the assumption of independence between the drug and the AE using their marginal totals and the dataset’s grand total counts. Association analyses then estimate and assess the significance of signal strengths, quantifying the differences between the observed and (null) expected counts or related quantities defined by parameters of some underlying statistical model.
Numerous SRS data mining and signal detection approaches have been developed over the past decades. Examples include the proportional reporting ratio (PRR) (Evans et al.,, 2001), the reporting odds ratio (ROR) (Rothman et al.,, 2004), formal frequentist methods such as the likelihood ratio test (LRT) and its variants (Huang et al., 2011; Chakraborty et al., 2022; Ding et al., 2020; and the references cited therein), and Bayesian methods such as the Gamma Poisson Shrinker (GPS) (DuMouchel,, 1999), Bayesian confidence propagation neural network (BCPNN) (Bate et al.,, 1998), and Dirichlet process (DP) Poisson Gamma mixture model (Hu et al.,, 2015). Some approaches—e.g., PRR and ROR—employ ad-hoc thresholds on empirical signal strength estimates, while the more formal methods—e.g., LRT and Bayesian approaches—employ rigorous probabilistic tools for statistically guaranteed signal detection, including controlled type I errors and false discovery rates (FDR). Despite these advances, signal detection for rare AEs and newer drugs remains challenging due to their small (marginal total) sample sizes, which leads to small observed and expected report counts in individual drugs, reducing the statistical power to detect the underlying signals. Yet, a large and growing body of research has documented the critical need for reliable pharmacovigilance signal detection precisely under these setups (see, e.g., Poluzzi et al., 2012; Sardella and Belcher, 2018; Raschi et al., 2021, and the references cited therein).
Bayesian methods facilitate principled inference and signal detection for SRS data by modeling drug-AE signal strengths as model parameters with suitable priors and utilizing their posterior distributions. However, prior selection is crucial. An overly restrictive prior—e.g., a single gamma shrinkage prior for all drug-AE signal strength parameters—may cause excessive shrinkage, pulling estimates toward a single common value (Hu et al.,, 2015). Conversely, an overly flexible prior—e.g., independent flat priors for all parameters—may inadequately regularize their estimation, resulting in overfitting and noisy estimates. State-of-the-art Bayesian data mining approaches balance flexibility and regularization using semi/nonparametric shrinkage priors—e.g., a two-component gamma mixture prior (the GPS model of DuMouchel, (1999)) or a DP mixture prior with a base gamma distribution (Hu et al.,, 2015)—for AE signal strengths separately for each drug. This stabilizes signal detection, provided the drugs are treated either all independently or as a single group.
However, real-world SRS databases record drugs with nuanced similarities in their AEs. For instance, various statin drugs, used to lower blood cholesterol and cardiovascular risk, share chemical compositions and function, leading to common AEs like rhabdomyolysis (Montastruc,, 2023). On the other hand, small differences in drug composition and function can cause distinct AEs or variations in AE severity; e.g., myoglobinuria is typically associated with Atorvastatin but not Lovastatin (Seneff,, 2016). Figure 1 shows the pairwise Spearman correlation coefficients for occurrences of 46 common AEs of six statin drugs as cataloged in FAERS (Q3 2014-Q4 2024). These correlations reveal strong yet nuanced between-drug associations: they are consistently above 0.70, ruling out independence, yet many are noticeably below one, suggesting against a single grouping of all drugs.
Figure 1: Spearman correlation matrix for occurrence counts of 46 common AEs (PTs) among statin drugs in the FDA FAERS data (2014 Q1-2020 Q4)
Existing SRS data mining approaches, both Bayesian and frequentist, rarely account for nuanced between-drug associations. Instead, they often assume either independence of drugs (the original LRT, BCPNN, GPS, and DP Poisson-Gamma) or a single group for all drugs (extended LRT of Huang et al., (2011)). These restrictive assumptions limit the statistical power and sensitivity for detecting signals, particularly for rare AEs and newer drugs with small observed and expected counts. Indeed, our extensive simulations (Section 4) show that existing methods suffer from poor sensitivity (after controlling FDR) when detecting signals from datasets containing multiple but weak true signals, rare AEs, and newer drugs.
We address this critical gap with a novel local-global mixture DP-based Bayesian method that balances flexibility and shrinkage in AE signal strengths for each drug while enabling data-adaptive information sharing across drugs based on their associations, enhancing statistical power for signal discovery. In this framework, a local DP models distinct AE signal strength patterns specific to each drug while a global process defines a common pattern across drugs; the final AE signal strengths for a drug are obtained from a probabilistic mixture of these local and global processes. By appropriately controlling the local-global contributions, our approach can flexibly capture complex between-drug associations in SRS data.
For full Bayesian inference from our model, we employ weakly informative priors customized for SRS data mining on various DP hyperparameters. This eliminates the need for any additional tuning and acknowledges all modeling uncertainty, including hyperparameters. Subsequently, we develop efficient Markov chain Monte Carlo (MCMC) algorithms for posterior computation. Signal detection is performed using a rigorous hypothesis testing framework that controls the false discovery rate (FDR) while minimizing the false negative rate (FNR), both estimated via MCMC-computed posterior probabilities. Our default implementation supports a wide range of data-generating settings for robust signal detection. Additional inferential goals, such as identifying structural zero (physically or biologically impossible drug-AE pairs), can be seamlessly incorporated through additional model layers without compromising signal detection accuracy, which we also discuss.
Our key contributions are as follows. First, we propose local-global mixture DP models for SRS signal detection (Section 3). Local-global priors are widely used in contemporary statistics (Bhadra et al., 2019 and references cited therein) and have been employed in DP modeling of grouped data, notably via hierarchical DP models (Teh et al.,, 2004). Our framework instead adopts a more transparent two-component mixture of DPs (Antoniak,, 1974) for local-global pattern decomposition. To our knowledge, the use of local-global DP models for principled SRS data mining is novel. We develop efficient MCMC algorithms for model implementation (Section 3.3 and Supplement S.3; software in the Supplement) and leverage FDR-controlled, FNR-optimized hypothesis tests for signal detection (Section 3.4). Second, we provide an intuitive overview of DP-based SRS data mining models to aid broader understanding (Section 3.1). Third, through extensive simulations (Section 4), we show that our methods achieve comparable or superior sensitivity to existing approaches (reviewed in Supplement S.1) when only a few strong signals are present across common AEs but yield substantial sensitivity gains—often twofold or more—when many weak signals exist among rare AEs and newer drugs, all while maintaining strong FDR and type I error control. Finally, our analysis of a real-world statin dataset (Section 5) from the FDA FAERS database demonstrates our methods’ ability to effectively detect AE signals in practice.
2 SRS DATA, OUR NOTATION, AND PROBABILITY MODELS
An SRS database stores a large number of individual case safety reports that can be summarized as a high-dimensional contingency table cataloging the occurrences of multiple AEs across the rows and various drugs across the columns. Let $I$ and $J$ denote the total number of AEs and drugs under consideration, respectively. Further, let $n_{ij}$ be the total number/count of reported cases for drug-$j$ and AE-$i$; $i=1,\dots,I$ and $j=1,\dots,J$ constituting the contingency table $((n_{ij}))$. Define $n_{i\bullet}=\sum_{j=1}^{J}n_{ij}$, $n_{\bullet j}=\sum_{i=1}^{I}n_{ij}$, and $n_{\bullet\bullet}=\sum_{i=1}^{I}\sum_{j=1}^{J}n_{ij}$.
Signal detection in SRS contingency tables often relies on the observed relative reporting rates, $n_{ij}/E_{ij}$, or a related quantity, as a measure of signal strength. Here, $E_{ij}=n_{i\bullet}n_{\bullet j}/n_{\bullet\bullet}$ represents the null expected count of drug-$j$ and AE-$i$, assuming independence between the drug and AE. Intuitively, if $n_{ij}/E_{ij}\gg 1$, the observed frequency for $(i,j)$ is much higher than the null expected count, suggesting a potential signal. Challenges arise for small $n_{\bullet j}$ (e.g., a new drug) and/or small $n_{i\bullet}$ (rare AE), leading to a small $E_{ij}$ and rendering $n_{ij}/E_{ij}$ to be noisy. Bayesian approaches parameterizing $n_{ij}/E_{ij}$ (or an analogous quantity) using a probability model for $n_{ij}$ can stabilize inference on signal strengths under these setups by incorporating suitable prior distributions that can facilitate principled information sharing across AEs and drugs, leading to improved signal detection.
The Poisson model $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$ is commonly used for its analytical simplicity and computational tractability. Here the signal strength parameter $\lambda_{ij}$ parametrizes $n_{ij}/E_{ij}$, with the pair $(i,j)$ considered a signal if $\lambda_{ij}\gg 1$ and non-signal otherwise. When the Poisson assumption is violated, particularly due to an excess of zero counts arising from structural zeros—i.e., drug-AE pairs that are physically or biologically impossible to co-occur—a zero-inflated Poisson (ZIP) model is often preferred. In this framework, $n_{ij}$ is modeled as a mixture of two components: a $\text{Poisson}(\lambda_{ij}E_{ij})$ component for drug-AE pairs that can co-occur and a point mass at 00, denoted by $\delta_{\{0\}}$ (the Kronecker delta), for structural zeros. Importantly, the structural zero positions are typically unknown and inferred from the SRS dataset.
A Bayesian SRS data mining approach begins with such a probability model for $n_{ij}$ (e.g., Poisson or ZIP), yielding a likelihood for the signal strengths $\lambda_{ij}$ that is combined with a prior distribution to derive the posterior distribution of $\lambda_{ij}$ for inference. The marginal distribution of $n_{ij}$, after integrating over the prior, can be more flexible than the likelihood model for $n_{ij}$. Indeed, with an appropriate nonparametric mixture prior, a flexible marginal distribution accommodating ZIP-like behavior can be achieved using only a Poisson likelihood for $n_{ij}$ (see Remark Remark under Section 3). Thus, Bayesian SRS data mining methods are less sensitive to the likelihood choice (Poisson or ZIP) when a flexible prior is used. However, if the zero-inflation probability parameters are of interest, the ZIP model remains useful.
3 NOVEL LOCAL-GLOBAL DP MODELS FOR SRS DATA MINING
We begin in Section 3.1 with a review of the drug-specific DP Poisson model of Hu et al., (2015), outlining its mechanism and providing guidance on the hyperparameters. We then introduce our novel local-global DP framework (Section 3.2) that accommodates between-drug associations and enables effective information sharing across drugs.
3.1 A conceptual overview on the DP Poisson model
Consider the Poisson likelihood $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$, with the DP prior (Hu et al.,, 2015):
$\lambda_{ij}\sim D_{j}\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}};\ D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$ (1)
for the signal strength parameters $\{\lambda_{1j}\dots,\lambda_{Ij}\}$ associated with each drug $j$. Here $\delta_{\{\theta\}}$ denotes a degenerate distribution at $\theta$ and DP$(\alpha,G(\beta))$ is a Dirichlet process with concentration hyperparameter $\alpha>0$ and base distribution $G(\beta)\equiv$ Gamma(shape = $\beta$, rate = $\beta$) with hyperparameter $\beta>0$. The discrete mixture structure ensures that $D_{j}$, the prior for $\lambda_{ij}$, is almost surely discrete/categorical placing probability weights $\eta_{hj}$; $10\leq\eta_{hj}\leq 1;\sum_{h=1}^{\infty}\eta_{hj}=$ 1 on atoms $\theta_{hj}$. This permits clustering, allowing multiple $\lambda_{ij}$ to share an atom $\theta_{hj}$ as a common value. The probability weights $\eta_{hj}$ are generated using stick-breaking (Sethuraman,, 1994):
$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj})\text{ for }h\geq 2;\ v_{hj}\sim\text{independent Beta}(1,\alpha_{j})\text{ for }h\geq 1$ (2)
based on the DP concentration hyperparameter $\alpha_{j}>0$. Meanwhile, the atoms are generated independently from the DP base distribution: $\theta_{hj}\sim G(\beta_{j})$. In practice, the infinite mixture defining $D_{j}$ can be well-approximated by a finite mixture $\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$ with a sufficiently large $K$ (Ishwaran and Zarepour,, 2002). This approximation simplifies model fitting and offers intuitive insights on the hyperparameters $\alpha_{j}$ and $\beta_{j}$, guiding their specification.
Role of $\alpha_{j}$. The stick-breaking representation (2) elucidates that when $\alpha_{j}<1$, each $v_{hj}$ is, on average, greater than $1/2$ with smaller $\alpha_{j}$ favoring $v_{hj}$ closer to one. Consequently, $\eta_{hj}$ decreases sharply in $h$, leading the mixture (1), or its finite approximation with $K$ components, to retain only a few non-empty components with sufficiently large weights $\eta_{hj}$. Under a finite mixture approximation with $K_{+}<K$ non-empty components, (1) models $\lambda_{ij}$ using a categorical distribution with $K_{+}$ categories, placing point masses at the atoms $\theta_{1j},\dots,\theta_{K_{+}j}$ with weights $\eta_{1j},\dots,\eta_{K_{+}j}$. This fosters clustering, restricting $\lambda_{1j},\dots,\lambda_{Ij}$ to at most $K_{+}$ distinct values. When $K_{+}<I$, multiple $\lambda_{ij}$ share the same cluster $h$ admitting the common value $\theta_{hj}$, facilitating their pooling for information sharing. While a smaller $\alpha$ encourages a smaller $K_{+}$ aiding greater information sharing, a larger $K_{+}$ ($>2$) is permitted in the posterior when supported by data. This contrasts with the GPS framework of DuMouchel, (1999) commonly used for SRS signal detection (see Section S.1 for a review) which strictly categorizes $\lambda_{ij}$ into two groups, intuitively, signal ($\lambda_{ij}\leq 1$) and non-signal ($\lambda_{ij}\gg 1$). The DP framework, by allowing $K_{+}>2$, enhances heterogeneity acknowledgment within both signal and non-signal groups. Figure 2a illustrates this with $I=6$ AEs and $K_{+}=3$ clusters.
Because a smaller $\alpha_{j}$ promotes greater information sharing, a prior assigning sufficient mass to small positive values is recommended for stable estimation of $\lambda_{ij}$. Hu et al., (2015) used a Uniform(0.2, 10) prior for $\alpha_{j}$. We suggest a more aggressive weakly informative default prior, viz., Exponential(rate = $\psi_{\alpha}$) with a moderately large $\psi_{\alpha}$, e.g., $\psi_{\alpha}=3$ (placing $\approx 95\%$ mass in $(0,1)$), which favors small $\alpha_{j}$ while allowing larger values if needed. This choice also ensures conjugacy for efficient Gibbs sampling updates for MCMC implementation.
Figure 2: Schematic diagrams visualizing the mechanisms of a “local only” DP prior with one drug (Panel A) and local-global mixture DP prior with $J=3$ drugs (panel B). Throughout, darker shades represent larger values. Panel A: The observed relative reporting rates $n_{ij}/E_{ij}$ for $I=6$ AEs for the drug $j$ (Panel A.1) show high variability across $i$ (different shades of green) due to randomness in observed data. The atoms $\theta_{hj}$ (Panel A.2) of the DP prior enable clustering of $I=6$ AE signal strengths into $K_{+}=3$ non-empty DP clusters. This ensures information sharing: the AEs $(i=1,i=3,i=5)$ in cluster 1 and separately the AEs $(i=2,i=6)$ in cluster 2 (cluster memberships are visualized via arrows between Panels A.1 and A.2) share their observed $(n_{ij},E_{ij})$ information to inform their cluster-specific DP atoms (single atom per cluster). These DP atoms, together with the cluster memberships, produce the final $I=6$ AE signal strengths $\lambda_{ij}$ (Panel A.3). Panel B: The noisy observed data $(n_{ij},E_{ij})$ leading to noisy observed relative reporting rates $n_{ij}/E_{ij}$ (Panel B.1; different shades of green) are used to produce “local” signal strengths $\lambda_{ij}^{l}$ for $I=6$ AEs and $J=3$ drugs using a separate, independent local DP prior for each drug (Panel B.2.1), and global AE signal strengths $\lambda_{i}^{g}$ for shared by all $J=3$ drugs based on a common global DP prior (Column B.2.2). Probabilistic two-component mixtures of $\lambda_{ij}^{l}$ and $\lambda_{i}$ (Panel B.3) based on the local distribution indicators $z_{ij}^{l}$ produce the final realized signal strengths $\lambda_{ij}$ (Panel B.4).
Role of $\beta_{j}$. The parameter $\beta_{j}$ controls variability among the atoms $\theta_{hj}$ drawn from the base $G(\beta_{j})\equiv$ Gamma(shape = $\beta_{j}$, rate = $\beta_{j}$) distribution, thereby controlling the shrinkage in $\lambda_{ij}$. Since $G(\beta_{j})$ has unit mean and variance $1/\beta_{j}$, larger (smaller) values of $\beta_{j}$ encourage all $\theta_{hj}$ to concentrate more (less) around $1$, inducing more (less) shrinkage in $\lambda_{ij}$ (around $1$).
[…]
Remark.
[…]
3.2 The local-global DP Poisson model
The DP model in Section 3.1 enables flexible and effective signal detection for each drug separately but does not account for associations between drugs in their signal strenghts. To capture these associations, we introduce the concept of local and global (prior) distributions for AE signal strengths. Specifically, we assume that the realized AE signal strengths $\{\lambda_{1j},\dots,\lambda_{Ij}\}$ for each drug $j$ is drawn from a mixture of a “local” distribution specific to the drug and a “global” distribution shared among drugs with similar compositions or functions.
Under the same likelihood $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$, we consider the following local-global mixture of DP prior for $\lambda_{ij}$:
$\lambda_{ij}=z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g};\ \lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D;\ z_{ij}^{l}\sim\text{Bernoulli}(\pi^{l})$
$D_{j}\sim DP(\alpha_{j},G(\beta_{j}));\ D\sim DP(\alpha,G(\beta)).$ (3)
Here, the realized signal strengths $\{\lambda_{1j},\dots,\lambda_{Ij}\}$ for a drug $j$ is generated by combining local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$ drawn from a drug-$j$-specific local distribution $D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$ as in (1) and global strengths $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ generated from a global distribution $D\sim\text{DP}(\alpha,G(\beta))$ shared by all drugs. The binary local distribution indicators $z_{ij}^{l}$ governed by the common local proportion $\pi^{l}$, determine the probabilistic mixing of $\lambda_{ij}^{l}$ and $\lambda_{i}^{g}$.
We note that the proposed local-global mixture DP prior (3) enables information sharing in signal strengths across both AEs and drugs. Specifically, for each drug $j$, the realized signal strengths $\{\lambda_{i^{\prime}j}:\text{all }i^{\prime}\text{ with }z_{i^{\prime}j}=1\}$ exhibit the same $D_{j}$-induced clustering (see Section 3.1) as the local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$, facilitating information sharing across AEs. Meanwhile, for any AE $i$, the signal strengths $\{\lambda_{ij^{\prime}}:\text{all }j^{\prime}\text{ with }z_{ij^{\prime}}=0\}$ form a cluster attaining the common global value $\lambda_{i}^{g}$, and enable information sharing across drugs (the global values $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ themselves admitting clustering via $D$). Figure 2B illustrates this mechanism using $I=6$ AEs and $J=3$ drugs. This bi-level clustering contrasts with the single-level (AE-only) clustering induced by the DP prior of Hu et al., (2015). As demonstrated in extensive simulations (Section 4), the bi-level clustering and information sharing play a critical role in stabilizing the posterior and enhancing signal detection for $\lambda_{ij}$, particularly for drug-AE pairs where $n_{i\bullet}$ is small (rare AEs) and/or $n_{\bullet j}$ is small (newer drugs), leading to small $E_{ij}$.
The drugs under consideration should be similar in composition or function to justify a single global distribution $D$. If distinct drug groups exist with no similarities, separate global processes can be used to enable separate within-group information sharing. Additionally, SRS data often include a collapsed, heterogeneous comparator category $J$ (e.g., “Other drugs”), serving as a baseline group with few, if any, distinct AE signals. In such cases, shrinking all $\lambda_{iJ}$ toward $1$ using a prior such as $\text{Gamma}(\text{shape}=\tau,\text{rate}=\tau)$ with an appropriate hyperparameter $\tau$ can improve signal detection and inference (see Supplement S.2.1).
As noted in Remark Remark, a separate ZIP likelihood $n_{ij}\sim\text{ZIP}(\lambda_{ij}E_{ij},\omega_{j})$ is unnecessary for signal detection due to the flexibility of the DP prior. However, it may be used if direct inference on structural zero positions/probabilities is required. Complete hierarchical Bayesian representations of our Poisson and ZIP-based models are provided in Supplement S.2.1.
Remark.
The Hu et al., (2015) model can be viewed as a “local-only” special case of our local-global model, with only drug-specific local distributions $D_{j}$ for $\lambda_{ij}$ and no global distribution (i.e., $\pi^{l}=1$). At the other extreme, a “global-only” model ($\pi^{l}=0$) yields common $\lambda_{i}$ across drugs. This is a restrictive model that cannot differentiate between drugs and only permits “global” signal detection—assessing whether an AE-$i$ is a signal for any drug under consideration (see Supplement S.2.2 for further discussion on this model).
Remark.
Extensive literature exists on local-global DP models for sharing information across groups while preserving DP-based clustering within each group. A common approach employs hierarchical DPs (Teh et al.,, 2004), where local DPs share a global DP as a base distribution. In contrast, our approach leverages a two-component DP mixture framework (Antoniak,, 1974), separately modeling local and global AE signal strengths and thus offering a more transparent bi-level clustering over AEs and drugs, enabling more flexible estimation of the local and global parameters for greater flexibility in SRS applications.
Remark.
In this paper, we focus on drug-based pharmacovigilance, identifying AE (row) signals given one or more drugs (columns in SRS contingency tables). An alternative AE-based approach discovers drugs strongly associated with a given set of AEs (Ding et al.,, 2020; Liu et al.,, 2024). A row-based version of our column-based prior (3), obtained by switching the roles of $i$ and $j$, can be used for such analyses.
3.3 Full Bayesian implementation using MCMC
For full Bayesian inference on our models, we place independent weakly informative proper priors on all model hyperparameters. Along with the exponential prior for $\alpha_{j}$ and $\alpha$, and half-Cauchy prior for $\beta_{j}$ and $\beta$ as suggested in Section 3.1, we consider a similar Cauchy+(0, $\psi_{\tau}$) prior for $\tau$ with $\psi_{\tau}=1/2$, and independent Uniform$(0,1)$ priors for the probability weight parameters, viz., $\pi_{l}$ and $\omega_{j}$ (if included in the model).
We develop efficient MCMC algorithms for rigorous posterior sampling, approximating the DPs using finite mixtures with $K=O(\log(IJ))$ components. Each MCMC iteration involves data augmentation to generate latent DP mixture allocations, followed by Gibbs sampling updates for the stick-breaking weights and atoms and the concentration hyperparameters for each local and global DP. The local distribution indicators $z_{ij}^{l}$, local distribution proportion $\pi^{l}$, AE signal strengths $\lambda_{iJ}$ for the reference drug $J$, and zero-inflation indicators $y_{ij}$ and probabilities $\omega_{j}$ (if included in the model) also admit standard Gibbs updates. However, the full conditionals for the DP base Gamma hyperparameters ($\beta_{j}$ and $\beta$) and the reference category shrinkage hyperparameter $\tau$ lack standard forms; for these we employ stepping-out slice sampling (Neal,, 2003). Detailed MCMC steps are provided in Supplementary Algorithms S.1, S.2 and S.3. For signal detection, we use an FDR-controlled, FNR-optimized hypothesis testing framework based on posterior draws for $\lambda_{ij}$, described next.
3.4 Signal detection using hypothesis testing
For signal detection we consider testing the hypotheses:
$H_{0,ij}:\lambda_{ij}\leq 1\text{ vs. }H_{1,ij}:\lambda_{ij}>1$
for all AEs $i=1,\dots,I$ and all drugs $j=1,\dots,J$. The null hypothesis $H_{0,ij}$ is rejected in favor of $H_{1,ij}$, identifying the drug-$j$ and AE-$i$ pair to be a signal, if $T_{ij}>k$, where $T_{ij}$ is a test statistic derived from the posterior distribution and $k>1$ is an appropriate threshold. A common choice for $T_{ij}$ is the marginal posterior posterior quantile $T_{ij}(p)$, defined by $\Pr{\!}_{\text{posterior}}(\lambda_{ij}\leq T_{ij}(p))=p$ for a small pre-specified quantile level $p\in(0,1)$, where $\Pr{\!}_{\text{posterior}}(\cdot)$ denotes posterior probability given the data, and can be computed via MCMC draws. Hu et al., (2015) suggest using the 5-th posterior quantiles $T_{ij}(0.05)$ as the test statistic with a rejection threshold of $k=2$; however, such deterministic thresholds can be overly conservative (curtailing sensitivity) or overly liberal (failing to control false discoveries). Instead, we consider a family of tests $\{T_{ij}(p)>k\}$ with varying $p$ (reasonably small) and $k>1$, selecting the optimal $(\hat{p},\hat{k})$ that controls the false discovery rate (FDR) while minimizing the false negative rate (FNR). Given $p$ and $k$, FDR and FNR are estimated using the posterior distribution as (Ahmed et al., 2010; Müller et al., 2006; in the expressions below $1(A)$ denotes the indicator of a set $A$, $\hat{=}$ denotes “estimated by”, and $T^{\text{obs}}$ is the computed value of $T$):
$\text{FDR}(p,k)\ \hat{=}\ \frac{\sum_{i,j}q_{ij}\ I(T_{ij}(p)^{\text{obs}}>k)}{\sum_{ij}1(T_{ij}(p)^{\text{obs}}>k)},\quad\text{FNR}(p,k)\ \hat{=}\ \frac{\sum_{i,j}(1-q_{ij})\ 1(T_{ij}(p)^{\text{obs}}\leq k)}{\sum_{ij}1(T_{ij}(p)^{\text{obs}}\leq k)},$ (4)
where $q_{ij}=\Pr{\!}_{\text{posterior}}(\lambda_{ij}\leq 1)$ is the posterior probability of $H_{0,ij}$, computable from MCMC draws. A grid search is used to determine the optimal $(\hat{p},\hat{k})$ that ensures $\text{FDR}(\hat{p},\hat{k})\leq\alpha$ for a pre-specified level $\alpha$ while minimizing FNR (see Supplement S.3.3 for details on the grids).
In practice, drug-AE pairs with $n_{ij}=1$ are often deemed non-signals, even if $\lambda_{ij}\gg 1$ with high posterior probability, leading to $T_{ij}(p)>k$ for suitably chosen $p$ and $k$. To prevent their detection as signals, the rejection region can be modified to $\{T_{ij}(p)>k\text{ and }n_{ij}>1\}$.
4 SIMULATION ASSESSMENTS
In this section, we conduct extensive simulations to assess the signal detection performance of our local-global DP Poisson and ZIP models under a wide array of data-generating settings. To generate random SRS contingency tables of dimension $I\times J$, we first define a prespecified true signal strength matrix (described below) $\Lambda^{0}=((\lambda_{ij}^{0}))$ encoding signals, structural zeros, non-signals, and between-drug dependency patterns. Using a reference $I\times J$ contingency table (described next) from a real-world SRS database, we extract row and column totals $\{\bm{n}_{i\bullet}:i=1,\dots,I\}$ and $\{\bm{n}_{\bullet j}:j=1,\dots,J\}$, respectively, to facilitate random contingency table generation. Next, we generate $\bm{p}_{i\bullet}\sim\text{Dirichlet}(\bm{n}_{i\bullet})$ and $\bm{p}_{\bullet j}\sim\text{Dirichlet}(\bm{n}_{\bullet j})$ for all $i$, $j$ defining the random probability matrix $\bm{p}=((p_{ij}))$ with $p_{ij}\propto\lambda_{ij}^{0}\bm{p}_{i\bullet}\bm{p}_{\bullet j}$ and $\sum_{i,j}p_{ij}=1$. Finally, we generate a random SRS contingency table by sampling from $\text{Multinomial}(\bm{n}_{\bullet\bullet};\bm{p})$, collapsing any generated $n_{ij}=1$ corresponding to a true signal (as defined by $\lambda^{0}_{ij}$) to $n_{ij}=0$ afterward to eliminate signal detection at $n_{ij}=1$.
Reference SRS dataset to compute row and column totals. We use the FDA FAERS database (Q3 2014-Q4 2020), which catalogs a total of 63,976,610 reported cases (Chakraborty et al.,, 2023), to construct our reference SRS contingency table. We focus on six statin medications: Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, collapsing all other drugs into a comparator “Other drugs” category. Using 46 preferred terms (PTs) for AEs, we form a $47\times 7$ contingency table summarizing the report counts for each drug-AE pair. This dataset has been previously analyzed in Chakraborty et al., (2022).
Constructing the true signal strength matrix $\Lambda^{0}$. To generate $\Lambda^{0}$, we define two signal patterns: (1) fixed row signals, where all drugs (columns) exhibit signals for the same AEs (PTs; rows), and (2) random signals, where signals appear on randomly chosen AEs (different row for each drug). A larger number of fixed row signals increase the between-drug associations, measured via correlations between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ for different statin drugs $j$ and $j^{\prime}$, while more random signals weaken them. The matrix $\Lambda^{0}$ is constructed by specifying fixed row signals and embedding additional random signals in the remaining rows across all statin columns. Signal strengths in these cells range from $1<\lambda_{ij}^{0}\leq 3$ (see Table 1). The remaining cells $(i^{\prime},j^{\prime})$ are designated as non-signals with $\lambda_{i^{\prime}j^{\prime}}^{0}=1$. To introduce zero inflation, we prespecify true $\omega_{j}^{0}$ (00 or $0.10$; Table 1), and set $\omega_{j}^{0}I$ randomly selected non-signal cells to $\lambda_{ij}^{0}=0$. This approach yields a controlled $\Lambda^{0}$ to generate random SRS contingency tables. Within this setup, we designed three simulation scenarios, each with and without zero-inflation (see Table 1).
Table 1: Simulation scenarios describing the construction of the true signal strength matrix $\Lambda^{0}$ encoding signals, non-signals, zero-inflation, and between-drug associations. The first two rows in the table show the number of fixed-AE/row signals and the number of random signals per (statin) column in $\Lambda^{0}$. The zero inflation rate (third row in the table) shows the drug-specific true zero-inflation probability $\omega_{j}^{0}$ in each simulation scenario. The comparator “Other drugs” column has all non-signal $(\lambda_{ij}^{0}=1)$ cells. Finally, the fourth row in the table shows the average Pearson’s correlation coefficient $\rho$ between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ over all statin drug pairs $j$ and $j^{\prime}$ (excluding the “Other drugs” category) as a measure of pairwise dependence between drugs.
Simulation Scenario 1a 1b 2a 2b 3a 3b
No. of fixed row signals 1 1 3 3 6 6
No. of random signals per column 5 5 3 3 1 1
Zero-inflation rate per column 0 0.1 0 0.1 0 0.1
Average $\rho$ between columns 0.04 0.04 0.43 0.43 0.83 0.83
  • The common/fixed PTs (fixed AEs/rows) that are considered signal in each case are as follows:
    • Case 1: Rhabdomyolysis
    • Case 2: Muscle Disorder, Myalgia, Rhabdomyolysis
    • Case 3: Blood Creatine Phosphokinase Increased, Muscle Disorder, Muscle Rupture, Musculoskeletal Discomfort, Myalgia, Rhabdomyolysis
  • A common signal strength value $\lambda^{0,\text{signal}}$ is assigned to all signal cells in a given simulation setting, and we vary $\lambda^{0,\text{signal}}\in\{1.1,1.2,1.3,1.4,1.5,1.7,1.9,2,2.5,3\}$.
Replicated Data Generation, Model Fitting, and signal detection. For each constructed signal strength matrix $\Lambda$ (Table 1), and using the random contingency table generation mechanism described earlier, we generate $R=1{,}000$ independent random contingency table replicates. In each replicate, we fit our models via MCMC sampling (Supplementary Algorithms S.1, S.2 and S.3) with $10{,}000$ post-burn-in draws after discarding the first $5,000$ burn-in draws; see Supplement S.3 for additional computational details, including starting points. From the computed posteriors, we conduct signal detection using hypothesis testing (Section 3.4).
We compare our signal detection performance against existing Bayesian SRS data mining methods, viz., BCPNN, GPS, and the local-only DP model of Hu et al., (2015) (henceforth DP Hu et al). For our models and DP Hu et al., signal detection uses an optimized posterior quantile level $\hat{p}$ (fixed at 0.05 for DP Hu et al.) and threshold $\hat{k}$ ensuring FDR $\leq 0.05$ while minimizing FNR (Section 3.4), with the additional constraint $n_{ij}>1$ for our methods. To ensure fair FDR control ($\leq 0.05$) in BCPNN and GPS, we leverage the independence of signal strengths in the (empirical) Bayes posteriors for these two models and compute Benjamini-Hochberg-adjusted (Müller et al.,, 2006) posterior probabilities of being a non-signal; a drug-$j$ and AE-$i$ pair is then classified as a signal if the adjusted probability is $\leq 0.05$. Additionally, we evaluate the state-of-the-art frequentist ZIP-based extended (pseudo) LRT (Chakraborty et al.,, 2022; Huang et al.,, 2011; Ding et al.,, 2020), which ensures FDR control ($\leq 0.05$) with high sensitivity while addressing zero inflation in the counts when present (see Supplement S.1 for a detailed review of competing approaches).
We assess signal detection for all methods using replication-based FDR, sensitivity, average Type I error, and F-score (Figure 3). For each method, these metrics are computed as FDR = $R^{-1}\sum_{r}\bigl{(}\text{FP}_{r}/(\text{FP}_{r}+\text{TP}_{r})\bigr{)}$, sensitivity = $R^{-1}\sum_{r}\bigl{(}\text{TP}_{r}/(\text{TP}_{r}+\text{FN}_{r})\bigr{)}$, average type I error = $R^{-1}\sum_{r}\bigl{(}\text{FP}_{r}/\text{TN}_{r}\bigr{)}$, and F-score = $R^{-1}\sum_{r}\bigl{(}2\text{TP}_{r}/(2\text{TP}_{r}+\text{FN}_{r}+\text{FP}_{r})\bigr{)}$ where $\text{TP}_{r}$ and $\text{FN}_{r}$ denote the number of true signals that are discovered and not discovered respectively, and $\text{FP}_{r}$ and $\text{TN}_{r}$ are the number of true non-signals discovered and not discovered in replicate $r$ based on the signal strength values as cataloged in the corresponding true $\Lambda$.
Figure 3: Simulation results: the horizontal axis shows the true signal strength $\lambda_{0}$ for signal cells, the vertical axis shows the value of evaluation metrics. Each column presents one metric and each row presents one simulation scenario in Table 1.
The first two rows of Figure 3 show comparable (when $\lambda^{0,\text{signal}}<1.5$) or slightly better (when $\lambda^{0,\text{signal}}\geq 1.5$) sensitivity for our local-global methods compared to all other methods when the between-drug association is near zero (Cases 1a,b). Under moderate between-drug association ($\rho=0.43$; Cases 2a,b), our methods perform noticeably better in detecting more true signals, achieving up to twice the sensitivity, particularly for weak or moderate true signal strengths ($1.1\leq\lambda^{0,\text{signal}}\leq 1.7$). The gain is even further pronounced under high between-drug association ($\rho=0.83$; Cases 3a,b), where our methods attain sensitivities of $\geq 0.70$ with only moderate true signal strength ($\lambda^{0,\text{signal}}\geq 1.7$), while all competitors fail to reach $0.70$ sensitivity even for $\lambda^{0,\text{signal}}=3$. These results highlight the substantial boost in signal detection from effectively leveraging between-drug similarities as facilitated by our models. In all settings, the FDR and average type I error is strictly controlled for all methods.
The Poisson and ZIP-based DP local-global models perform similarly across all scenarios regardless of the presence of zero-inflated observed counts (sub-cases a and b), due to the flexibility of the DP-based nonparametric framework (see Remark Remark).
All computations were performed on the cluster computing facility at the Center for Computational Research at the University at Buffalo, with each process running on 16 GB RAM and a clock speed of 1.6-3 GHz. The median computation time per replication for our DP local-global methods (both Poisson and ZIP) was 145 seconds.
5 REAL DATA ANALYSIS
We use the statin data from Section 4 to demonstrate our method’s utility in real-world SRS data mining. Applying our DP local-global models (using the same MCMC and test statistic generation settings as used for simulated data analysis in Section 4) we compare detected signals against those from the DP Hu et al., BCPNN, GPS, and LRT. Table 2 presents the number of signals identified across statin drugs. The results indicate identical performance for our Poisson and ZIP models, detecting more signals than all competing methods.
Table 2: Summary table of discoveries from our proposed DP local-global models, and comparing them to the other existing methods. The count in each cell represents the number of AE signals detected by the method (along the columns), first separately for each drug and then in total (rows). The number of discoveries made uniquely by our DP Poisson local-global method but not by the other methods are presented in parentheses.
Drug DP Poisson DP ZIP DP Hu et al. BCPNN GPS LRT
Atorvastatin 33 33 (0) 33 (0) 28 (5) 26 (7) 28 (5)
Fluvastatin 14 14 (0) 13 (1) 10 (4) 12 (2) 12 (2)
Lovastatin 12 12 (0) 10 (2) 9 (3) 9 (3) 9 (3)
Pravastatin 23 23 (0) 17 (6) 11 (12) 11 (12) 11 (12)
Rosuvastatin 29 29 (0) 28 (1) 25 (4) 25 (4) 25 (4)
Simvastatin 29 29 (0) 28 (1) 26 (3) 25 (4) 25 (4)
Total 140 140 (0) 129 (11) 110 (31) 110 (32) 110 (30)
To further examine the results, we focus on 20 PTs (AEs/$i$) with the highest average (across statin drugs) posterior medians of $\lambda_{ij}$ from our DP Poisson local-global model (Figure 4). The heatmap indicates non-detection (gray) or the degree of information sharing between drugs, measured via the posterior probabilities of $1-z_{ij}^{l}$ (white-blue gradient; darker blue signifies greater information sharing at each AE). Our method detects several rare AEs with $n_{ij}=2$, some also identified by certain competing methods (though no other method detects all these AEs simultaneously) and some not detected by any competing method (not shown in Figure 4; see Table 2). This, along with strong FDR control, highlights our method’s effectiveness in SRS signal detection.
Figure 4: Heatmap of 20 selected PTs. Cells not detected as signals by our method are colored gray. White-to-blue shades represent the posterior probability of $1-z_{ij}^{l}$, with darker (lighter) blue indicating a higher (lower) probability that the signal strength originates from the global signal strength distribution. The letters “H”, “B”, “G”, and “L” indicate signal detection by the DP Hu et al., BCPNN, GPS, and likelihood ratio test, respectively.
Finally, we focus on inferring structural zeros through the estimated drug-specific zero-inflation probabilities $\omega_{j}$ as returned by our DP ZIP local-global model. For comparison, we also monitor the corresponding estimates provided by the frequentist ZIP (pseudo) LRT method. The two methods yield largely similar results: our approach produces point estimates (posterior mean) for $\omega_{j}$ as 0.02, 0.10, 0.11, 0.08, 0.02, and 0.03 for Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, respectively. The corresponding frequentist maximum likelihood estimates are 0.00, 0.07, 0.16, 0.06, 0.00, and 0.00, respectively.
6 DISCUSSION
We propose nonparametric Bayesian models and their MCMC implementation for SRS data mining, balancing flexibility with shrinkage while capturing between-drug associations. Signal detection is performed through hypothesis tests that control FDR and minimize FNR, both estimated from MCMC posteriors. Our approach achieves significantly higher sensitivity than existing methods, particularly for rarer AEs and/or new drugs.
There are several limitations with our current approach that motivate interesting future directions. First, we focus here primarily on signal detection, i.e., to identify critical AEs of concern for each drug. The clusters of AEs produced by our DP priors have not been investigated but may reveal additional, deeper insights by identifying groups of “similar AEs”. A reformulation via mixed membership models could further enhance interpretability by allowing the same AE to appear in multiple clusters/groups. Second, while we consider data-driven characterization of similarities in AE signal strengths between drugs, a more flexible approach acknowledging known Drug-Drug Interactions (DDI; Poleksic and Xie, 2019), possibly via appropriate prior distributions, may further enhance signal detection. Finally, our current framework aggregates drug-AE data across time points for a cross-sectional analysis. Extending the model to accommodate temporal variations in SRS data could enable formal identification of evolving AE patterns.
Acknowledgment. The authors are deeply grateful to Prof. Marianthi Markatou at the University at Buffalo for her insightful comments and helpful suggestions on the manuscript.
Supplementary Material
This supplement provides a brief review of existing SRS data mining methods, additional details about the DP local-global models, detailed steps for the MCMC algorithms implementing our models, additional computation details, and additional simulation results. Sections, tables, and figures in this supplement are labeled with a prefix “S.”, such as, Section S.1, Table S.1, and Figure S.1, etc.
Appendix S.1 A REVIEW OF EXISTING BAYESIAN/EMPIRICAL BAYESIAN SRS DATA MINING METHODS AND THE ZIP MODEL-BASED LIKELIHOOD RATIO TEST METHOD
S.1.1 The GPS model
The GPS method (DuMouchel,, 1999) employs an empirical Bayes approach for stabilized signal detection within a Poisson model-based framework for the report counts $n_{ij}$. Specifically, we assume $n_{ij}\sim$ Poisson$(E_{ij}\lambda_{ij})$, with $\lambda_{ij}$ parametrizing the signal strength for the drug-$j$ and AE-$i$ pair. A two-component gamma mixture prior for $\lambda_{ij}$ is considered:
$\lambda_{ij}\sim\kappa\text{Gamma}(\alpha_{1},\beta_{1})+(1-\kappa)\text{Gamma}(\alpha_{2},\beta_{2}),$
where $\kappa\in[0,1]$ is the mixing weight hyperparameter, and $\alpha_{1},\beta_{1}>0$, and $\alpha_{2},\beta_{2}>0$ are mixture component-specific gamma hyperparemeters. Here Gamma($\alpha$, $\beta$) denotes a probability distribution with density $f_{\text{gamma}}(y\mid\alpha,\beta)=\frac{1}{\Gamma(\alpha)}\beta^{\alpha}y^{\alpha-1}e^{-\beta y}$, $y>0$; $\alpha,\beta>0$. The posterior distribution of $\lambda_{ij}$ given $n_{ij}$, $E_{ij}$, and the hyperparameters is:
$\lambda_{ij}\mid n_{ij},E_{ij}\sim\kappa_{ij}\text{Gamma}(\hat{\alpha_{1}}+n_{ij},\hat{\beta_{1}}+E_{ij})+(1-\kappa_{ij})\text{Gamma}(\hat{\alpha_{2}}+n_{ij},\hat{\beta_{2}}+E_{ij})$
where $\kappa_{ij}=\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})/[\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})+(1-\kappa)f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{2},\beta_{2})]$. Under an empirical Bayes approach, the point estimates $\hat{\kappa}$, $\hat{\alpha}_{1}$, $\hat{\beta}_{1}$, $\hat{\alpha}_{2}$, and $\hat{\beta}_{2}$ for the hyperparameters are obtained by maximizing their marginal likelihoods given $n_{ij}$ and $E_{ij}$. The resulting posterior of $\lambda_{ij}$ conditional on these estimated hyperparameters is used for inference and signal detection. Specifically, DuMouchel and Pregibon, (2001) define the empirical Bayesian score $E(\log_{2}(\lambda_{ij})\mid n_{ij},E_{ij},\hat{\kappa},\hat{\alpha}_{1},\hat{\beta}_{1},\hat{\alpha}_{2},\hat{\beta}_{2})$ as a point estimate of signal strength for the drug-$j$ and AE-$i$ pair, and also suggest determining the pair to be a signal if the 5th percentile (EB05) of the above conditional posterior for $\lambda_{ij}$ exceeds 2.
S.1.2 The BCPNN method
The BCPNN approach (Bate et al.,, 1998) assumes that conditional on the grand total count $n_{\bullet\bullet}$, the cell-specific counts $n_{ij}$, the marginal row totals $n_{i\bullet}$, and the marginal column totals $n_{\bullet j}$ each distributed as binomial-beta: (a) $n_{ij}\mid p_{ij}\sim\text{Binomial}(n_{\bullet\bullet},p_{ij}),\ p_{ij}\sim\text{Beta}(\alpha_{ij},\beta_{ij})$, (b) $n_{i\bullet}\mid p_{i\bullet}\sim\text{Binomial}(n_{\bullet\bullet},p_{i\bullet}),\allowbreak\ p_{i\bullet}\sim\text{Beta}(\alpha_{i\bullet},\beta_{i\bullet})$, (c) $n_{\bullet j}\mid p_{\bullet j}\sim\text{Binomial}(n_{\bullet\bullet},p_{\bullet j}),\allowbreak\ p_{\bullet j}\sim\text{Beta}(\alpha_{\bullet j},\beta_{\bullet j})$, where $p_{ij}$ is the cell reporting rate, and $p_{i\bullet}$ and $p_{\bullet j}$ are the marginal row and column reporting rates. (Note that the binomial model assumptions for $n_{ij}$, $n_{i\bullet}$ and $n_{\bullet j}$ can be understood as consequences of an independent Poisson model for $n_{ij}$ with certain choices of $p_{ij}$, $p_{i\bullet}$, and $p_{\bullet j}$.). Bate et al., (1998) consider the information content $\text{IC}_{ij}=\log_{2}\left[p_{ij}/(p_{i\bullet}p_{\bullet j})\right]$ as a measure of signal strength for the drug-$j$ AE-$i$ pair, with a higher IC indicating a stronger association. The authors suggest an empirical Bayes estimation of these quantities using $\hat{\alpha}_{ij}=\hat{\alpha}_{i\bullet}=\hat{\beta}_{i\bullet}=\hat{\alpha}_{\bullet j}=\hat{\beta}_{\bullet j}=1$ and $\hat{\beta}_{ij}=1/(E(p_{i\bullet}\mid n_{i\bullet})E(p_{\bullet j}\mid n_{\bullet j}))-1$, with the denominator of $\hat{\beta}_{ij}$ being computed using the posterior distributions $p_{i\bullet}\mid n_{i\bullet}\sim\text{Beta}(\hat{\alpha}_{i\bullet}+n_{i\bullet},\hat{\beta}_{i\bullet}+n_{\bullet\bullet}-n_{i\bullet})$ and $p_{\bullet j}\mid n_{\bullet j}\sim\text{Beta}(\hat{\alpha}_{\bullet j}+n_{\bullet j},\hat{\beta}_{\bullet j}+n_{\bullet\bullet}-n_{\bullet j})$. The consequent (empirical Bayes) posterior distribution of $\text{IC}_{ij}$ can be approximated using Monte Carlo simulation or an asymptotic normal distribution. A signal is detected at the drug-$j$ and AE-$i$ pair if the 2.5th posterior percentile for $\text{IC}_{ij}$ exceeds 0.
S.1.3 DP Poisson method of Hu et al., (2015)
Hu et al., (2015) consider the same likelihood $n_{ij}\sim$ Poisson($\lambda_{ij}E_{ij}$) as in GPS, but extend the two-gamma mixture prior for $\lambda_{ij}$ to a DP (infinite) mixture prior:
$\lambda_{ij}\mid(\alpha_{j},M_{0j})\sim\text{DP}(\alpha,M_{0j})=\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}}$
where the base distribution $M_{0j}\equiv$ Gamma$(a_{j},a_{j})$ generates $\theta_{hj}\sim M_{0j}$ and the concentration hyperparameter $\alpha>0$ generates the mixture weights $\{\eta_{hj}\}_{h=1}^{\infty}$ satisfying $\sum_{h=1}^{\infty}\eta_{hj}=1$ with
$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj}),\ h=2,\dots;\ v_{hj}\sim\text{independent Beta}(1,\alpha)\text{ for }h=1,\dots$ (5)
leveraging a stick-breaking representation (Sethuraman,, 1994) for each $j$. Hu et al., (2015) consider full Bayesian inference on $\lambda_{ij}$ by first eliciting a prior distribution on the hyperparameters $a_{j}$ and $\alpha_{j}$—specifically, independent vague priors $a_{j}\sim\text{Uniform}(0,1)$ and $\alpha_{j}\sim\text{Uniform}(0.2,10)$; $j=1,\dots J$—and then utilizing MCMC sampling from the resulting joint posterior. Using MCMC posterior draws for $\lambda_{ij}$, a signal at the drug-$j$ and AE-$i$ pair is identified if the 5th posterior percentile of $\lambda_{ij}$ exceeds a threshold of 2.
S.1.4 ZIP model based Pseudo LRT (frequentist) method
The likelihood ratio test based method and its variants (Huang et al.,, 2011, 2013, 2014) have an advantage in FDR control while maintaining high sensitivity. Additionally, the use of a post-hoc threshold makes it more suitable for practical applications. Consider the reporting count $n_{ij}\sim\text{ZIP}(n=n_{ij}\mid\lambda_{ij}\times E_{ij},\omega_{j})$, where $\lambda_{ij}$ is the relative reporting rate and $E_{ij}=n_{i\cdot}n_{\cdot j}/n_{\cdot\cdot}$ represents the expected count under independence. The pseudo LRT, as proposed by Chakraborty et al., (2022), is a recent method designed to accommodate the zero-inflated nature of SRS data while easing computational demands. Similar to the standard LRT method, this approach first tests a global alternative hypothesis $H_{aj}:\lambda_{ij}>1$ for at least one $i\in\{1,\dots,I\}$ against the null hypothesis $H_{0j}:\lambda_{ij}=1$ for all $i\in\{1,\dots,I\}$ for a given drug-$j$. If $H_{aj}$ is rejected, a step-down procedure is used to conduct individual tests on $H_{aij}:\lambda_{ij}>1$.
The ordinary LRT methods estimate the nuisance parameter $\omega_{j}$ from $H_{0j}$ and $H_{0j}\cup H_{aj}$ individually, the pseudo LRT is constructed using a global estimate $\hat{\omega}_{j}$ of $\omega_{j}$ consistent under $H_{0j}\cup H_{aj}$ in both the numerator and the denominator of the likelihood ratio test statistic. The resulting statistic is given by
$\text{LR}_{ij}=\exp(-(\hat{\lambda}_{ij}-1)E_{ij})\hat{\lambda}_{ij}^{n}.$
The maximum likelihood ratio statistic for drug-$j$ is then defined as $\text{MLR}_{j}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\}\end{subarray}}\text{LR}_{ij}$. In the global test, a parametric bootstrap can be used to approximate the p-value $P_{j}=\Pr(\text{MLR}_{j}\geq\text{MLR}_{j}^{\text{obs}})$ using (parametric bootstrap) draws for $\text{LR}_{ij}$ under $H_{0j}$ and hence $\text{MLR}_{j}$, where $\text{MLR}_{j}^{\text{obs}}$ is the observed test statistic. If the global test is rejected, the step-down p-values can be computed for the individual tests using the same parametric bootstrap samples for $\text{MLR}_{j}$ and using the individual observed likelihood ratio test statistic values, such that $P_{ij}=\Pr(\text{LR}_{ij}\geq\text{MLR}_{ij}^{\text{obs}})$. For multiple drugs, the same approach can be applied, defining an overall maximum likelihood ratio as $\text{MLR}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\},j\in\{1,\dots,J\}\end{subarray}}\text{LR}_{ij}$. This step-down procedure avoids $\alpha$-inflation by using the same global test statistic, thus controlling the false discovery rate at the $\alpha$ level.
A consistent estimate of $\omega_{j}$ is obtained by maximizing the profile likelihood $\ell_{j}(\omega_{j})$, with $\hat{\lambda}_{ij}=\max\{1,n/E_{ij}\}$ as the MLE of $\lambda_{ij}$. The inference on zero-inflation is based on the estimate $\hat{\eta}_{ij}(\lambda_{ij},\hat{\omega}_{j})=\frac{\hat{\omega}_{j}}{\hat{\omega}_{j}+(1-\hat{\omega}_{j})\exp(-\lambda_{ij}E_{ij})}$, representing the estimated probability of zero-inflation for cells with zero count. A prediction of the zero-inflation indicator $z_{ij}$ can be made by generating Bernoulli samples from $\hat{\eta}_{ij}$.
Appendix S.2 ADDITIONAL DETAILS FOR THE DP LOCAL-GLOBAL FRAMEWORK
In this section, we generalize the model to encompass different variants depending on the direction of information sharing. Moreover, we provide additional details of the DP local-global framework. And we propose a simpler shrinkage model as an alternative for the “Other drugs” as described in Section 3.2.
S.2.1 Complete hierarchical Bayesian representation of our models
In Section 3.2 in the main text, we present the concept of the DP Poisson local-global model. The fully formulated hierarchical representations are provided in the following two sections for the DP Poisson and the DP ZIP local-global model.
S.2.1.1 DP Poison local-global model
A hierarchical representation of the complete model, including vague priors for the hyperparameters, is presented as follows.
likelihood:
$n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$
[…]
$\lambda_{ij}\sim z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$
$\lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D$
$z_{ij}^{l}\sim\text{Bernoulli}(\pi^{l})$
DP local prior for drug-$j$:
$D_{j}\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}}\approx\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$
$\theta_{hj}\sim G(\beta_{j})\equiv\text{Gamma}(\text{shape}=\beta_{j},\text{rate}=\beta_{j})$
$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj}),h\geq 2$
$v_{hj}\sim\text{Beta}(1,\alpha_{j}),h=1,\dots,K-1;\ v_{Kj}=1$
DP global prior:
$D\equiv\sum_{h=1}^{\infty}\eta_{h}\delta_{\{\theta_{h}\}}\approx\sum_{h=1}^{K}\eta_{h}\delta_{\{\theta_{h}\}}\sim DP(\alpha,G(\beta))$
$\theta_{h}\sim G(\beta)\equiv\text{Gamma}(\text{shape}=\beta,\text{rate}=\beta)$
$\eta_{1}=v_{1},\ \eta_{h}=v_{h}\prod_{t=1}^{h-1}(1-v_{t}),h\geq 2$
$v_{h}\sim\text{Beta}(1,\alpha)$
Gamma shrinkage prior for a baseline drug-$J$:
$\lambda_{iJ}\sim\text{Gamma}(\text{shape}=\tau,\text{rate}=\tau)$
vague proper priors for hyperparameters:
$\pi^{g}\sim\text{Uniform}(0,1)$
$\alpha_{j},\alpha\sim\text{Gamma}(\text{shape}=1,\text{rate}=3)$
$\beta_{j},\beta\sim\text{Cauchy}^{+}(0,0.5)$
$\tau\sim\text{Cauchy}^{+}(0,0.5)$
S.2.1.2 DP ZIP local-global model
A hierarchical representation of the complete model, including vague priors for the hyperparameters, is presented as follows.
likelihood:
$n_{ij}\sim y\text{Poisson}(\lambda_{ij}E_{ij})+(1-y)\delta_{\{0\}}$
$y\sim\text{Bernoulli}(\omega_{j})$
[…]
$\lambda_{ij}\sim z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$
$\lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D$
$z_{ij}^{l}\sim\text{Bernoulli}(\pi^{l})$
DP local prior for drug-$j$:
$D_{j}\equiv DP(\alpha_{j},G(\beta_{j}))\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}}\approx\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$
$\theta_{hj}\sim G_{0j}\equiv\text{Gamma}(\text{shape}=\beta_{j},\text{rate}=\beta_{j})$
$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj}),h\geq 2$
$v_{hj}\sim\text{Beta}(1,\alpha_{j});v_{Kj}=1$
DP global prior:
$D\equiv DP(\alpha,G(\beta))\equiv\sum_{h=1}^{\infty}\eta_{h}\delta_{\{\theta_{h}\}}\approx\sum_{h=1}^{K}\eta_{h}\delta_{\{\theta_{h}\}}$
$\theta_{h}\sim G_{0}\equiv\text{Gamma}(\beta,\beta)$
$\eta_{1}=v_{1},\ \eta_{h}=v_{h}\prod_{t=1}^{h-1}(1-v_{t}),h\geq 2$
$v_{h}\sim\text{Beta}(1,\alpha)$
Gamma shrinkage prior for a baseline drug-$J$:
$\lambda_{iJ}\sim\text{Gamma}(\text{shape}=\tau,\text{rate}=\tau)$
vague proper priors for hyperparameters:
$\pi^{g}\sim\text{Uniform}(0,1)$
$\alpha_{j},\alpha\sim\text{Gamma}(\text{shape}=1,\text{rate}=3)$
$\beta_{j},\beta\sim\text{Cauchy}^{+}(0,0.5)$
$\tau\sim\text{Cauchy}^{+}(0,0.5)$
S.2.2 Some notes on the “global-only” model
A special case of our local-global models leads to a reduced “global-only” model when $\pi^{g}=0$, i.e., $\lambda_{ij}=\lambda_{i}^{g}$ for $j=1,\dots,J-1$. Under this model there is no provision for different $\lambda_{ij}$ for any $j\neq J$. The global model has a simplified likelihood $n_{ij}\sim\text{Poisson}(\lambda_{i}E_{ij})$ here $\lambda_{i}^{g}\sim D$ if $j\neq J$, which produces $\sum_{j=1}^{J-1}n_{ij}\sim\text{Poisson}(\lambda_{i}\sum_{j=1}^{J-1}E_{ij})$. The model reduction also impacts the signal detection. For the local-global framework, hypothesis testing is conducted for each individual $\lambda_{ij}$. On the other hand, under the global model, the test is simplified to evaluating whether $\lambda_{i}^{g}>1$. Testing $\lambda_{i}^{g}>1$ entails assessing if the summarized Poisson count $\sum_{i}n_{ij}$ is large compared to the expected count $\sum_{i}E_{ij}$, owing to the reduced model. This approach represents a “global test”, whereas our proposed local-global model allows conducting hypothesis testing for each individual drug-AE pair. As a result, directly comparing the global model with the local-global model is not feasible, as they target different types of assessment.
Appendix S.3 COMPUTATIONAL DETAILS
In this section, we first provide MCMC algorithms implementing our models in Section S.3.1. The MCMC algorithms are derived using finite mixture approximation of the underlying DP processes (Ishwaran and Zarepour,, 2002; Frühwirth-Schnatter and Malsiner-Walli,, 2019), and combine Gibbs sampling updates with stepping-out slice sampling updates (Neal,, 2003) for various parameters. Some additional notes on the MCMC implementation, including initial value generation for our MCMC algorithms are then provided in Section S.3.2. Next we provide detailed notes on the grid-search method identifying optimal quantile-level $p$ and threshold $k$ used for our hypothesis tests for signal detection (Section S.3.3). The JAGS code (Plummer et al.,, 2003) for Hu et al., (2015)’s models is shown as a reference in Section S.1.3. The example statin dataset is included in Section S.3.5.
In our notation below, $\text{DPSB}_{K}(\bm{\eta}=(\eta_{1},\dots,\eta_{K}),\bm{\theta}=(\theta_{1},\dots,\theta_{K}),\alpha,\beta)$ denotes parameters associated with the model $\sum_{h=1}^{K}\eta_{h}1_{\{\theta_{h}\}}\sim\text{DP}(\alpha,G(\beta))$, comprising the stick-breaking weights $\bm{\eta}=(\eta_{1},\dots,\eta_{K})$ and atoms $\bm{\theta}=(\theta_{1},\dots,\theta_{K})$ and DP concentration hyperparamter $\alpha$ and base (gamma) distribution hyperparameter $\beta$.
S.3.1 MCMC algorithms
Algorithm S.1 One Iteration of the MCMC algorithm updating $\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$ parameters from its posterior given data $(n_{i},E_{i},z_{i})$ under the model: $n_{i}\sim\text{Pois}(\lambda_{i}E_{i})$, $\lambda_{i}\sim\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$, $\alpha\sim\text{Exponential}(\text{rate}=\psi_{\alpha})$, $\beta\sim\pi_{\beta}(\beta)$, and $z_{i}$ indicating if observation $i$ contributes to the model ($z_{i}=1$) or not ($z_{i}=0$); $i=1,\dots,N$.
  1. 1.
    Generate allocations $\bm{S}=(S_{1},\dots,S_{N})$ using
    $P(S_{i}=h\mid\bm{\eta},\bm{\theta},\bm{z},\bm{n},\bm{E})=\frac{\eta_{h}\{f_{\text{Poisson}}(n_{i}\mid\theta_{h}E_{i})\}^{z_{i}}}{\sum_{h^{\prime}=1}^{K}\eta_{h^{\prime}}\{f_{\text{Poisson}}(n_{i}\mid\theta_{h^{\prime}}E_{i})\}^{z_{i}}}$
    for $h=1,\dots,K$. Here $f_{\text{Poisson}}(n\mid\theta)$ denotes the probability mass function of the Poisson($\theta$) distribution evaluated at $n$.
  2. 2.
    Generate $\theta_{1},\dots,\theta_{K}$ independently from
    $\theta_{h}\mid\bm{S},\alpha,\bm{n},\bm{E},\bm{z}\sim\text{Gamma}\!\Biggl{(}\text{shape}=\sum_{i=1}^{N}n_{i}z_{i}1(S_{i}=h)+\alpha,\text{rate}=\sum_{i=1}^{N}E_{i}z_{i}1(S_{i}=h)+\alpha\Biggr{)}$
    1. 2.1.
      Calculate $\lambda_{i}=\theta_{S_{i}}$, $i=1,\dots,N$
  3. 3.
    Generate $\beta$ from the density
    $\pi(\beta\mid\bm{\theta})\propto\prod_{h=1}^{K}f_{\text{Gamma}}(\theta_{h}\mid\text{shape}=\beta,\text{rate}=\beta)\pi_{\beta}(\beta)$
    Here $f_{\text{Gamma}}(\theta\mid\text{shape}=\beta,\text{rate}=\beta)$ denotes the probability density function of the Gamma($\beta,\beta$) distribution at $\theta$. $\pi(\beta)$ is the prior distribution of $\beta$. Direct random sampling from this density is not feasible. Instead, we use stepping-out slice sampling (Neal,, 2003) to generate Markov chain draws for $\beta$.
  4. 4.
    Generate $v_{1},\dots,v_{K-1}$ independently from:
    $v_{h}\mid\bm{S},\alpha\sim\text{Beta}\left(1+\sum_{i=1}^{N}1(S_{i}=k),\alpha+\sum_{i=1}^{N}1(S_{i}>k)\right).$
    Then set $v_{K}=1$ and calculate $\eta_{1},\dots,\eta_{K}$ from $v_{1},\dots,v_{K}$ using
    $\eta_{1}=v_{1},\quad\eta_{h}=v_{h}\prod_{t=1}^{h-1}(1-v_{t}),\quad h=2,\dots,K.$
  5. 5.
    Generate $\alpha$ from
    $\alpha\mid v_{1},\dots,v_{K}\sim\text{Gamma}\left(\text{shape}=K,\text{rate}=-\sum_{h=1}^{K-1}\log(1-v_{h})+\psi_{\alpha}\right).$
Algorithm S.2 One Step Iteration of DP Poisson local-global column model (Given data $\{(n_{ij},E_{ij},z_{ij}^{l}):i=1,\dots,I;j=1,\dots,J\}$ )
  1. 1.
    Generate global $\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$ parameters using Algorithm S.1, with $\{(\tilde{n}_{i},\tilde{E}_{i},1(\sum_{j=1}^{J-1}z_{ij}^{l}<J-1)):i=1,\dots,I\}$ as data where
    $\tilde{n}_{j}=\sum_{j=1}^{J-1}n_{ij}(1-z_{ij}^{l}),\quad\tilde{E}_{j}=\sum_{j=1}^{J-1}E_{ij}(1-z_{ij}^{l}).$
    1. 1.1.
      Set $\lambda_{i}^{\text{g}}=\lambda_{i}$ returned from Algorithm S.1.
  2. 2.
    For $j=1,\dots,J-1$, generate local $\text{DPSB}(\bm{\eta}_{j},\bm{\theta}_{j},\alpha_{j},\beta_{j})$ using Algorithm S.1 using $\{(n_{ij},E_{ij},z_{ij});i=1,\dots,I\}$ as data.
    1. 2.1.j
      For this $j$, set $\lambda_{ij}^{l}=\lambda_{i}$ returned from Algorithm S.1.
  3. 3.
    Generate “local” allocations:
    $z_{ij}^{l}\mid n_{ij},E_{ij},z_{ij}^{l},\lambda_{ij}^{l},\lambda_{i}^{g}\sim\text{Bernoulli}(p_{ij}^{l}),$
    where
    $p_{ij}^{l}=\frac{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})}{f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})+(1-\pi^{l})f_{\text{Poisson}}(n_{ij}\mid\lambda_{i}^{g}E_{ij})}.$
  4. 4.
    Define $\lambda_{ij}=z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$ for all $i=1,\dots,I$ and $j=1,\dots,(J-1)$.
  5. 5.
    Generate $\pi^{l}$ using:
    $\pi^{l}\mid z_{ij}^{l}\sim\text{Beta}\left(a^{l}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}z_{ij},b^{l}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}(1-z_{ij})\right),$
    Notes: $a^{l}$ = 1, $b^{l}$ = 1
  6. 6.
    Generate $\lambda_{1J},\dots,\lambda_{IJ}$ from:
    $\lambda_{iJ}\mid\tau,n_{iJ},E_{iJ}\sim\text{Gamma}(\text{shape}=n_{iJ}+\tau,\text{rate}=E_{iJ}+\tau).$
  7. 7.
    Generate $\beta_{J}$ from:
    $\pi(\beta_{J}\mid\lambda_{1J},\dots,\lambda_{IJ})\propto\prod_{i=1}^{I}f_{\text{Gamma}}(\lambda_{ij}\mid\text{shape}=\tau,\text{rate}=\tau)\pi_{\tau}(\tau).$
Algorithm S.3 One Iteration of DP ZIP local-global Model Given data $(n_{ij},E_{ij},z_{ij}^{l})$
  1. 1.
    Generate the ZI indicator $y_{ij}$ for $i=1,\dots,I$, $j=1,\dots,J$, where:
    $y_{ij}\mid\omega_{j},\lambda_{ij},n_{ij},E_{ij}\sim\text{Bernoulli}(p_{ij}^{\text{ZI}}),$
    where
    $p_{ij}^{\text{ZI}}=\begin{cases}\frac{\omega_{j}}{\omega_{j}+(1-\omega_{j})\exp(-\lambda_{ij}E_{ij})},&\text{if }n_{ij}=0,\\ 0,&\text{if }n_{ij}\geq 1.\end{cases}$
  2. 2.
    Generate $\omega_{1},\dots,\omega_{J}$ independently with:
    $\omega_{j}\mid y_{1j}\dots y_{Ij}\sim\text{Beta}\left(1+\sum_{i=1}^{I}y_{ij},1+\sum_{i=1}^{I}(1-y_{ij})\right).$
  3. 3.
    Run Algorithm S.2 with all $(n_{ij},E_{ij})$ such that $y_{ij}=0$.
S.3.2 Initial value generation for the MCMC algorithms
Given the observed data $n_{ij}$ and $E_{ij}$ for $i=1,\dots,I$, $j=1,\dots,J$, and a prespecified number $K$ of DP mixture components for finite approximation, we initialize the number of nonempty components $K_{+}$ to the smallest integer greater than $\min\{K/2,I-1\}$. The local-global indicators $z_{ij}^{l}$ are randomly initialized by independently sampling each cell $(i,j)$ from a $\text{Bernoulli}(0.5)$ distribution. The DP parameters for each local and global DP process are then initialized using the subset of data $(n_{ij},E_{ij})$ assigned to that process, as determined by the indicators $z_{ij}^{l}$, for $i=1,\dots,J-1$.
For the initialization of parameters in the local DP prior for the $j$-th drug ($j=1,\dots,J-1$), we perform $k$-means clustering with the initialized $K_{+}$ number of clusters on the set ${(n_{ij}+0.5)/(E_{ij}+0.5):z_{ij}=1,\ i=1,\dots,I}$, and use the resulting cluster centers as initial values for the atoms $\theta_{hj}$, $h=1,\dots,K_{+}$, of the non-empty clusters. The probability weights $\eta_{hj}$ for these non-empty clusters are initialized using the relative frequencies of the $k$-means clusters, and the clusters are then relabeled to ensure $\eta_{1j}\geq\eta_{2j}\geq\dots\geq\eta_{K_{+}j}$. These atoms $\theta_{hj}$ are then used for numerical maximum likelihood estimation of $\beta$ with the model $\theta_{1j},\dots,\theta_{K_{+},j}\sim\text{independent~{}Gamma}(\text{shape}=\beta,\text{rate}=\beta)$, and the resulting estimate $\hat{\beta}$ is used to initialize $\beta_{j}$. The atoms $\theta_{h^{\prime}j}$ for the remaining empty clusters ($h^{\prime}=K_{+}+1,\dots,K$) are randomly drawn from $\text{Gamma}(\text{shape}=\beta_{j},\text{rate}=\beta_{j})$ using the initialized $\beta_{j}$. Their corresponding weights $\eta_{h^{\prime}j}$ are initialized to $(\min_{h^{\prime}\in{1,\dots,K_{+}}}\eta_{h^{\prime}j})/10$, and all weights $\eta_{hj}$ for $h=1,\dots,K$ are subsequently renormalized to ensure $\sum_{h=1}^{K}\eta_{hj}=1$. The local signal strengths $\lambda_{ij}^{l}$ are then initialized for each $i$ using the cluster center (atom) corresponding to the $k$-means cluster containing point $i$ if $z_{ij}=1$, or by random sampling from ${\theta_{1j},\dots,\theta_{Kj}}$ with probabilities ${\eta_{1j},\dots,\eta_{Kj}}$. Finally, the concentration hyperparameter $\alpha_{j}$ is initialized by random sampling from a $\text{Gamma}(\text{shape}=1,\text{rate}=2)$ distribution.
The parameters for the global DP prior are analogously initialized, but using the aggregated observed and expected counts $\{\tilde{n}_{i}=\sum_{j=1}^{J-1}n_{ij}(1-z_{ij}),\tilde{E}_{i}=\sum_{j=1}^{J-1}E_{ij}(1-z_{ij})\}$ as the data with all AEs $i$ such that $\sum_{j=1}^{J-1}(1-z_{ij})>0$.
Finally, the reference drug $J$ shrinkage hyperparameter $\tau$ is initialized at $1$, while the zero-inflation parameters $\omega_{j}$ (if included in the model) are initialized each at $0.01$.
S.3.3 Optimization method in pFDR calculation
A sequence of threshold values for $k$, ranging from 1.1 to 3 in small increments, is established. For our computations we specified this grid to be {1.1, 1.11, 1.12, 1.13, 1.14, 1.16, 1.17, 1.18, 1.19, 1.2, 1.25, 1.28, 1.31, 1.33, 1.36, 1.39, 1.42, 1.44, 1.47, 1.5, 1.6, 1.76, 1.91, 2.07, 2.22, 2.38, 2.53, 2.69, 2.84, 3}. Similarly, a set of distinct small quantile levels for $p$ is chosen ({0.01, 0.015, 0.02, 0.025, 0.03, 0.035, 0.04, 0.045, 0.05, 0.055, 0.06, 0.065, 0.07, 0.075, 0.08, 0.085, 0.09, 0.095, 0.1} in our computations). These define a range of possible combinations $(p,k)$ to be considered.
For every quantile level in the $p$-grid, we obtain a matrix of test statistics $T_{ij}(p)$ based on the $p$-th posterior quantiles for $\lambda_{ij}$ computed from their MCMC draws. Subsequently, we compute FDR$(p,k)$ and FNR$(p,k)$ estimates using the formulas presented in the main text (Section 3.4) for all possible $p$ and $k$ defined by the grid. From the set of all $(p,k)$ pairs, we then retain only those pairs with computed FDR $\leq\alpha$ for a preset threshold $\alpha$ ($=0.05$ in our analyses). Finally, among the retained $(p,k)$, we identify the pair that minimizes FNR. If multiple $(p,k)$ yield the same minimum FNR, we select the pair with the smallest $p$ and the largest $k$.
S.3.4 JAGS code for DP Hu et al. model
1 # Likelihood
2 for (i in 1:I) {
3 for (j in 1:J) {
4 n[i, j] ~ dpois(E[i, j] * lambda[i, j])
5 }
6 }
7
8 # Prior
9 for (j in 1:J) {
10 alpha[j] ~ dunif(0.2, 10)
11 a[j] ~ dunif(0, 1)
12
13 v[j, 1] ~ dbeta(1, alpha[j]); T(0.00001,0.99999) # to ensure in (0, 1)
14 w[j, 1] <- v[j, 1]
15
16 for (h in 2:K) {
17 v[j, h] ~ dbeta(1, alpha[j]); T(0.00001,0.99999) # to ensure in (0, 1)
18 w[j, h] <- v[j, h] * prod( (1 - c(v[j, 1:(h-1)])) )
19 }
20
21 for (h in 1:K) {
22 theta[j, h] ~ dgamma(a[j], a[j]); T(0.000001, ) # to ensure > 0
23 }
24
25 for (i in 1:I) {
26 ind[i, j] ~ dcat(c(w[j,]))
27 lambda[i, j] <- theta[j, ind[i, j]]
28 }
29 }
Listing 1: JAGS exmpale
S.3.5 Example statin dataset
We present the reference SRS contingency table for statin drug in Table S.1. It is used for data generation in Section 4 and data analysis in Section 5.
Table S.1: The SRS contingency table summarized from the FDA FEARS database (Q3 2014-Q4 2020) which catalogs a total of 63,976,610 reported cases.
Adverse Event Atorvastatin Fluvastatin Lovastatin Pravastatin Rosuvastatin Simvastatin Other
Acute Kidney Injury 1353 42 7 154 689 823 355651
Anuria 71 0 0 2 43 62 10403
Blood Calcium Decreased 14 2 0 0 110 17 15918
Blood Creatine Phosphokinase Abnormal 34 0 0 0 8 11 261
Blood Creatine Phosphokinase Increased 1175 125 32 200 562 768 23805
Blood Creatine Phosphokinase Mm Increased 2 0 0 0 9 0 14
Blood Creatinine Abnormal 27 0 0 0 5 3 3385
Blood Creatinine Increased 227 10 0 17 210 97 74742
Chromaturia 340 10 6 33 174 114 19294
Chronic Kidney Disease 152 16 2 19 177 37 339179
Compartment Syndrome 53 0 0 1 12 12 2644
Creatinine Renal Clearance Decreased 6 0 0 2 124 6 7768
Diaphragm Muscle Weakness 14 0 0 0 8 1 94
Electromyogram Abnormal 2 0 0 2 0 0 132
End Stage Renal Disease 30 0 0 0 19 6 97553
Glomerular Filtration Rate Abnormal 8 0 0 0 1 0 1069
Glomerular Filtration Rate Decreased 59 1 0 8 39 29 13190
Hypercreatininaemia 0 0 0 0 8 0 648
Hypocalcaemia 36 0 0 16 8 18 23102
Muscle Disorder 291 2 7 21 191 87 7329
Muscle Enzyme Increased 48 1 0 0 13 9 410
Muscle Fatigue 85 0 2 16 30 39 4257
Muscle Haemorrhage 24 0 0 5 13 4 3806
Muscle Necrosis 68 2 0 1 10 20 662
Muscle Rupture 181 25 0 61 36 120 3219
Muscular Weakness 1857 45 31 152 808 859 111003
Musculoskeletal Discomfort 137 18 15 25 187 93 19931
Musculoskeletal Disorder 56 3 0 9 65 73 25881
Musculoskeletal Pain 420 3 2 38 324 228 82576
Myalgia 5362 341 151 939 2757 3216 143819
Myasthenic Syndrome 1 0 9 0 3 6 643
Myoglobin Blood Increased 71 4 0 4 16 39 1003
Myoglobin Blood Present 2 0 0 0 0 0 8
Myoglobin Urine Present 0 0 0 0 1 2 70
Myoglobinaemia 15 0 0 0 0 0 62
Myoglobinuria 26 4 0 1 1 9 296
Myopathy 849 64 45 145 377 544 6695
Myopathy Toxic 31 0 0 1 2 21 457
Myositis 219 8 10 28 62 141 7482
Necrotising Myositis 279 0 0 2 10 52 278
Oliguria 52 0 0 4 24 37 7590
Renal Failure 534 26 11 69 225 195 250710
Renal Impairment 390 52 11 37 161 169 103343
Renal Tubular Necrosis 40 0 0 12 10 24 12762
Rhabdomyolysis 2041 52 44 163 936 1376 31707
Tendon Discomfort 9 0 0 3 22 10 794
Other Pt 180699 4886 2845 20296 113960 76068 61724222
Appendix S.4 ADDITIONAL SIMULATION RESULTS
We present detailed values of the simulation to compare among our drug (column)-based models (Poisson/ZIP), DP Hu et al., and posterior quantile optimized DP Hu et al. method. Four evaluation metrics (FDR, sensitivity, average type I error, and F-score) are evaluated as well as the main text with same simulation settings in Table 1. Frequentist and Bayesian MSE are compared to further evaluate the loss of the estimation.
S.4.1 Simulation results for DP-based models
We present the four evaluation metrics in Tables S.2, S.3, S.4, and S.5 to provide a clearer numerical comparison beyond Figure 3. Additionally, we further explored the effect of posterior quantile optimization on the DP Hu et al. in the process of FDR control and minimizing FNR. This allows for a more comprehensive understand the strategy of optimizing the posterior quantile of the test statistic and deterministic threshold simultaneously.
We found that in the DP Hu et al. method, simultaneously optimizing the posterior quantile results in only a slight increase in FDR, which remains well-controlled within a very small level (around 0.01). In return, this optimization leads to an improvement in sensitivity, benefiting all cases. When comparing across all DP-based models, the sensitivity gain improvement brought by incorporating the local-global structure is more significant than from this optimization.
Table S.2: Simulation of FDR for DP-based models
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0035 0.0022 0.0050 0.0102
1.2 0.0072 0.0063 0.0056 0.0171
1.3 0.0081 0.0088 0.0053 0.0172
1.4 0.0074 0.0077 0.0061 0.0168
1.5 0.0081 0.0079 0.0059 0.0149
1.7 0.0057 0.0057 0.0050 0.0133
1.9 0.0049 0.0047 0.0039 0.0108
2.0 0.0053 0.0056 0.0046 0.0103
2.5 0.0045 0.0046 0.0043 0.0106
3.0 0.0031 0.0029 0.0034 0.0086
Case 1b Weak correlation ZI 1.1 0.0010 0.0017 0.0060 0.0197
1.2 0.0044 0.0034 0.0096 0.0214
1.3 0.0067 0.0070 0.0070 0.0156
1.4 0.0086 0.0085 0.0051 0.0157
1.5 0.0089 0.0092 0.0059 0.0138
1.7 0.0067 0.0065 0.0056 0.0136
1.9 0.0058 0.0057 0.0048 0.0117
2.0 0.0049 0.0050 0.0045 0.0118
2.5 0.0032 0.0034 0.0032 0.0089
3.0 0.0036 0.0037 0.0042 0.0095
Case 2a Moderate correlation No ZI 1.1 0.0000 0.0010 0.0065 0.0193
1.2 0.0107 0.0128 0.0063 0.0150
1.3 0.0114 0.0132 0.0069 0.0160
1.4 0.0120 0.0134 0.0060 0.0146
1.5 0.0095 0.0098 0.0069 0.0157
1.7 0.0072 0.0079 0.0059 0.0145
1.9 0.0052 0.0056 0.0058 0.0128
2.0 0.0042 0.0045 0.0048 0.0125
2.5 0.0035 0.0035 0.0041 0.0102
3.0 0.0033 0.0032 0.0032 0.0093
Case 2b Moderate correlation ZI 1.1 0.0000 0.0000 0.0055 0.0122
1.2 0.0093 0.0106 0.0065 0.0186
1.3 0.0117 0.0130 0.0069 0.0168
1.4 0.0102 0.0114 0.0061 0.0153
1.5 0.0066 0.0076 0.0054 0.0149
1.7 0.0053 0.0057 0.0058 0.0134
1.9 0.0039 0.0040 0.0059 0.0126
2.0 0.0047 0.0050 0.0045 0.0110
2.5 0.0039 0.0037 0.0042 0.0103
3.0 0.0030 0.0032 0.0041 0.0094
Case 3a Strong correlation No ZI 1.1 0.0000 0.0000 0.0073 0.0133
1.2 0.0084 0.0114 0.0074 0.0183
1.3 0.0073 0.0084 0.0080 0.0184
1.4 0.0086 0.0096 0.0069 0.0145
1.5 0.0061 0.0069 0.0067 0.0150
1.7 0.0045 0.0050 0.0066 0.0136
1.9 0.0045 0.0050 0.0051 0.0129
2.0 0.0039 0.0042 0.0049 0.0118
2.5 0.0033 0.0035 0.0037 0.0096
3.0 0.0030 0.0029 0.0031 0.0085
Case 3b Strong correlation ZI 1.1 0.0000 0.0000 0.0040 0.0171
1.2 0.0106 0.0120 0.0084 0.0162
1.3 0.0077 0.0090 0.0063 0.0161
1.4 0.0078 0.0083 0.0061 0.0153
1.5 0.0068 0.0073 0.0051 0.0125
1.7 0.0044 0.0049 0.0049 0.0119
1.9 0.0030 0.0032 0.0050 0.0112
2.0 0.0035 0.0038 0.0052 0.0113
2.5 0.0033 0.0034 0.0040 0.0092
3.0 0.0036 0.0037 0.0039 0.0094
Table S.3: Simulation of sensitivity for DP-based models
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0016 0.0014 0.0009 0.0021
1.2 0.0427 0.0422 0.0346 0.0435
1.3 0.1227 0.1243 0.0899 0.1041
1.4 0.1923 0.1942 0.1432 0.1606
1.5 0.2444 0.2472 0.1915 0.2115
1.7 0.3153 0.3180 0.2685 0.2897
1.9 0.3692 0.3707 0.3258 0.3475
2.0 0.3926 0.3957 0.3495 0.3720
2.5 0.4706 0.4729 0.4346 0.4554
3.0 0.5251 0.5273 0.4920 0.5128
Case 1b Weak correlation ZI 1.1 0.0012 0.0012 0.0011 0.0022
1.2 0.0406 0.0399 0.0331 0.0416
1.3 0.1216 0.1232 0.0848 0.1006
1.4 0.1948 0.1968 0.1401 0.1581
1.5 0.2426 0.2451 0.1888 0.2102
1.7 0.3182 0.3196 0.2675 0.2898
1.9 0.3709 0.3733 0.3261 0.3457
2.0 0.3921 0.3939 0.3467 0.3678
2.5 0.4717 0.4742 0.4327 0.4564
3.0 0.5227 0.5253 0.4910 0.5124
Case 2a Moderate correlation No ZI 1.1 0.0061 0.0060 0.0012 0.0026
1.2 0.2096 0.2170 0.0526 0.0674
1.3 0.3304 0.3377 0.1334 0.1530
1.4 0.3991 0.4036 0.1998 0.2210
1.5 0.4489 0.4537 0.2499 0.2751
1.7 0.5172 0.5197 0.3359 0.3637
1.9 0.5663 0.5681 0.4043 0.4323
2.0 0.5837 0.5847 0.4310 0.4589
2.5 0.6396 0.6411 0.5363 0.5608
3.0 0.6734 0.6746 0.5984 0.6214
Case 2b Moderate correlation ZI 1.1 0.0067 0.0072 0.0011 0.0025
1.2 0.2067 0.2133 0.0523 0.0659
1.3 0.3309 0.3367 0.1348 0.1535
1.4 0.3968 0.4021 0.1973 0.2186
1.5 0.4438 0.4477 0.2475 0.2723
1.7 0.5138 0.5157 0.3340 0.3599
1.9 0.5601 0.5620 0.4000 0.4274
2.0 0.5772 0.5790 0.4297 0.4570
2.5 0.6329 0.6339 0.5313 0.5561
3.0 0.6696 0.6708 0.5967 0.6205
Case 3a Strong correlation No ZI 1.1 0.0065 0.0076 0.0009 0.0020
1.2 0.2680 0.2784 0.0329 0.0455
1.3 0.4595 0.4682 0.1088 0.1313
1.4 0.5675 0.5746 0.1807 0.2100
1.5 0.6272 0.6330 0.2480 0.2791
1.7 0.7015 0.7049 0.3546 0.3868
1.9 0.7460 0.7482 0.4300 0.4612
2.0 0.7629 0.7644 0.4608 0.4940
2.5 0.8096 0.8106 0.5793 0.6094
3.0 0.8330 0.8336 0.6530 0.6799
Case 3b Strong correlation ZI 1.1 0.0083 0.0094 0.0010 0.0021
1.2 0.2627 0.2769 0.0338 0.0471
1.3 0.4588 0.4688 0.1098 0.1326
1.4 0.5638 0.5717 0.1826 0.2115
1.5 0.6275 0.6333 0.2492 0.2801
1.7 0.7044 0.7080 0.3534 0.3848
1.9 0.7485 0.7507 0.4276 0.4599
2.0 0.7630 0.7651 0.4574 0.4893
2.5 0.8100 0.8109 0.5787 0.6097
3.0 0.8341 0.8348 0.6556 0.6825
Table S.4: Simulation of average type I error for DP-based models
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0000
1.2 0.0001 0.0001 0.0001 0.0002
1.3 0.0002 0.0003 0.0001 0.0003
1.4 0.0002 0.0002 0.0001 0.0004
1.5 0.0003 0.0003 0.0002 0.0005
1.7 0.0002 0.0003 0.0002 0.0005
1.9 0.0002 0.0002 0.0002 0.0005
2.0 0.0003 0.0003 0.0002 0.0005
2.5 0.0003 0.0003 0.0002 0.0006
3.0 0.0002 0.0002 0.0002 0.0006
Case 1b Weak correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0001 0.0001 0.0001 0.0002
1.3 0.0002 0.0002 0.0001 0.0002
1.4 0.0003 0.0003 0.0001 0.0004
1.5 0.0003 0.0004 0.0002 0.0004
1.7 0.0003 0.0003 0.0002 0.0005
1.9 0.0003 0.0003 0.0002 0.0005
2.0 0.0002 0.0002 0.0002 0.0006
2.5 0.0002 0.0002 0.0002 0.0005
3.0 0.0003 0.0003 0.0003 0.0006
Case 2a Moderate correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0005 0.0007 0.0001 0.0002
1.3 0.0006 0.0008 0.0002 0.0004
1.4 0.0008 0.0009 0.0002 0.0005
1.5 0.0006 0.0007 0.0002 0.0006
1.7 0.0005 0.0006 0.0003 0.0007
1.9 0.0004 0.0004 0.0003 0.0007
2.0 0.0003 0.0004 0.0003 0.0008
2.5 0.0003 0.0003 0.0003 0.0007
3.0 0.0003 0.0003 0.0002 0.0007
Case 2b Moderate correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0005 0.0006 0.0001 0.0002
1.3 0.0007 0.0007 0.0001 0.0004
1.4 0.0006 0.0007 0.0002 0.0005
1.5 0.0004 0.0005 0.0002 0.0006
1.7 0.0004 0.0004 0.0003 0.0006
1.9 0.0003 0.0003 0.0003 0.0007
2.0 0.0004 0.0004 0.0003 0.0007
2.5 0.0003 0.0003 0.0003 0.0007
3.0 0.0003 0.0003 0.0003 0.0008
Case 3a Strong correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0005 0.0007 0.0001 0.0002
1.3 0.0006 0.0007 0.0001 0.0004
1.4 0.0009 0.0010 0.0002 0.0005
1.5 0.0007 0.0007 0.0003 0.0007
1.7 0.0005 0.0006 0.0004 0.0008
1.9 0.0005 0.0006 0.0003 0.0009
2.0 0.0005 0.0005 0.0003 0.0009
2.5 0.0004 0.0004 0.0003 0.0009
3.0 0.0004 0.0004 0.0003 0.0009
Case 3b Strong correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0006 0.0007 0.0001 0.0002
1.3 0.0007 0.0008 0.0001 0.0004
1.4 0.0008 0.0009 0.0002 0.0005
1.5 0.0007 0.0008 0.0002 0.0006
1.7 0.0005 0.0006 0.0003 0.0007
1.9 0.0004 0.0004 0.0003 0.0008
2.0 0.0004 0.0005 0.0004 0.0009
2.5 0.0004 0.0004 0.0004 0.0009
3.0 0.0005 0.0005 0.0004 0.0010
Table S.5: Simulation of F-score for DP-based models
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0030 0.0025 0.0018 0.0040
1.2 0.0783 0.0772 0.0650 0.0810
1.3 0.2129 0.2152 0.1615 0.1844
1.4 0.3180 0.3204 0.2466 0.2722
1.5 0.3887 0.3924 0.3175 0.3444
1.7 0.4761 0.4791 0.4200 0.4446
1.9 0.5358 0.5375 0.4883 0.5115
2.0 0.5602 0.5634 0.5148 0.5380
2.5 0.6369 0.6390 0.6027 0.6213
3.0 0.6859 0.6879 0.6566 0.6738
Case 1b Weak correlation ZI 1.1 0.0023 0.0022 0.0022 0.0042
1.2 0.0748 0.0733 0.0622 0.0774
1.3 0.2109 0.2130 0.1531 0.1787
1.4 0.3212 0.3240 0.2415 0.2682
1.5 0.3863 0.3894 0.3138 0.3429
1.7 0.4793 0.4808 0.4187 0.4449
1.9 0.5378 0.5403 0.4886 0.5094
2.0 0.5602 0.5621 0.5118 0.5333
2.5 0.6382 0.6406 0.6011 0.6226
3.0 0.6837 0.6859 0.6556 0.6731
Case 2a Moderate correlation No ZI 1.1 0.0104 0.0102 0.0024 0.0050
1.2 0.3333 0.3429 0.0973 0.1231
1.3 0.4894 0.4973 0.2322 0.2616
1.4 0.5642 0.5684 0.3299 0.3578
1.5 0.6141 0.6186 0.3966 0.4269
1.7 0.6775 0.6796 0.4997 0.5287
1.9 0.7198 0.7212 0.5724 0.5988
2.0 0.7344 0.7351 0.5992 0.6242
2.5 0.7777 0.7789 0.6954 0.7141
3.0 0.8025 0.8034 0.7463 0.7622
Case 2b Moderate correlation ZI 1.1 0.0114 0.0122 0.0021 0.0049
1.2 0.3292 0.3375 0.0969 0.1204
1.3 0.4900 0.4964 0.2345 0.2626
1.4 0.5624 0.5677 0.3265 0.3548
1.5 0.6101 0.6136 0.3938 0.4237
1.7 0.6750 0.6766 0.4975 0.5248
1.9 0.7150 0.7166 0.5681 0.5941
2.0 0.7289 0.7303 0.5980 0.6226
2.5 0.7727 0.7735 0.6911 0.7102
3.0 0.7999 0.8007 0.7447 0.7614
Case 3a Strong correlation No ZI 1.1 0.0114 0.0131 0.0018 0.0039
1.2 0.4044 0.4167 0.0623 0.0852
1.3 0.6199 0.6280 0.1940 0.2290
1.4 0.7176 0.7231 0.3035 0.3437
1.5 0.7660 0.7701 0.3947 0.4325
1.7 0.8212 0.8234 0.5209 0.5537
1.9 0.8515 0.8527 0.5988 0.6270
2.0 0.8629 0.8638 0.6284 0.6571
2.5 0.8929 0.8934 0.7314 0.7531
3.0 0.9072 0.9076 0.7881 0.8055
Case 3b Strong correlation ZI 1.1 0.0145 0.0162 0.0020 0.0040
1.2 0.3987 0.4159 0.0639 0.0881
1.3 0.6192 0.6287 0.1955 0.2312
1.4 0.7147 0.7211 0.3062 0.3455
1.5 0.7658 0.7702 0.3964 0.4342
1.7 0.8231 0.8255 0.5197 0.5520
1.9 0.8537 0.8551 0.5966 0.6261
2.0 0.8633 0.8645 0.6251 0.6529
2.5 0.8931 0.8936 0.7308 0.7536
3.0 0.9076 0.9080 0.7897 0.8071

### v2

@@
 While clinical trials remain the gold standard for assessing medical product safety before approval, they have limitations due to factors such as small sample sizes, ethical considerations, restricted timelines, etc., and the critical need for postmarket monitoring is increasingly recognized.
-For example, severe AEs such as congenital malformations from Thalidomide and liver injury from the Ketoconazole (Lavertu et al.,, 2021) were identified only after these drugs were approved for the U.S.
+For example, severe AEs such as congenital malformations from Thalidomide and liver injury from the Ketoconazole lavertu2021new were identified only after these drugs were approved for the U.S.
 market.
@@
 For exposition, we focus only on drugs hereinafter, though analogous considerations apply to other medical products monitored in SRS databases.
-The SRS databases serve as key resources for assessing real-world evidence in postmarket drug safety by cataloging a large number of case safety reports linking drugs to relevant AEs.
+The SRS databases serve as key resources for assessing real-world evidence in postmarket drug safety by cataloging a large number of case safety reports, each listing the specific drugs taken by a user along with AEs encountered.
 For example, the FAERS database has compiled over $29$ million reports by 2024.
-These databases are inherently observational and suffer from selection bias, measurement error, confounding, and missing data, among other factors (Markatou and Ball,, 2014), precluding causal inference.
-Instead, the focus is on detecting signals: identifying drug-AE pairs that appear disproportionately more often than their null expected counts, derived under the assumption of independence between the drug and the AE using their marginal totals and the dataset’s grand total counts.
-Association analyses then estimate and assess the significance of signal strengths, quantifying the differences between the observed and (null) expected counts or related quantities defined by parameters of some underlying statistical model.
-Numerous SRS data mining and signal detection approaches have been developed over the past decades.
-Examples include the proportional reporting ratio (PRR) (Evans et al.,, 2001), the reporting odds ratio (ROR) (Rothman et al.,, 2004), formal frequentist methods such as the likelihood ratio test (LRT) and its variants (Huang et al., 2011; Chakraborty et al., 2022; Ding et al., 2020; and the references cited therein), and Bayesian methods such as the Gamma Poisson Shrinker (GPS) (DuMouchel,, 1999), Bayesian confidence propagation neural network (BCPNN) (Bate et al.,, 1998), and Dirichlet process (DP) Poisson Gamma mixture model (Hu et al.,, 2015).
+These databases are inherently observational, and the data curated in them suffer from several statistical challenges, including selection bias, measurement error, confounding, missing data, and absence of information on non-events, non-users, and general health characteristics of users, among other factors markatou2014pattern , precluding formal causal inference.
+In SRS data mining, the focus instead is on detecting signals— identifying drug-AE pairs whose observed counts are disproportionately higher than their null expected counts, the theoretical baseline counts for the pairs under the assumption that the drug and the AE are independent.
+Then signal detection entails first quantifying the signal strength, typically in terms of some parameterization of the ratio of the observed counts to the null expected counts (or some related quantities) for each drug-AE pair, and then identifying pairs that exhibit sufficiently large signal strengths.
+An important consideration in this context is the derivation of these null expected counts for drug-AE pairs.
+Unlike traditional contingency table-based case-control data analysis, a key challenge here is the absence of information on non-events: SRS databases record only reported drugs and AEs and provide no information on the total number of individuals taking the drugs who did not experience or report a specific AE.
+This absence means that the true denominators—the actual counts of patients exposed to each drug—are unavailable, precluding direct calculation of absolute risks or incidence rates as in traditional epidemiological designs.
+Instead, SRS data mining methods rely on statistical comparators (all other reports in the database not involving the drug or AE in question) to approximate marginal total report counts for each drug and AE.
+These approximated margins are then used to estimate the null expected counts for the drug-AE pairs gravel2023considerations ; kontsioti2023exploring ; zhang2018three ; markatou2014pattern .
+Despite these limitations, meaningful safety signals can be detected from these databases when appropriate statistical methods are applied, as demonstrated by a growing body of research.
+For principled and accurate signal detection from SRS databases, numerous statistical data mining methods have been developed over the past decades.
+Examples include the proportional reporting ratio (PRR) evans2001use , the reporting odds ratio (ROR) rothman2004reporting , formal frequentist methods such as the likelihood ratio test (LRT) and its variants huang2011likelihood ; chakraborty2022use ; ding2020evaluation , and Bayesian methods such as the Gamma Poisson Shrinker (GPS) dumouchel1999bayesian , Bayesian confidence propagation neural network (BCPNN) bate1998bayesian , and Dirichlet process (DP) Poisson mixture model Hu et al.hu2015signal .
 Some approaches—e.g., PRR and ROR—employ ad-hoc thresholds on empirical signal strength estimates, while the more formal methods—e.g., LRT and Bayesian approaches—employ rigorous probabilistic tools for statistically guaranteed signal detection, including controlled type I errors and false discovery rates (FDR).
 Despite these advances, signal detection for rare AEs and newer drugs remains challenging due to their small (marginal total) sample sizes, which leads to small observed and expected report counts in individual drugs, reducing the statistical power to detect the underlying signals.
-Yet, a large and growing body of research has documented the critical need for reliable pharmacovigilance signal detection precisely under these setups (see, e.g., Poluzzi et al., 2012; Sardella and Belcher, 2018; Raschi et al., 2021, and the references cited therein).
-Bayesian methods facilitate principled inference and signal detection for SRS data by modeling drug-AE signal strengths as model parameters with suitable priors and utilizing their posterior distributions.
-However, prior selection is crucial.
-An overly restrictive prior—e.g., a single gamma shrinkage prior for all drug-AE signal strength parameters—may cause excessive shrinkage, pulling estimates toward a single common value (Hu et al.,, 2015).
+Yet, a large and growing body of researchpoluzzi2012data ; sardella2018pharmacovigilance ; raschi2021value has documented the critical need for reliable early signal detection precisely under these settings involving rare AEs and newer drugs.
+In this paper, we argue that an appropriately formulated Bayesian framework can yield important insights for pharmacovigilance under these challenging settings.
+Bayesian SRS data mining methods enable principled inference and signal detection by modeling drug-AE signal strengths as parameters with suitable priors and utilizing their posterior distributions.
+Prior specification, however, is crucial for optimal performance.
+An overly restrictive prior—e.g., a single, common gamma prior for all drug-AE signal strength parameters—may cause excessive shrinkage, pulling estimates toward a single common valuehu2015signal .
 Conversely, an overly flexible prior—e.g., independent flat priors for all parameters—may inadequately regularize their estimation, resulting in overfitting and noisy estimates.
-State-of-the-art Bayesian data mining approaches balance flexibility and regularization using semi/nonparametric shrinkage priors—e.g., a two-component gamma mixture prior (the GPS model of DuMouchel, (1999)) or a DP mixture prior with a base gamma distribution (Hu et al.,, 2015)—for AE signal strengths separately for each drug.
-This stabilizes signal detection, provided the drugs are treated either all independently or as a single group.
-However, real-world SRS databases record drugs with nuanced similarities in their AEs.
-For instance, various statin drugs, used to lower blood cholesterol and cardiovascular risk, share chemical compositions and function, leading to common AEs like rhabdomyolysis (Montastruc,, 2023).
-On the other hand, small differences in drug composition and function can cause distinct AEs or variations in AE severity; e.g., myoglobinuria is typically associated with Atorvastatin but not Lovastatin (Seneff,, 2016).
-Figure 1 shows the pairwise Spearman correlation coefficients for occurrences of 46 common AEs of six statin drugs as cataloged in FAERS (Q3 2014-Q4 2024).
-These correlations reveal strong yet nuanced between-drug associations: they are consistently above 0.70, ruling out independence, yet many are noticeably below one, suggesting against a single grouping of all drugs.
-Figure 1: Spearman correlation matrix for occurrence counts of 46 common AEs (PTs) among statin drugs in the FDA FAERS data (2014 Q1-2020 Q4)
-Existing SRS data mining approaches, both Bayesian and frequentist, rarely account for nuanced between-drug associations.
-Instead, they often assume either independence of drugs (the original LRT, BCPNN, GPS, and DP Poisson-Gamma) or a single group for all drugs (extended LRT of Huang et al., (2011)).
+State-of-the-art Bayesian data mining approaches balance flexibility and regularization using semi/nonparametric shrinkage priors, such as a two-component gamma mixture prior (the GPS modeldumouchel1999bayesian ) or a DP mixture prior with a base gamma distributionhu2015signal , for AE signal strengths separately for each drug.
+These approaches stabilize signal detection; however, they still impose restrictive assumptions, treating drugs in question either entirely independently or as a single homogeneous group.
+Real-world SRS databases, however, record drugs with nuanced similarities in their AEs.
+For instance, various statin drugs, used to lower blood cholesterol and cardiovascular risk, share chemical compositions and function, leading to common AEs like rhabdomyolysis montastruc2023rhabdomyolysis .
+Yet, due to small differences in drug composition and function they may cause distinct AEs or have variations in AE severity; e.g., myoglobinuria is typically associated with Atorvastatin but not Lovastatin seneff2016statins .
+Figure 1 shows the pairwise Kendall’s $\tau$ correlation coefficients between six statin drugs with respect to the occurrence counts of 1491 AEs as cataloged in the FAERS database (Q3 2014-Q4 2024; see Section 5 for details on the dataset).
+These correlations reveal strong yet nuanced between-drug associations: they are consistently above $0.40$ with some as high as $0.68$, ruling out independence, yet many are noticeably around or below $0.50$, suggesting against a single grouping of all drugs.
+Figure 1: Kendall’s $\tau$ correlation matrix for occurrence counts of 1491 commonly occurring AEs (PTs) among six statin drugs in the FDA FAERS data (2014 Q1-2020 Q4).
+Existing SRS data mining approaches, both Bayesian and frequentist, rarely account for such nuanced between-drug associations.
+Instead, they often assume either independence of drugs (the original LRT, BCPNN, GPS, and DP Poisson-Gamma) or a single group for all drugs (extended LRT of Huang et al.huang2011likelihood ).
 These restrictive assumptions limit the statistical power and sensitivity for detecting signals, particularly for rare AEs and newer drugs with small observed and expected counts.
 Indeed, our extensive simulations (Section 5) show that existing methods suffer from poor sensitivity (after controlling FDR) when detecting signals from datasets containing multiple but weak true signals, rare AEs, and newer drugs.
-We address this critical gap with a novel local-global mixture DP-based Bayesian method that balances flexibility and shrinkage in AE signal strengths for each drug while enabling data-adaptive information sharing across drugs based on their associations, enhancing statistical power for signal discovery.
+We address this critical gap with a novel local-global mixture DP-based Bayesian method that balances flexibility and shrinkage in AE signal strengths for each drug while enabling data-adaptive information sharing across drugs with respect to their AE patterns, enhancing statistical power for signal discovery.
 In this framework, a local DP models distinct AE signal strength patterns specific to each drug while a global process defines a common pattern across drugs; the final AE signal strengths for a drug are obtained from a probabilistic mixture of these local and global processes.
@@
 Our key contributions are as follows.
-First, we propose local-global mixture DP models for SRS signal detection (Section 3).
-Local-global priors are widely used in contemporary statistics (Bhadra et al., 2019 and references cited therein) and have been employed in DP modeling of grouped data, notably via hierarchical DP models (Teh et al.,, 2004).
-Our framework instead adopts a more transparent two-component mixture of DPs (Antoniak,, 1974) for local-global pattern decomposition.
+First, we propose local-global mixture DP models for SRS signal detection (Section 4.2).
+Local-global priors are widely used in contemporary statistics (see Bhadra et al.bhadra2019lasso and references cited therein) and have been employed in DP modeling of grouped data, most prominently through hierarchical DP models teh2004sharing .
+Such models require handling complex combinatorial partition structures—often represented via Chinese restaurant franchise-style distributions teh2004sharing —making implementation challenging, particularly for sizable datasets.
+In contrast, our framework adopts a more transparent two-component mixture of DPs antoniak1974mixtures for local-global pattern decomposition.
+This enables our framework to achieve efficient MCMC implementation using standard distributions while retaining the ability to capture both shared global structure and drug-specific local variations.
 To our knowledge, the use of local-global DP models for principled SRS data mining is novel.
-We develop efficient MCMC algorithms for model implementation (Section 3.3 and Supplement S.3; software in the Supplement) and leverage FDR-controlled, FNR-optimized hypothesis tests for signal detection (Section 3.4).
-Second, we provide an intuitive overview of DP-based SRS data mining models to aid broader understanding (Section 3.1).
-Third, through extensive simulations (Section 4), we show that our methods achieve comparable or superior sensitivity to existing approaches (reviewed in Supplement S.1) when only a few strong signals are present across common AEs but yield substantial sensitivity gains—often twofold or more—when many weak signals exist among rare AEs and newer drugs, all while maintaining strong FDR and type I error control.
-Finally, our analysis of a real-world statin dataset (Section 5) from the FDA FAERS database demonstrates our methods’ ability to effectively detect AE signals in practice.
+We develop efficient MCMC algorithms for model implementation (Section 4.3 and Supplement S.2; software in the Supplement) and leverage FDR-controlled, FNR-optimized hypothesis tests for signal detection (Section 4.4).
+Second, we provide an intuitive overview of DP-based SRS data mining models and guidance on prior specification for hyperparameters to aid broader understanding (Section 4.1).
+Third, through extensive simulations (Section 5), we show that our methods achieve comparable or superior sensitivity to existing approaches (reviewed in Section 3) when only a few strong signals are present across common AEs but yield substantial sensitivity gains—often twofold or more—when many weak signals exist among rare AEs and newer drugs, all while maintaining strong FDR and type I error control.
+Finally, our analysis of a real-world statin dataset (Section 6) from the FDA FAERS database demonstrates our methods’ ability to effectively detect AE signals in practice from large-scale SRS databases.
 2 SRS DATA, OUR NOTATION, AND PROBABILITY MODELS
-An SRS database stores a large number of individual case safety reports that can be summarized as a high-dimensional contingency table cataloging the occurrences of multiple AEs across the rows and various drugs across the columns.
-Let $I$ and $J$ denote the total number of AEs and drugs under consideration, respectively.
+An SRS database stores a large number of individual case safety reports, each documenting the list of drugs a user was exposed to and the AEs encountered.
+These reports can be summarized as a high-dimensional contingency table cataloging the occurrences of all AEs noted in the database across the rows and various drugs across the columns.
 Further, let $n_{ij}$ be the total number/count of reported cases for drug-$j$ and AE-$i$; $i=1,\dots,I$ and $j=1,\dots,J$ constituting the contingency table $((n_{ij}))$.
 Define $n_{i\bullet}=\sum_{j=1}^{J}n_{ij}$, $n_{\bullet j}=\sum_{i=1}^{I}n_{ij}$, and $n_{\bullet\bullet}=\sum_{i=1}^{I}\sum_{j=1}^{J}n_{ij}$.
-Signal detection in SRS contingency tables often relies on the observed relative reporting rates, $n_{ij}/E_{ij}$, or a related quantity, as a measure of signal strength.
-Here, $E_{ij}=n_{i\bullet}n_{\bullet j}/n_{\bullet\bullet}$ represents the null expected count of drug-$j$ and AE-$i$, assuming independence between the drug and AE.
+As noted in the Introduction, the true total counts of users of a drug or the total number of individuals suffering from a specific AE are not available in SRS data due to the lack of information on non-events and non-users.
+Instead, SRS data mining methods consider comparator drugs—all drugs other than the drug-$j$ in question—and comparator AEs — all AEs other than the AE $i$ in question—cataloged in the SRS database, and compare the observed prevalence of the drug-$j$, AE-$i$ pair, which is substantially larger than other drug-AE pairs.
+Using the comparator drugs, and the reported margins $n_{i\bullet}$, $n_{\bullet j}$, and $n_{\bullet\bullet}$ measuring the total number of reported cases specific to AE-$i$, drug-$j$, and all drugs and AEs combined, respectively, they approximate the expected null baseline count $E_{ij}=n_{i\bullet}n_{\bullet j}/n_{\bullet\bullet}$, assuming independence between the drug and the AE.
+Of note, SRS data mining often focuses on identifying significant AEs of certain groups or subgroups of drugs, constituting only a (small, somewhat homogeneous) subset of all drugs cataloged in the database.
+Here, an artificial, heterogeneous group of drugs, obtained by collapsing all other drugs in the database, is often created as a ‘baseline’ or ‘reference’ category.
+This category of “Other drugs” is conceptually too heterogeneous to have any meaningful, unique AEs, but is still included to aid better comparison for disproportionality analysis, and a greater AE (row) total ($n_{i\bullet}$) counts for more accurate $E_{ij}$ estimation.
+Analogously, a heterogeneous group of ’Other AEs’, obtained by collapsing all AEs outside those of interest (or with very small sample sizes), may be constructed ding2020evaluation .
+Signal detection in SRS contingency tables typically relies on the observed relative reporting rates, $n_{ij}/E_{ij}$, or a related quantity, as a measure of signal strength.
 Intuitively, if $n_{ij}/E_{ij}\gg 1$, the observed frequency for $(i,j)$ is much higher than the null expected count, suggesting a potential signal.
@@
 When the Poisson assumption is violated, particularly due to an excess of zero counts arising from structural zeros—i.e., drug-AE pairs that are physically or biologically impossible to co-occur—a zero-inflated Poisson (ZIP) model is often preferred.
-In this framework, $n_{ij}$ is modeled as a mixture of two components: a $\text{Poisson}(\lambda_{ij}E_{ij})$ component for drug-AE pairs that can co-occur and a point mass at 00, denoted by $\delta_{\{0\}}$ (the Kronecker delta), for structural zeros.
+In this framework, $n_{ij}$ is modeled as a mixture of two components: a $\text{Poisson}(\lambda_{ij}E_{ij})$ component for drug-AE pairs that can co-occur and a point mass at 0, denoted by $\delta_{\{0\}}$ (the Kronecker delta), for structural zeros.
 Importantly, the structural zero positions are typically unknown and inferred from the SRS dataset.
@@
 The marginal distribution of $n_{ij}$, after integrating over the prior, can be more flexible than the likelihood model for $n_{ij}$.
-Indeed, with an appropriate nonparametric mixture prior, a flexible marginal distribution accommodating ZIP-like behavior can be achieved using only a Poisson likelihood for $n_{ij}$ (see Remark Remark under Section 3).
+Indeed, with an appropriate nonparametric mixture prior, a flexible marginal distribution accommodating ZIP-like behavior can be achieved using only a Poisson likelihood for $n_{ij}$ (see Section 4.1).
 Thus, Bayesian SRS data mining methods are less sensitive to the likelihood choice (Poisson or ZIP) when a flexible prior is used.
 However, if the zero-inflation probability parameters are of interest, the ZIP model remains useful.
-3 NOVEL LOCAL-GLOBAL DP MODELS FOR SRS DATA MINING
-We begin in Section 3.1 with a review of the drug-specific DP Poisson model of Hu et al., (2015), outlining its mechanism and providing guidance on the hyperparameters.
-We then introduce our novel local-global DP framework (Section 3.2) that accommodates between-drug associations and enables effective information sharing across drugs.
-3.1 A conceptual overview on the DP Poisson model
-Consider the Poisson likelihood $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$, with the DP prior (Hu et al.,, 2015):
-$\lambda_{ij}\sim D_{j}\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}};\ D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$ (1)
-for the signal strength parameters $\{\lambda_{1j}\dots,\lambda_{Ij}\}$ associated with each drug $j$.
-Here $\delta_{\{\theta\}}$ denotes a degenerate distribution at $\theta$ and DP$(\alpha,G(\beta))$ is a Dirichlet process with concentration hyperparameter $\alpha>0$ and base distribution $G(\beta)\equiv$ Gamma(shape = $\beta$, rate = $\beta$) with hyperparameter $\beta>0$.
-The discrete mixture structure ensures that $D_{j}$, the prior for $\lambda_{ij}$, is almost surely discrete/categorical placing probability weights $\eta_{hj}$; $10\leq\eta_{hj}\leq 1;\sum_{h=1}^{\infty}\eta_{hj}=$ 1 on atoms $\theta_{hj}$.
-This permits clustering, allowing multiple $\lambda_{ij}$ to share an atom $\theta_{hj}$ as a common value.
-The probability weights $\eta_{hj}$ are generated using stick-breaking (Sethuraman,, 1994):
-$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj})\text{ for }h\geq 2;\ v_{hj}\sim\text{independent Beta}(1,\alpha_{j})\text{ for }h\geq 1$ (2)
-based on the DP concentration hyperparameter $\alpha_{j}>0$.
-Meanwhile, the atoms are generated independently from the DP base distribution: $\theta_{hj}\sim G(\beta_{j})$.
-In practice, the infinite mixture defining $D_{j}$ can be well-approximated by a finite mixture $\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$ with a sufficiently large $K$ (Ishwaran and Zarepour,, 2002).
-This approximation simplifies model fitting and offers intuitive insights on the hyperparameters $\alpha_{j}$ and $\beta_{j}$, guiding their specification.
-Role of $\alpha_{j}$.
-The stick-breaking representation (2) elucidates that when $\alpha_{j}<1$, each $v_{hj}$ is, on average, greater than $1/2$ with smaller $\alpha_{j}$ favoring $v_{hj}$ closer to one.
-Consequently, $\eta_{hj}$ decreases sharply in $h$, leading the mixture (1), or its finite approximation with $K$ components, to retain only a few non-empty components with sufficiently large weights $\eta_{hj}$.
-Under a finite mixture approximation with $K_{+}<K$ non-empty components, (1) models $\lambda_{ij}$ using a categorical distribution with $K_{+}$ categories, placing point masses at the atoms $\theta_{1j},\dots,\theta_{K_{+}j}$ with weights $\eta_{1j},\dots,\eta_{K_{+}j}$.
+3 A REVIEW OF A FEW EXISTING SRS DATA MINING METHODS
+As noted in the Introduction, numerous SRS data mining methods have been proposed in the literature over the past decades.
+As representatives of the state-of-the-art and competitors of our proposed approach, we focus on four prominent and commonly used SRS data mining methods—both Bayesian and frequentist—that enable (upon appropriate adjustment, as needed) strong false discovery rate controls while permitting reasonable sensitivity for AE signal detection.
+This section reviews these methods, highlighting the underlying modeling assumptions.
+3.1 The Bayesian Propagation Neural Network (BCPNN) method
+The BCPNN approach (Bate et al.bate1998bayesian ) assumes that conditional on the grand total count $n_{\bullet\bullet}$, the cell-specific counts $n_{ij}$, the marginal row totals $n_{i\bullet}$, and the marginal column totals $n_{\bullet j}$ each distributed as binomial-beta: (a) $n_{ij}\mid p_{ij}\sim\text{Binomial}(n_{\bullet\bullet},p_{ij}),\ p_{ij}\sim\text{Beta}(\alpha_{ij},\beta_{ij})$, (b) $n_{i\bullet}\mid p_{i\bullet}\sim\text{Binomial}(n_{\bullet\bullet},p_{i\bullet}),\allowbreak\ p_{i\bullet}\sim\text{Beta}(\alpha_{i\bullet},\beta_{i\bullet})$, (c) $n_{\bullet j}\mid p_{\bullet j}\sim\text{Binomial}(n_{\bullet\bullet},p_{\bullet j}),\allowbreak\ p_{\bullet j}\sim\text{Beta}(\alpha_{\bullet j},\beta_{\bullet j})$, where $p_{ij}$ is the cell-specific reporting rate, and $p_{i\bullet}$ and $p_{\bullet j}$ are the marginal row and column reporting rates.
+We note that the binomial model assumptions for $n_{ij}$, $n_{i\bullet}$ and $n_{\bullet j}$ can be understood as consequences of an independent Poisson model for $n_{ij}$ with certain choices of $p_{ij}$, $p_{i\bullet}$, and $p_{\bullet j}$.
+Bate et al.bate1998bayesian consider the information content $\text{IC}_{ij}=\log_{2}\left[p_{ij}/(p_{i\bullet}p_{\bullet j})\right]$ as a measure of signal strength for the drug-$j$ AE-$i$ pair, with a higher IC indicating a stronger association.
+The authors suggest an empirical Bayes estimation of these quantities using $\hat{\alpha}_{ij}=\hat{\alpha}_{i\bullet}=\hat{\beta}_{i\bullet}=\hat{\alpha}_{\bullet j}=\hat{\beta}_{\bullet j}=1$ and $\hat{\beta}_{ij}=1/(E(p_{i\bullet}\mid n_{i\bullet})E(p_{\bullet j}\mid n_{\bullet j}))-1$, with the denominator of $\hat{\beta}_{ij}$ being computed using the posterior distributions $p_{i\bullet}\mid n_{i\bullet}\sim\text{Beta}(\hat{\alpha}_{i\bullet}+n_{i\bullet},\hat{\beta}_{i\bullet}+n_{\bullet\bullet}-n_{i\bullet})$ and $p_{\bullet j}\mid n_{\bullet j}\sim\text{Beta}(\hat{\alpha}_{\bullet j}+n_{\bullet j},\hat{\beta}_{\bullet j}+n_{\bullet\bullet}-n_{\bullet j})$.
+The consequent (empirical Bayes) posterior distribution of $\text{IC}_{ij}$ can be approximated using Monte Carlo draws or asymptotic normality.
+A signal is detected at the drug-$j$ and AE-$i$ pair if the 2.5th posterior percentile for $\text{IC}_{ij}$ exceeds 0.
+While this fixed threshold of 0 is not necessarily guaranteed to control false discoveries, methods to generate data-driven thresholds ensuring a predefined FDR level (e.g., $0.05$) have been proposed.ahmed2010false
+3.2 The Gamma Poisson Shrinker (GPS) model
+[…]
+$\lambda_{ij}\sim\kappa\text{Gamma}(\alpha_{1},\beta_{1})+(1-\kappa)\text{Gamma}(\alpha_{2},\beta_{2}),$
+where $\kappa\in[0,1]$ is the mixing weight hyperparameter, and $\alpha_{1},\beta_{1}>0$, and $\alpha_{2},\beta_{2}>0$ are mixture component-specific gamma hyperparemeters.
+Here Gamma($\alpha$, $\beta$) denotes a probability distribution with density $f_{\text{gamma}}(y\mid\alpha,\beta)=\frac{1}{\Gamma(\alpha)}\beta^{\alpha}y^{\alpha-1}e^{-\beta y}$, $y>0$; $\alpha,\beta>0$.
+Intuitively, the two components of the mixture are expected to represent clusters of $\lambda_{ij}$ corresponding to large values (“signals”) and small values (“non-signals”).
+The posterior distribution of $\lambda_{ij}$ given $n_{ij}$, $E_{ij}$, and the hyperparameters is:
+$\lambda_{ij}\mid n_{ij},E_{ij}\sim\kappa_{ij}\text{Gamma}(\hat{\alpha_{1}}+n_{ij},\hat{\beta_{1}}+E_{ij})+(1-\kappa_{ij})\text{Gamma}(\hat{\alpha_{2}}+n_{ij},\hat{\beta_{2}}+E_{ij})$
+where $\kappa_{ij}=\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})/[\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})+(1-\kappa)f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{2},\beta_{2})]$.
+Under an empirical Bayes approach, the point estimates $\hat{\kappa}$, $\hat{\alpha}_{1}$, $\hat{\beta}_{1}$, $\hat{\alpha}_{2}$, and $\hat{\beta}_{2}$ for the hyperparameters are obtained by maximizing their marginal likelihoods given $n_{ij}$ and $E_{ij}$.
+The resulting posterior of $\lambda_{ij}$ conditional on these estimated hyperparameters is used for inference and signal detection.
+Specifically, dumouchel2001empirical define the empirical Bayesian score $E(\log_{2}(\lambda_{ij})\mid n_{ij},E_{ij},\hat{\kappa},\hat{\alpha}_{1},\hat{\beta}_{1},\hat{\alpha}_{2},\hat{\beta}_{2})$ as a point estimate of signal strength for the drug-$j$ and AE-$i$ pair, and also suggest determining the pair to be a signal if the 5th percentile (EB05) of the above conditional posterior for $\lambda_{ij}$ exceeds a specific threshold, commonly 2.
+Similar to the BCPNN method, the deterministic threshold of $2$ may not necessarily ensure strong FDR control; however, a data-driven threshold controlling FDR to a pre-specified level (e.g., 0.05) may be obtained.ahmed2010false
+3.3 DP Poisson method of Hu et al.hu2015signal
+Several extensions and generalizations of the two-component mixture prior used in the GPS method have been suggested in the literaturehu2015signal ; zhang2018three ; tan2025flexible .
+Of these, we focus on one of the most flexible generalizations as proposed in Hu et al.hu2015signal , which employs a Dirichlet process mixture as the prior for $\lambda_{ij}$.
+Hu et al.hu2015signal consider the same Poisson likelihood $n_{ij}\sim$ Poisson($\lambda_{ij}E_{ij}$) as in GPS, but replace the two-gamma mixture prior for $\lambda_{ij}$ by a DP (infinite) mixture prior:
+$\lambda_{ij}\sim D_{j}\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}};\ D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j})).$ (1)
+[…]
+$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj})\text{ for }h\geq 2;\ v_{hj}\sim\text{independent Beta}(1,\alpha_{j})\text{ for }h\geq 1.$ (2)
+Meanwhile, the atoms $\theta_{hj}$ are generated independently from the DP base gamma distribution: $\theta_{hj}\sim G(\beta_{j})$; $h\geq 1$.
+In practice, the infinite mixture (1) defining $D_{j}$ can be well-approximated ishwaran2002exact by a finite mixture $\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$ with a sufficiently large number of components $K$; this finite mixture approximation provides a computationally tractable approach to implementing this model.
+Hu et al.hu2015signal consider full Bayesian inference on $\lambda_{ij}$ by first eliciting independent vague prior distributions on the hyperparameters $\alpha_{j}$ (specifically a Uniform(0.2, 10) prior) and $\beta_{j}$ (a Uniform(0, 1) prior) and then utilizing MCMC sampling from the resulting joint posterior.
+Subsequently, the drug-$j$ and AE-$i$ pair is identified as a signal if the 5th posterior percentile of $\lambda_{ij}$ computed using its posterior MCMC draws exceeds a certain threshold.
+The authors suggest using the deterministic threshold of $2$ for this purpose; a more data-driven threshold controlling FDR at a prespecified level (e.g., $0.05$) may be obtained.ahmed2010false
+Our proposed approach is motivated by this method, and we revisit this approach elucidating the key roles played by the model hyperparameters $\alpha_{j}$ and $\beta_{j}$, and guiding their prior specification for full Bayesian inference in Section 4.1.
+3.4 ZIP model based Pseudo LRT (frequentist) method
+In parallel to Bayesian developments, major advances on SRS data mining for AE signal detection have also been made using frequentist methods, notably using the machinery of the likelihood ratio test huang2011likelihood ; huang2013likelihood ; huang2014likelihood .
+These approaches often have the advantage of built-in FDR control while ensuring high sensitivity.
+We consider the ZIP model for the reporting count $n_{ij}\sim\text{ZIP}(n=n_{ij}\mid\lambda_{ij}\times E_{ij},\omega_{j})$.
+The pseudo LRT, as proposed by Chakraborty et al.chakraborty2022use , is a recent method designed to accommodate the zero-inflated nature of SRS data while easing computational demands.
+Similar to the standard LRT method, this approach first tests a global alternative hypothesis $H_{aj}:\lambda_{ij}>1$ for at least one $i\in\{1,\dots,I\}$ against the null hypothesis $H_{0j}:\lambda_{ij}=1$ for all $i\in\{1,\dots,I\}$ for a given drug-$j$.
+If $H_{aj}$ is rejected, a step-down procedure is used to conduct individual tests on $H_{aij}:\lambda_{ij}>1$.
+The ordinary LRT estimates the nuisance parameter $\omega_{j}$ separately for $H_{0j}$ and $H_{0j}\cup H_{aj}$, while the pseudo LRT is constructed using a global point estimate $\hat{\omega}_{j}$ of $\omega_{j}$ consistent under $H_{0j}\cup H_{aj}$ in both the numerator and the denominator of the likelihood ratio test statistic.
+Chakraborty et al consider the point estimate of $\omega_{j}$ is obtained by maximizing the profile likelihood $\ell_{j}(\omega_{j})$ for this purpose, with $\hat{\lambda}_{ij}=\max\{1,n/E_{ij}\}$ as the maximum likelihood estimate of $\lambda_{ij}$.
+The resulting statistic is given by
+$\text{LR}_{ij}=\exp\left[-(\hat{\lambda}_{ij}-1)E_{ij}\right]\hat{\lambda}_{ij}^{n_{ij}}.$
+The maximum likelihood ratio statistic for drug-$j$ is then defined as $\text{MLR}_{j}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\}\end{subarray}}\text{LR}_{ij}$.
+In the global test, a parametric bootstrap is used to approximate the p-value $P_{j}=\Pr(\text{MLR}_{j}\geq\text{MLR}_{j}^{\text{obs}})$ using (parametric bootstrap) draws for $\text{LR}_{ij}$ under $H_{0j}$ and hence $\text{MLR}_{j}$, where $\text{MLR}_{j}^{\text{obs}}$ is the observed test statistic.
+If the global test is rejected, step-down p-values chakraborty2022use ; huang2011likelihood are computed for the individual tests using the same parametric bootstrap samples for $\text{MLR}_{j}$ and using the individual observed likelihood ratio test statistic values, such that $P_{ij}=\Pr(\text{LR}_{ij}\geq\text{MLR}_{ij}^{\text{obs}})$.
+For multiple drugs, the same approach can be applied, defining an overall maximum likelihood ratio as $\text{MLR}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\},j\in\{1,\dots,J\}\end{subarray}}\text{LR}_{ij}$.
+This step-down procedure employing the same global test statistic strictly controls the false discovery rate at the same level $\alpha$ of the global test performed using the MLR huang2011likelihood ; chakraborty2022use .
+When multiple drugs $j$ are tested simultaneously, one considers the extended version with global maximum likelihood ratio $\max_{j\in\{1,\dots,J\}}\text{MLR}_{j}$.
+4 A NOVEL LOCAL-GLOBAL DP MODEL FOR SRS DATA MINING
+We begin this section by revisiting the DP Poisson model of Hu et al.
+hu2015signal (Section 4.1), highlighting the role of its hyperparameters in SRS signal discovery with the aid of a schematic plot (Figure 2a) and providing guidance on prior specification for these hyperparameters.
+We then introduce our novel local-global DP framework (Section 4.2), which accommodates between-drug associations and enables effective information sharing across drugs, thereby enhancing SRS data mining.
+4.1 The DP Poisson model revisited: the roles of the hyperparameters in SRS signal discovery and the effect of zero-inflation
+Role and prior specification for the hyperparameter $\alpha_{j}$
+The stick-breaking representation (2) elucidates that when $\alpha_{j}<1$, each $v_{hj}$, on average, is greater than $1/2$ with smaller values of $\alpha_{j}$ favoring $v_{hj}$ closer to one.
+Consequently, $\eta_{hj}$ decreases sharply in $h$, leading the mixture (1), or its finite approximation with $K$ components, to retain only a few non-empty components with sufficiently large and positive weights $\eta_{hj}$.
+Under a finite mixture approximation with $K_{+}<K$ non-empty components, (1) models $\lambda_{1j},\dots,\lambda_{Ij}$ using a categorical distribution with $K_{+}$ categories, placing point masses at the atoms $\theta_{1j},\dots,\theta_{K_{+}j}$ with weights $\eta_{1j},\dots,\eta_{K_{+}j}$.
 This fosters clustering, restricting $\lambda_{1j},\dots,\lambda_{Ij}$ to at most $K_{+}$ distinct values.
 When $K_{+}<I$, multiple $\lambda_{ij}$ share the same cluster $h$ admitting the common value $\theta_{hj}$, facilitating their pooling for information sharing.
-While a smaller $\alpha$ encourages a smaller $K_{+}$ aiding greater information sharing, a larger $K_{+}$ ($>2$) is permitted in the posterior when supported by data.
-This contrasts with the GPS framework of DuMouchel, (1999) commonly used for SRS signal detection (see Section S.1 for a review) which strictly categorizes $\lambda_{ij}$ into two groups, intuitively, signal ($\lambda_{ij}\leq 1$) and non-signal ($\lambda_{ij}\gg 1$).
+It follows that a smaller $\alpha$ encourages a smaller $K_{+}$, resulting in fewer clusters and thus greater information sharing among the $\lambda_{ij}$, while a larger $\alpha$ induces a larger $K_{+}$, resulting in more clusters that collectively capture more heterogeneity among the $\lambda_{i}j$ values.
+Importantly, when supported by the data and induced through a (moderately) large $\alpha$, values of $K_{+}>2$ are permitted in the posterior of this method.
+This contrasts with the GPS method of dumouchel1999bayesian (see Section 3 for a review) which strictly categorizes $\lambda_{ij}$ into two groups, intuitively, signal ($\lambda_{ij}\leq 1$) and non-signal ($\lambda_{ij}\gg 1$).
 The DP framework, by allowing $K_{+}>2$, enhances heterogeneity acknowledgment within both signal and non-signal groups.
@@
 Because a smaller $\alpha_{j}$ promotes greater information sharing, a prior assigning sufficient mass to small positive values is recommended for stable estimation of $\lambda_{ij}$.
-Hu et al., (2015) used a Uniform(0.2, 10) prior for $\alpha_{j}$.
-We suggest a more aggressive weakly informative default prior, viz., Exponential(rate = $\psi_{\alpha}$) with a moderately large $\psi_{\alpha}$, e.g., $\psi_{\alpha}=3$ (placing $\approx 95\%$ mass in $(0,1)$), which favors small $\alpha_{j}$ while allowing larger values if needed.
+Hu et al.hu2015signal used a Uniform(0.2, 10) prior for $\alpha_{j}$.
+We suggest a more aggressive weakly informative default prior, viz., Exponential(rate = $\psi_{\alpha}$) with a moderately large $\psi_{\alpha}$, e.g., $\psi_{\alpha}=3$ (placing $\approx 95\%$ mass in $(0,1)$).
+This prior favors small $\alpha_{j}$ while allowing larger values if needed.
 This choice also ensures conjugacy for efficient Gibbs sampling updates for MCMC implementation.
@@
 Probabilistic two-component mixtures of $\lambda_{ij}^{l}$ and $\lambda_{i}$ (Panel B.3) based on the local distribution indicators $z_{ij}^{l}$ produce the final realized signal strengths $\lambda_{ij}$ (Panel B.4).
-Role of $\beta_{j}$.
+Role and prior specification for the hyperparameter $\beta_{j}$
 The parameter $\beta_{j}$ controls variability among the atoms $\theta_{hj}$ drawn from the base $G(\beta_{j})\equiv$ Gamma(shape = $\beta_{j}$, rate = $\beta_{j}$) distribution, thereby controlling the shrinkage in $\lambda_{ij}$.
@@
 […]
-Remark.
-[…]
+Effect of zero-inflation
+Zero inflation can be handled to some extent in the DP Poisson model through a small atom $\theta_{hj}\approx 0$, causing $\lambda_{ij}$ and hence Poisson($\lambda_{ij}E_{ij}$) to concentrate near zero for all AEs $i$ in cluster $h$, effectively mimicking a degenerate-at-zero $\delta_{\{0\}}$ distribution.
+This yields a model similar to $n_{ij}\sim$ ZIP($\lambda_{ij}E_{ij},\omega_{j})$, $\omega_{j}$ being the structural zero probability, with an analogous DP prior on $\lambda_{ij}$, producing similar posterior for $\lambda_{ij}$.
+Consequently, signal detection and inference on $\lambda_{ij}$ are also expected to be broadly similar in both models (Section 5 provides numerical validation).
+However, some (small) differences may arise in practice, particularly when analyzing large real-world SRS datasets with complex signal patterns, due to numerical rounding errors for handling zero atoms.
+Furthermore, for direct inference on the structural zero positions, the Poisson approximation of $\delta_{\{0\}}$ is insufficient, and a ZIP model $n_{ij}\sim\text{ZIP}(\lambda_{ij}E_{ij},\omega_{j})$ is needed.
+The same prior (3) may still be used on $\lambda_{ij}$ assuming $\lambda_{ij}=0$ whenever the underlying cell-specific zero-inflation indicator, say $y_{ij}$, indicates a structural zero to ensure identifiability.
+For any drug-$j$, the posterior distributions of $y_{ij}$ and/or the column-specific zero-inflation probability $\omega_{j}$ provide inference on whether AE-$i$ and/or any AE is a structural zero, respectively.
+In practice, it is thus important to assess whether structural zeros/zero inflation are likely in the SRS dataset being analyzed, and whether inference on them is of interest, in order to guide appropriate model selection.
+The DP-Poisson framework offers some robustness against model misspecification due to zero inflation, but explicit ZIP modeling may be preferred when structural zeros are a primary concern.
 4.2 The local-global DP Poisson model
-The DP model in Section 3.1 enables flexible and effective signal detection for each drug separately but does not account for associations between drugs in their signal strenghts.
+The DP model in Section 4.1 enables flexible and effective signal detection for each drug separately but does not account for associations between drugs in their signal strengths.
 To capture these associations, we introduce the concept of local and global (prior) distributions for AE signal strengths.
@@
 $\lambda_{ij}=z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g};\ \lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D;\ z_{ij}^{l}\sim\text{Bernoulli}(\pi^{l})$
-$D_{j}\sim DP(\alpha_{j},G(\beta_{j}));\ D\sim DP(\alpha,G(\beta)).$ (3)
+$D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}));\ D\sim\text{DP}(\alpha,G(\beta)).$ (3)
 Here, the realized signal strengths $\{\lambda_{1j},\dots,\lambda_{Ij}\}$ for a drug $j$ is generated by combining local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$ drawn from a drug-$j$-specific local distribution $D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$ as in (1) and global strengths $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ generated from a global distribution $D\sim\text{DP}(\alpha,G(\beta))$ shared by all drugs.
@@
 Figure 2B illustrates this mechanism using $I=6$ AEs and $J=3$ drugs.
-This bi-level clustering contrasts with the single-level (AE-only) clustering induced by the DP prior of Hu et al., (2015).
+This bi-level clustering contrasts with the single-level (AE-only) clustering induced by the DP prior of Hu et al.hu2015signal .
 As demonstrated in extensive simulations (Section 5), the bi-level clustering and information sharing play a critical role in stabilizing the posterior and enhancing signal detection for $\lambda_{ij}$, particularly for drug-AE pairs where $n_{i\bullet}$ is small (rare AEs) and/or $n_{\bullet j}$ is small (newer drugs), leading to small $E_{ij}$.
@@
 Additionally, SRS data often include a collapsed, heterogeneous comparator category $J$ (e.g., “Other drugs”), serving as a baseline group with few, if any, distinct AE signals.
-In such cases, shrinking all $\lambda_{iJ}$ toward $1$ using a prior such as $\text{Gamma}(\text{shape}=\tau,\text{rate}=\tau)$ with an appropriate hyperparameter $\tau$ can improve signal detection and inference (see Supplement S.2.1).
-As noted in Remark Remark, a separate ZIP likelihood $n_{ij}\sim\text{ZIP}(\lambda_{ij}E_{ij},\omega_{j})$ is unnecessary for signal detection due to the flexibility of the DP prior.
-However, it may be used if direct inference on structural zero positions/probabilities is required.
-Complete hierarchical Bayesian representations of our Poisson and ZIP-based models are provided in Supplement S.2.1.
+In such cases, shrinking all $\lambda_{iJ}$ toward $1$ using a prior such as $\text{Gamma}(\text{shape}=\tau,\text{rate}=\tau)$ with an appropriate hyperparameter $\tau$ can improve signal detection and inference (see Supplement S.1.1).
+As noted in Section 4.1, while the DP-Poisson framework is expected to provide some robustness against zero-inflation, a separate ZIP likelihood $n_{ij}\sim\text{ZIP}(\lambda_{ij}E_{ij},\omega_{j})$ may be used if substantial zero inflation is present in the data, or direct inference on structural zero positions/probabilities is required.
+Complete hierarchical Bayesian representations of our Poisson and ZIP-based models are provided in Supplement S.1.1.
 Remark.
-The Hu et al., (2015) model can be viewed as a “local-only” special case of our local-global model, with only drug-specific local distributions $D_{j}$ for $\lambda_{ij}$ and no global distribution (i.e., $\pi^{l}=1$).
+The Hu et al.hu2015signal model can be viewed as a “local-only” special case of our local-global model, with only drug-specific local distributions $D_{j}$ for $\lambda_{ij}$ and no global distribution (i.e., $\pi^{l}=1$).
 At the other extreme, a “global-only” model ($\pi^{l}=0$) yields common $\lambda_{i}$ across drugs.
-This is a restrictive model that cannot differentiate between drugs and only permits “global” signal detection—assessing whether an AE-$i$ is a signal for any drug under consideration (see Supplement S.2.2 for further discussion on this model).
+This is a restrictive model that cannot differentiate between drugs and only permits “global” signal detection—assessing whether an AE-$i$ is a signal for any drug under consideration (see Supplement S.1.2 for further discussion on this model).
 Remark.
 Extensive literature exists on local-global DP models for sharing information across groups while preserving DP-based clustering within each group.
-A common approach employs hierarchical DPs (Teh et al.,, 2004), where local DPs share a global DP as a base distribution.
-In contrast, our approach leverages a two-component DP mixture framework (Antoniak,, 1974), separately modeling local and global AE signal strengths and thus offering a more transparent bi-level clustering over AEs and drugs, enabling more flexible estimation of the local and global parameters for greater flexibility in SRS applications.
+A common approach employs hierarchical DPs teh2004sharing , where local DPs share a global DP as a base distribution.
+In contrast, our approach leverages a two-component DP mixture framework antoniak1974mixtures , separately modeling local and global AE signal strengths and thus offering a more transparent bi-level clustering over AEs and drugs, enabling more flexible estimation of the local and global parameters for greater flexibility in SRS applications.
 Remark.
 In this paper, we focus on drug-based pharmacovigilance, identifying AE (row) signals given one or more drugs (columns in SRS contingency tables).
-An alternative AE-based approach discovers drugs strongly associated with a given set of AEs (Ding et al.,, 2020; Liu et al.,, 2024).
-A row-based version of our column-based prior (3), obtained by switching the roles of $i$ and $j$, can be used for such analyses.
+An alternative AE-based approach can be performed to discover drugs strongly associated with a given set of AEs ding2020evaluation ; liu2024pattern .
+Our proposed approach can be readily adapted for such analyses: a row-based version of our column-based prior (3), obtained by switching the roles of $i$ and $j$, can be used for this purpose.
 4.3 Full Bayesian implementation using MCMC
 For full Bayesian inference on our models, we place independent weakly informative proper priors on all model hyperparameters.
-Along with the exponential prior for $\alpha_{j}$ and $\alpha$, and half-Cauchy prior for $\beta_{j}$ and $\beta$ as suggested in Section 3.1, we consider a similar Cauchy+(0, $\psi_{\tau}$) prior for $\tau$ with $\psi_{\tau}=1/2$, and independent Uniform$(0,1)$ priors for the probability weight parameters, viz., $\pi_{l}$ and $\omega_{j}$ (if included in the model).
-We develop efficient MCMC algorithms for rigorous posterior sampling, approximating the DPs using finite mixtures with $K=O(\log(IJ))$ components.
+Along with the exponential prior for $\alpha_{j}$ and $\alpha$, and half-Cauchy prior for $\beta_{j}$ and $\beta$ as suggested in Section 4.1, we consider a similar Cauchy+(0, $\psi_{\tau}$) prior for $\tau$ with $\psi_{\tau}=1/2$, and independent Uniform(0,1)(0,1) priors for the probability weight parameters, viz., $\pi_{l}$ and $\omega_{j}$ (if included in the model).
+We develop efficient MCMC algorithms for rigorous posterior sampling, approximating the DPs using finite mixtures with $K=O(\log(IJ))$ components ishwaran2001gibbs .
 Each MCMC iteration involves data augmentation to generate latent DP mixture allocations, followed by Gibbs sampling updates for the stick-breaking weights and atoms and the concentration hyperparameters for each local and global DP.
 The local distribution indicators $z_{ij}^{l}$, local distribution proportion $\pi^{l}$, AE signal strengths $\lambda_{iJ}$ for the reference drug $J$, and zero-inflation indicators $y_{ij}$ and probabilities $\omega_{j}$ (if included in the model) also admit standard Gibbs updates.
-However, the full conditionals for the DP base Gamma hyperparameters ($\beta_{j}$ and $\beta$) and the reference category shrinkage hyperparameter $\tau$ lack standard forms; for these we employ stepping-out slice sampling (Neal,, 2003).
+However, the full conditionals for the DP base Gamma hyperparameters ($\beta_{j}$ and $\beta$) and the reference category shrinkage hyperparameter $\tau$ lack standard forms; for these we employ univariate stepping-out slice sampling neal2003slice steps.
 Detailed MCMC steps are provided in Supplementary Algorithms S.1, S.2 and S.3.
@@
 $H_{0,ij}:\lambda_{ij}\leq 1\text{ vs. }H_{1,ij}:\lambda_{ij}>1$
-for all AEs $i=1,\dots,I$ and all drugs $j=1,\dots,J$.
-The null hypothesis $H_{0,ij}$ is rejected in favor of $H_{1,ij}$, identifying the drug-$j$ and AE-$i$ pair to be a signal, if $T_{ij}>k$, where $T_{ij}$ is a test statistic derived from the posterior distribution and $k>1$ is an appropriate threshold.
-A common choice for $T_{ij}$ is the marginal posterior posterior quantile $T_{ij}(p)$, defined by $\Pr{\!}_{\text{posterior}}(\lambda_{ij}\leq T_{ij}(p))=p$ for a small pre-specified quantile level $p\in(0,1)$, where $\Pr{\!}_{\text{posterior}}(\cdot)$ denotes posterior probability given the data, and can be computed via MCMC draws.
-Hu et al., (2015) suggest using the 5-th posterior quantiles $T_{ij}(0.05)$ as the test statistic with a rejection threshold of $k=2$; however, such deterministic thresholds can be overly conservative (curtailing sensitivity) or overly liberal (failing to control false discoveries).
-Instead, we consider a family of tests $\{T_{ij}(p)>k\}$ with varying $p$ (reasonably small) and $k>1$, selecting the optimal $(\hat{p},\hat{k})$ that controls the false discovery rate (FDR) while minimizing the false negative rate (FNR).
-Given $p$ and $k$, FDR and FNR are estimated using the posterior distribution as (Ahmed et al., 2010; Müller et al., 2006; in the expressions below $1(A)$ denotes the indicator of a set $A$, $\hat{=}$ denotes “estimated by”, and $T^{\text{obs}}$ is the computed value of $T$):
+[…]
 $\text{FDR}(p,k)\ \hat{=}\ \frac{\sum_{i,j}q_{ij}\ I(T_{ij}(p)^{\text{obs}}>k)}{\sum_{ij}1(T_{ij}(p)^{\text{obs}}>k)},\quad\text{FNR}(p,k)\ \hat{=}\ \frac{\sum_{i,j}(1-q_{ij})\ 1(T_{ij}(p)^{\text{obs}}\leq k)}{\sum_{ij}1(T_{ij}(p)^{\text{obs}}\leq k)},$ (4)
-where $q_{ij}=\Pr{\!}_{\text{posterior}}(\lambda_{ij}\leq 1)$ is the posterior probability of $H_{0,ij}$, computable from MCMC draws.
-A grid search is used to determine the optimal $(\hat{p},\hat{k})$ that ensures $\text{FDR}(\hat{p},\hat{k})\leq\alpha$ for a pre-specified level $\alpha$ while minimizing FNR (see Supplement S.3.3 for details on the grids).
+[…]
 In practice, drug-AE pairs with $n_{ij}=1$ are often deemed non-signals, even if $\lambda_{ij}\gg 1$ with high posterior probability, leading to $T_{ij}(p)>k$ for suitably chosen $p$ and $k$.
@@
 In this section, we conduct extensive simulations to assess the signal detection performance of our local-global DP Poisson and ZIP models under a wide array of data-generating settings.
-To generate random SRS contingency tables of dimension $I\times J$, we first define a prespecified true signal strength matrix (described below) $\Lambda^{0}=((\lambda_{ij}^{0}))$ encoding signals, structural zeros, non-signals, and between-drug dependency patterns.
+We follow the SRS contingency table generation approach previously used in the literature chakraborty2022use ; liu2024pattern .
+Specifically, to generate random SRS contingency tables of dimension $I\times J$, we first define a prespecified true signal strength matrix (described below) $\Lambda^{0}=((\lambda_{ij}^{0}))$ encoding signals, structural zeros, non-signals, and between-drug dependency patterns.
 Using a reference $I\times J$ contingency table (described next) from a real-world SRS database, we extract row and column totals $\{\bm{n}_{i\bullet}:i=1,\dots,I\}$ and $\{\bm{n}_{\bullet j}:j=1,\dots,J\}$, respectively, to facilitate random contingency table generation.
@@
 Finally, we generate a random SRS contingency table by sampling from $\text{Multinomial}(\bm{n}_{\bullet\bullet};\bm{p})$, collapsing any generated $n_{ij}=1$ corresponding to a true signal (as defined by $\lambda^{0}_{ij}$) to $n_{ij}=0$ afterward to eliminate signal detection at $n_{ij}=1$.
-Reference SRS dataset to compute row and column totals.
-We use the FDA FAERS database (Q3 2014-Q4 2020), which catalogs a total of 63,976,610 reported cases (Chakraborty et al.,, 2023), to construct our reference SRS contingency table.
+Reference SRS dataset to compute row and column totals
+We use the FDA FAERS database (Q3 2014-Q4 2020), which catalogs a total of 63,976,61063{,}976{,}610 reported cases chakraborty2023likelihood , to construct our reference SRS contingency table.
 We focus on six statin medications: Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, collapsing all other drugs into a comparator “Other drugs” category.
-Using 46 preferred terms (PTs) for AEs, we form a $47\times 7$ contingency table summarizing the report counts for each drug-AE pair.
-This dataset has been previously analyzed in Chakraborty et al., (2022).
-Constructing the true signal strength matrix $\Lambda^{0}$.
+For expository purposes, and to create a sufficiently large SRS contingency table, we focused on 1491 preferred terms (PTs) as AEs that are common (appear at least once) for these statin drugs as well for a separate group of drugs unrelated to statin, namely, Gadolinium-based contrast agents (GBCA) drugs chakraborty2022use .
+All other PTs recorded in the database within this period were collapsed into a heterogeneous category labeled “Other AE”, yielding a $1492\times 7$ contingency table summarizing the report counts for each drug-AE pair.
+Constructing the true signal strength matrix $\Lambda^{0}$
 To generate $\Lambda^{0}$, we define two signal patterns: (1) fixed row signals, where all drugs (columns) exhibit signals for the same AEs (PTs; rows), and (2) random signals, where signals appear on randomly chosen AEs (different row for each drug).
@@
 The remaining cells $(i^{\prime},j^{\prime})$ are designated as non-signals with $\lambda_{i^{\prime}j^{\prime}}^{0}=1$.
-To introduce zero inflation, we prespecify true $\omega_{j}^{0}$ (00 or $0.10$; Table 1), and set $\omega_{j}^{0}I$ randomly selected non-signal cells to $\lambda_{ij}^{0}=0$.
+To introduce zero inflation, we prespecify true $\omega_{j}^{0}$ (0 or $0.30$; Table 1), and set $\omega_{j}^{0}I$ randomly selected non-signal cells to $\lambda_{ij}^{0}=0$.
 This approach yields a controlled $\Lambda^{0}$ to generate random SRS contingency tables.
@@
 The comparator “Other drugs” column has all non-signal $(\lambda_{ij}^{0}=1)$ cells.
-Finally, the fourth row in the table shows the average Pearson’s correlation coefficient $\rho$ between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ over all statin drug pairs $j$ and $j^{\prime}$ (excluding the “Other drugs” category) as a measure of pairwise dependence between drugs.
+Finally, the fourth row in the table shows the average Kendall’s $\tau$ between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ across all statin drug pairs $j$ and $j^{\prime}$ (excluding the “Other drugs” category) as a measure of pairwise association of their AE signal strengths.
 Simulation Scenario 1a 1b 2a 2b 3a 3b
-No. of fixed row signals 1 1 3 3 6 6
-No. of random signals per column 5 5 3 3 1 1
-Zero-inflation rate per column 0 0.1 0 0.1 0 0.1
-Average $\rho$ between columns 0.04 0.04 0.43 0.43 0.83 0.83
-  • The common/fixed PTs (fixed AEs/rows) that are considered signal in each case are as follows:
-    • Case 1: Rhabdomyolysis
-    • Case 2: Muscle Disorder, Myalgia, Rhabdomyolysis
-    • Case 3: Blood Creatine Phosphokinase Increased, Muscle Disorder, Muscle Rupture, Musculoskeletal Discomfort, Myalgia, Rhabdomyolysis
+No. of fixed row signals 5 5 10 10 20 20
+No. of random signals per column 17 17 10 10 3 3
+Zero-inflation rate per column 0 0.3 0 0.3 0 0.3
+Average Kendall’s $\tau$ between columns 0.12 0.12 0.49 0.49 0.77 0.77
   • A common signal strength value $\lambda^{0,\text{signal}}$ is assigned to all signal cells in a given simulation setting, and we vary $\lambda^{0,\text{signal}}\in\{1.1,1.2,1.3,1.4,1.5,1.7,1.9,2,2.5,3\}$.
-Replicated Data Generation, Model Fitting, and signal detection.
-For each constructed signal strength matrix $\Lambda$ (Table 1), and using the random contingency table generation mechanism described earlier, we generate $R=1{,}000$ independent random contingency table replicates.
-In each replicate, we fit our models via MCMC sampling (Supplementary Algorithms S.1, S.2 and S.3) with $10{,}000$ post-burn-in draws after discarding the first $5,000$ burn-in draws; see Supplement S.3 for additional computational details, including starting points.
+Replicated Data Generation, Model Fitting, and signal detection
+For each constructed signal strength matrix $\Lambda$ (Table 1), and using the random contingency table generation mechanism described earlier, we generate R=1, [^0] R=1{,}000 independent random contingency table replicates.
+In each replicate, we fit our models via MCMC sampling (Supplementary Algorithms S.1, S.2 and S.3) with 10,00010{,}000 post-burn-in draws after discarding the first $5,000$ burn-in draws; see Supplement S.2 for additional computational details, including starting points for the MCMC algorithms.
 From the computed posteriors, we conduct signal detection using hypothesis testing (Section 4.4).
-We compare our signal detection performance against existing Bayesian SRS data mining methods, viz., BCPNN, GPS, and the local-only DP model of Hu et al., (2015) (henceforth DP Hu et al).
+We compare our signal detection performance against existing Bayesian SRS data mining methods, viz., BCPNN, GPS, and the local-only DP model of Hu et al.hu2015signal (henceforth DP Hu et al).
 For our models and DP Hu et al., signal detection uses an optimized posterior quantile level $\hat{p}$ (fixed at 0.05 for DP Hu et al.) and threshold $\hat{k}$ ensuring FDR $\leq 0.05$ while minimizing FNR (Section 4.4), with the additional constraint $n_{ij}>1$ for our methods.
-To ensure fair FDR control ($\leq 0.05$) in BCPNN and GPS, we leverage the independence of signal strengths in the (empirical) Bayes posteriors for these two models and compute Benjamini-Hochberg-adjusted (Müller et al.,, 2006) posterior probabilities of being a non-signal; a drug-$j$ and AE-$i$ pair is then classified as a signal if the adjusted probability is $\leq 0.05$.
-Additionally, we evaluate the state-of-the-art frequentist ZIP-based extended (pseudo) LRT (Chakraborty et al.,, 2022; Huang et al.,, 2011; Ding et al.,, 2020), which ensures FDR control ($\leq 0.05$) with high sensitivity while addressing zero inflation in the counts when present (see Supplement S.1 for a detailed review of competing approaches).
+To ensure fair FDR control ($\leq 0.05$) in BCPNN and GPS, we leverage the independence of signal strengths in the (empirical) Bayes posteriors for these two models and compute Benjamini-Hochberg-adjusted muller2006fdr posterior probabilities of being a non-signal; a drug-$j$ and AE-$i$ pair is then classified as a signal if the adjusted probability is $\leq 0.05$.
+Additionally, we evaluate the state-of-the-art frequentist ZIP-based extended (pseudo) LRT chakraborty2022use ; huang2011likelihood ; ding2020evaluation , which ensures FDR control ($\leq 0.05$) with high sensitivity while addressing zero inflation in the counts when present.
 We assess signal detection for all methods using replication-based FDR, sensitivity, average Type I error, and F-score (Figure 3).
@@
 Each column presents one metric and each row presents one simulation scenario in Table 1.
-The first two rows of Figure 3 show comparable (when $\lambda^{0,\text{signal}}<1.5$) or slightly better (when $\lambda^{0,\text{signal}}\geq 1.5$) sensitivity for our local-global methods compared to all other methods when the between-drug association is near zero (Cases 1a,b).
-Under moderate between-drug association ($\rho=0.43$; Cases 2a,b), our methods perform noticeably better in detecting more true signals, achieving up to twice the sensitivity, particularly for weak or moderate true signal strengths ($1.1\leq\lambda^{0,\text{signal}}\leq 1.7$).
-The gain is even further pronounced under high between-drug association ($\rho=0.83$; Cases 3a,b), where our methods attain sensitivities of $\geq 0.70$ with only moderate true signal strength ($\lambda^{0,\text{signal}}\geq 1.7$), while all competitors fail to reach $0.70$ sensitivity even for $\lambda^{0,\text{signal}}=3$.
+The first two rows of Figure 3 show comparable (when $\lambda^{0,\text{signal}}<1.5$) or slightly better (when $\lambda^{0,\text{signal}}\geq 1.5$) sensitivity for our local-global methods compared to all other methods when the between-drug association is very small ($\tau=0.12$; Cases 1a,b).
+Under moderate between-drug association ($\tau=0.49$; Cases 2a,b), our methods perform noticeably better in detecting more true signals, achieving up to twice the sensitivity, particularly for weak or moderate true signal strengths ($1.1\leq\lambda^{0,\text{signal}}\leq 1.7$).
+The gain is even further pronounced under high between-drug association ($\tau=0.77$; Cases 3a,b), where our methods attain sensitivities of $\geq 0.70$ with only moderate true signal strength ($\lambda^{0,\text{signal}}\geq 1.7$), while all competitors fail to reach $0.70$ sensitivity even for $\lambda^{0,\text{signal}}=3$.
 These results highlight the substantial boost in signal detection from effectively leveraging between-drug similarities as facilitated by our models.
 In all settings, the FDR and average type I error is strictly controlled for all methods.
-The Poisson and ZIP-based DP local-global models perform similarly across all scenarios regardless of the presence of zero-inflated observed counts (sub-cases a and b), due to the flexibility of the DP-based nonparametric framework (see Remark Remark).
+The Poisson and ZIP-based DP local-global models perform similarly across all scenarios regardless of the presence of zero-inflated observed counts (sub-cases a and b), due to the flexibility of the DP-based nonparametric framework (see Section 4.1).
 All computations were performed on the cluster computing facility at the Center for Computational Research at the University at Buffalo, with each process running on 16 GB RAM and a clock speed of 1.6-3 GHz.
-The median computation time per replication for our DP local-global methods (both Poisson and ZIP) was 145 seconds.
+The median computation time per replication for our DP local-global methods (both Poisson and ZIP) is 48 minutes.
 6 REAL DATA ANALYSIS
 We use the statin data from Section 5 to demonstrate our method’s utility in real-world SRS data mining.
-Applying our DP local-global models (using the same MCMC and test statistic generation settings as used for simulated data analysis in Section 4) we compare detected signals against those from the DP Hu et al., BCPNN, GPS, and LRT.
-Table 2 presents the number of signals identified across statin drugs.
-The results indicate identical performance for our Poisson and ZIP models, detecting more signals than all competing methods.
+We fit our DP-Poisson and DP-ZIP local-global models to the dataset, setting the finite mixture approximation of the DP with $K=20$ components, and using the same MCMC and test statistic generation settings as used for simulated data analysis in Section 5).
+We compare the detected signals against those from the DP Hu et al., BCPNN, GPS, and LRT.
+Table 2 presents the number of signals identified across statin drugs by all these methods.
+The results indicate very similar performance for our Poisson and ZIP local-global models, both of which detect a substantially larger number of signals than all competing methods.
 Table 2: Summary table of discoveries from our proposed DP local-global models, and comparing them to the other existing methods.
 The count in each cell represents the number of AE signals detected by the method (along the columns), first separately for each drug and then in total (rows).
-The number of discoveries made uniquely by our DP Poisson local-global method but not by the other methods are presented in parentheses.
-Drug DP Poisson DP ZIP DP Hu et al. BCPNN GPS LRT
-Atorvastatin 33 33 (0) 33 (0) 28 (5) 26 (7) 28 (5)
-Fluvastatin 14 14 (0) 13 (1) 10 (4) 12 (2) 12 (2)
-Lovastatin 12 12 (0) 10 (2) 9 (3) 9 (3) 9 (3)
-Pravastatin 23 23 (0) 17 (6) 11 (12) 11 (12) 11 (12)
-Rosuvastatin 29 29 (0) 28 (1) 25 (4) 25 (4) 25 (4)
-Simvastatin 29 29 (0) 28 (1) 26 (3) 25 (4) 25 (4)
-Total 140 140 (0) 129 (11) 110 (31) 110 (32) 110 (30)
-To further examine the results, we focus on 20 PTs (AEs/$i$) with the highest average (across statin drugs) posterior medians of $\lambda_{ij}$ from our DP Poisson local-global model (Figure 4).
-The heatmap indicates non-detection (gray) or the degree of information sharing between drugs, measured via the posterior probabilities of $1-z_{ij}^{l}$ (white-blue gradient; darker blue signifies greater information sharing at each AE).
-Our method detects several rare AEs with $n_{ij}=2$, some also identified by certain competing methods (though no other method detects all these AEs simultaneously) and some not detected by any competing method (not shown in Figure 4; see Table 2).
-This, along with strong FDR control, highlights our method’s effectiveness in SRS signal detection.
-Figure 4: Heatmap of 20 selected PTs.
+The numbers in parentheses represent the differences in the number of discoveries made by each method when compared to our DP Poisson local-global method.
+Drug DP Poisson local-global DP ZIP local-global DP Hu et al. BCPNN GPS LRT
+Atorvastatin 308 306 (2) 307 (1) 234 (74) 203 (105) 126 (182)
+Fluvastatin 136 136 (0) 122 (14) 68 (68) 60 (76) 26 (110)
+Lovastatin 90 94 (-4) 59 (31) 37 (53) 30 (60) 11 (79)
+Pravastatin 226 227 (-1) 210 (16) 122 (104) 91 (135) 35 (191)
+Rosuvastatin 314 314 (0) 297 (17) 228 (86) 195 (119) 110 (204)
+Simvastatin 314 317 (-3) 257 (57) 220 (94) 183 (131) 94 (220)
+Total 1388 1394 (-6) 1252 (136) 909 (479) 762 (626) 402 (986)
+To further examine the signal detection results, we focus on our DP Poisson local-global model and zoom in on 15 selected PTs, whose signal strength estimates are plotted using heatmaps in Figure 4.
+The figure is divided into three panels based on the frequencies of the AEs, and the colors in the heatmaps indicate either non-detection (gray) or the estimated signal strengths measured through the posterior medians of $\lambda_{ij}$ (white-to-blue gradient with a darker blue signifying a stronger signal strength).
+Panel A displays the five most common AEs with the largest frequencies (total counts for the AEs across all six statin drugs being greater than $2000$) and the strongest estimated signal strengths (having the largest posterior medians of $\lambda_{ij}$ across all statin drugs $j$) that are detected as a signal for at least one statin drug by our method.
+The AEs displayed in this panel are all well-documented muscle-related conditions, such as rhabdomyolysis antons2006clinical ; tournadre2020statins and myopathy ballantyne2003risk ; abd2011statin , and their detection by our method is consistent with the performance of all competing methods.
+Panel B contains five semi-rare AEs (with observed counts across statin drugs ranging between 500 and 2000) that exhibit the strongest estimated signal strengths.
+This category primarily includes some specific types of disorders, such as nervous system disorders, and interestingly, several of these detected signals are not identified by any competing method for the newer or less frequently used statin drugs.
+For example, unlike other methods, our model detects lovastatin-induced drug intolerance as a signal, which has been flagged as a potential side effect of lovastatinahmad2014statin .
+Additionally, our method detects signals for paraesthesia and hypoaesthesia, symptoms of peripheral neuropathy, associated with lovastatin use; this finding is also consistent with previous research chong2004statin ; wang2021association .
+By appropriately borrowing strengths from the more commonly used statins, our method is successfully able to identify potential lovastatin-induced AE signals in these specific examples.
+Finally, Panel C focuses on five rare AEs (with observed counts across statin drugs being lower than 500) with the strongest estimated signal strengths.
+Within this group, our method identifies several AEs related to neurological and movement issues, such as movement disorder and swollen tongue, and some other mixed disorders.
+Many of these AEs were not originally discovered in clinical trials, but later were flagged as potential signals in specific, targeted pharmacovigilance investigations.
+For example, fluvastatin-induced swollen tongue, also known as angioedema, has been recently recorded kumar2024fluvastatin .
+Similar to the semi-rare AEs displayed in Panel B, our method is able to identify these rare AE signals for the less frequently used drug fluvastatin by appropriately sharing AE signal strength information across all statin drugs.
+Figure 4: Heatmaps of 15 selected PTs.
 Cells not detected as signals by our method are colored gray.
-White-to-blue shades represent the posterior probability of $1-z_{ij}^{l}$, with darker (lighter) blue indicating a higher (lower) probability that the signal strength originates from the global signal strength distribution.
+White-to-blue shades represent the median of the posterior of signal strength parameter $\lambda_{ij}$, with darker (lighter) blue indicating a stronger (weaker) signal strength.
+Panel A contains 5 common PTs among all statin drugs.
+Panel B contains 5 semi-rare PTs with new discovered signals by our proposed method.
+Panel C contains 5 rare PTs with new discovered signals by our proposed method.
+Count $n$, expected count $E$, and the 90% credible interval of the posterior distribution of $\lambda_{ij}$ are shown in each cell.
 The letters “H”, “B”, “G”, and “L” indicate signal detection by the DP Hu et al., BCPNN, GPS, and likelihood ratio test, respectively.
-Finally, we focus on inferring structural zeros through the estimated drug-specific zero-inflation probabilities $\omega_{j}$ as returned by our DP ZIP local-global model.
-For comparison, we also monitor the corresponding estimates provided by the frequentist ZIP (pseudo) LRT method.
-The two methods yield largely similar results: our approach produces point estimates (posterior mean) for $\omega_{j}$ as 0.02, 0.10, 0.11, 0.08, 0.02, and 0.03 for Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, respectively.
-The corresponding frequentist maximum likelihood estimates are 0.00, 0.07, 0.16, 0.06, 0.00, and 0.00, respectively.
 7 DISCUSSION
@@
 Signal detection is performed through hypothesis tests that control FDR and minimize FNR, both estimated from MCMC posteriors.
-Our approach achieves significantly higher sensitivity than existing methods, particularly for rarer AEs and/or new drugs.
+Our approach achieves significantly higher sensitivity than existing methods, particularly for rare AEs and/or newer drugs, with small observed counts.
+Application to the statin dataset demonstrates our model’s superior performance, which detects more plausible signals involving rare AEs and/or newer drugs than competing methods while reliably identifying well-established AEs.
+By enabling between-drug information sharing, our model can also discover potential signals among rare AEs and provide more nuanced detection for semi-rare AEs.
 There are several limitations with our current approach that motivate interesting future directions.
@@
 A reformulation via mixed membership models could further enhance interpretability by allowing the same AE to appear in multiple clusters/groups.
-Second, while we consider data-driven characterization of similarities in AE signal strengths between drugs, a more flexible approach acknowledging known Drug-Drug Interactions (DDI; Poleksic and Xie, 2019), possibly via appropriate prior distributions, may further enhance signal detection.
+Second, while we consider data-driven characterization of similarities in AE signal strengths between drugs, a more flexible approach acknowledging known Drug-Drug Interactions (DDI poleksic2019database ), possibly via appropriate prior distributions, may further enhance signal detection.
 Finally, our current framework aggregates drug-AE data across time points for a cross-sectional analysis.
 Extending the model to accommodate temporal variations in SRS data could enable formal identification of evolving AE patterns.
-Acknowledgment. The authors are deeply grateful to Prof. Marianthi Markatou at the University at Buffalo for her insightful comments and helpful suggestions on the manuscript.
+Acknowledgment. The authors are deeply grateful to Prof. Marianthi Markatou at the University at Buffalo for her insightful comments and helpful suggestions on an earlier version of the manuscript.
 Supplementary Material
-This supplement provides a brief review of existing SRS data mining methods, additional details about the DP local-global models, detailed steps for the MCMC algorithms implementing our models, additional computation details, and additional simulation results.
+This supplement provides additional details about the DP local-global models, detailed steps for the MCMC algorithms implementing our models, additional computation details, and additional simulation results.
 Sections, tables, and figures in this supplement are labeled with a prefix “S.”, such as, Section S.1, Table S.1, and Figure S.1, etc.
-Appendix S.1 A REVIEW OF EXISTING BAYESIAN/EMPIRICAL BAYESIAN SRS DATA MINING METHODS AND THE ZIP MODEL-BASED LIKELIHOOD RATIO TEST METHOD
-S.1.1 The GPS model
-The GPS method (DuMouchel,, 1999) employs an empirical Bayes approach for stabilized signal detection within a Poisson model-based framework for the report counts $n_{ij}$.
-Specifically, we assume $n_{ij}\sim$ Poisson$(E_{ij}\lambda_{ij})$, with $\lambda_{ij}$ parametrizing the signal strength for the drug-$j$ and AE-$i$ pair.
-A two-component gamma mixture prior for $\lambda_{ij}$ is considered:
-$\lambda_{ij}\sim\kappa\text{Gamma}(\alpha_{1},\beta_{1})+(1-\kappa)\text{Gamma}(\alpha_{2},\beta_{2}),$
-where $\kappa\in[0,1]$ is the mixing weight hyperparameter, and $\alpha_{1},\beta_{1}>0$, and $\alpha_{2},\beta_{2}>0$ are mixture component-specific gamma hyperparemeters.
-Here Gamma($\alpha$, $\beta$) denotes a probability distribution with density $f_{\text{gamma}}(y\mid\alpha,\beta)=\frac{1}{\Gamma(\alpha)}\beta^{\alpha}y^{\alpha-1}e^{-\beta y}$, $y>0$; $\alpha,\beta>0$.
-The posterior distribution of $\lambda_{ij}$ given $n_{ij}$, $E_{ij}$, and the hyperparameters is:
-$\lambda_{ij}\mid n_{ij},E_{ij}\sim\kappa_{ij}\text{Gamma}(\hat{\alpha_{1}}+n_{ij},\hat{\beta_{1}}+E_{ij})+(1-\kappa_{ij})\text{Gamma}(\hat{\alpha_{2}}+n_{ij},\hat{\beta_{2}}+E_{ij})$
-where $\kappa_{ij}=\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})/[\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})+(1-\kappa)f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{2},\beta_{2})]$.
-Under an empirical Bayes approach, the point estimates $\hat{\kappa}$, $\hat{\alpha}_{1}$, $\hat{\beta}_{1}$, $\hat{\alpha}_{2}$, and $\hat{\beta}_{2}$ for the hyperparameters are obtained by maximizing their marginal likelihoods given $n_{ij}$ and $E_{ij}$.
-The resulting posterior of $\lambda_{ij}$ conditional on these estimated hyperparameters is used for inference and signal detection.
-Specifically, DuMouchel and Pregibon, (2001) define the empirical Bayesian score $E(\log_{2}(\lambda_{ij})\mid n_{ij},E_{ij},\hat{\kappa},\hat{\alpha}_{1},\hat{\beta}_{1},\hat{\alpha}_{2},\hat{\beta}_{2})$ as a point estimate of signal strength for the drug-$j$ and AE-$i$ pair, and also suggest determining the pair to be a signal if the 5th percentile (EB05) of the above conditional posterior for $\lambda_{ij}$ exceeds 2.
-S.1.2 The BCPNN method
-The BCPNN approach (Bate et al.,, 1998) assumes that conditional on the grand total count $n_{\bullet\bullet}$, the cell-specific counts $n_{ij}$, the marginal row totals $n_{i\bullet}$, and the marginal column totals $n_{\bullet j}$ each distributed as binomial-beta: (a) $n_{ij}\mid p_{ij}\sim\text{Binomial}(n_{\bullet\bullet},p_{ij}),\ p_{ij}\sim\text{Beta}(\alpha_{ij},\beta_{ij})$, (b) $n_{i\bullet}\mid p_{i\bullet}\sim\text{Binomial}(n_{\bullet\bullet},p_{i\bullet}),\allowbreak\ p_{i\bullet}\sim\text{Beta}(\alpha_{i\bullet},\beta_{i\bullet})$, (c) $n_{\bullet j}\mid p_{\bullet j}\sim\text{Binomial}(n_{\bullet\bullet},p_{\bullet j}),\allowbreak\ p_{\bullet j}\sim\text{Beta}(\alpha_{\bullet j},\beta_{\bullet j})$, where $p_{ij}$ is the cell reporting rate, and $p_{i\bullet}$ and $p_{\bullet j}$ are the marginal row and column reporting rates.
-(Note that the binomial model assumptions for $n_{ij}$, $n_{i\bullet}$ and $n_{\bullet j}$ can be understood as consequences of an independent Poisson model for $n_{ij}$ with certain choices of $p_{ij}$, $p_{i\bullet}$, and $p_{\bullet j}$.).
-Bate et al., (1998) consider the information content $\text{IC}_{ij}=\log_{2}\left[p_{ij}/(p_{i\bullet}p_{\bullet j})\right]$ as a measure of signal strength for the drug-$j$ AE-$i$ pair, with a higher IC indicating a stronger association.
-The authors suggest an empirical Bayes estimation of these quantities using $\hat{\alpha}_{ij}=\hat{\alpha}_{i\bullet}=\hat{\beta}_{i\bullet}=\hat{\alpha}_{\bullet j}=\hat{\beta}_{\bullet j}=1$ and $\hat{\beta}_{ij}=1/(E(p_{i\bullet}\mid n_{i\bullet})E(p_{\bullet j}\mid n_{\bullet j}))-1$, with the denominator of $\hat{\beta}_{ij}$ being computed using the posterior distributions $p_{i\bullet}\mid n_{i\bullet}\sim\text{Beta}(\hat{\alpha}_{i\bullet}+n_{i\bullet},\hat{\beta}_{i\bullet}+n_{\bullet\bullet}-n_{i\bullet})$ and $p_{\bullet j}\mid n_{\bullet j}\sim\text{Beta}(\hat{\alpha}_{\bullet j}+n_{\bullet j},\hat{\beta}_{\bullet j}+n_{\bullet\bullet}-n_{\bullet j})$.
-The consequent (empirical Bayes) posterior distribution of $\text{IC}_{ij}$ can be approximated using Monte Carlo simulation or an asymptotic normal distribution.
-A signal is detected at the drug-$j$ and AE-$i$ pair if the 2.5th posterior percentile for $\text{IC}_{ij}$ exceeds 0.
-S.1.3 DP Poisson method of Hu et al., (2015)
-Hu et al., (2015) consider the same likelihood $n_{ij}\sim$ Poisson($\lambda_{ij}E_{ij}$) as in GPS, but extend the two-gamma mixture prior for $\lambda_{ij}$ to a DP (infinite) mixture prior:
-$\lambda_{ij}\mid(\alpha_{j},M_{0j})\sim\text{DP}(\alpha,M_{0j})=\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}}$
-where the base distribution $M_{0j}\equiv$ Gamma$(a_{j},a_{j})$ generates $\theta_{hj}\sim M_{0j}$ and the concentration hyperparameter $\alpha>0$ generates the mixture weights $\{\eta_{hj}\}_{h=1}^{\infty}$ satisfying $\sum_{h=1}^{\infty}\eta_{hj}=1$ with
-$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj}),\ h=2,\dots;\ v_{hj}\sim\text{independent Beta}(1,\alpha)\text{ for }h=1,\dots$ (5)
-leveraging a stick-breaking representation (Sethuraman,, 1994) for each $j$.
-Hu et al., (2015) consider full Bayesian inference on $\lambda_{ij}$ by first eliciting a prior distribution on the hyperparameters $a_{j}$ and $\alpha_{j}$—specifically, independent vague priors $a_{j}\sim\text{Uniform}(0,1)$ and $\alpha_{j}\sim\text{Uniform}(0.2,10)$; $j=1,\dots J$—and then utilizing MCMC sampling from the resulting joint posterior.
-Using MCMC posterior draws for $\lambda_{ij}$, a signal at the drug-$j$ and AE-$i$ pair is identified if the 5th posterior percentile of $\lambda_{ij}$ exceeds a threshold of 2.
-S.1.4 ZIP model based Pseudo LRT (frequentist) method
-The likelihood ratio test based method and its variants (Huang et al.,, 2011, 2013, 2014) have an advantage in FDR control while maintaining high sensitivity.
-Additionally, the use of a post-hoc threshold makes it more suitable for practical applications.
-Consider the reporting count $n_{ij}\sim\text{ZIP}(n=n_{ij}\mid\lambda_{ij}\times E_{ij},\omega_{j})$, where $\lambda_{ij}$ is the relative reporting rate and $E_{ij}=n_{i\cdot}n_{\cdot j}/n_{\cdot\cdot}$ represents the expected count under independence.
-The pseudo LRT, as proposed by Chakraborty et al., (2022), is a recent method designed to accommodate the zero-inflated nature of SRS data while easing computational demands.
-Similar to the standard LRT method, this approach first tests a global alternative hypothesis $H_{aj}:\lambda_{ij}>1$ for at least one $i\in\{1,\dots,I\}$ against the null hypothesis $H_{0j}:\lambda_{ij}=1$ for all $i\in\{1,\dots,I\}$ for a given drug-$j$.
-If $H_{aj}$ is rejected, a step-down procedure is used to conduct individual tests on $H_{aij}:\lambda_{ij}>1$.
-The ordinary LRT methods estimate the nuisance parameter $\omega_{j}$ from $H_{0j}$ and $H_{0j}\cup H_{aj}$ individually, the pseudo LRT is constructed using a global estimate $\hat{\omega}_{j}$ of $\omega_{j}$ consistent under $H_{0j}\cup H_{aj}$ in both the numerator and the denominator of the likelihood ratio test statistic.
-The resulting statistic is given by
-$\text{LR}_{ij}=\exp(-(\hat{\lambda}_{ij}-1)E_{ij})\hat{\lambda}_{ij}^{n}.$
-The maximum likelihood ratio statistic for drug-$j$ is then defined as $\text{MLR}_{j}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\}\end{subarray}}\text{LR}_{ij}$.
-In the global test, a parametric bootstrap can be used to approximate the p-value $P_{j}=\Pr(\text{MLR}_{j}\geq\text{MLR}_{j}^{\text{obs}})$ using (parametric bootstrap) draws for $\text{LR}_{ij}$ under $H_{0j}$ and hence $\text{MLR}_{j}$, where $\text{MLR}_{j}^{\text{obs}}$ is the observed test statistic.
-If the global test is rejected, the step-down p-values can be computed for the individual tests using the same parametric bootstrap samples for $\text{MLR}_{j}$ and using the individual observed likelihood ratio test statistic values, such that $P_{ij}=\Pr(\text{LR}_{ij}\geq\text{MLR}_{ij}^{\text{obs}})$.
-For multiple drugs, the same approach can be applied, defining an overall maximum likelihood ratio as $\text{MLR}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\},j\in\{1,\dots,J\}\end{subarray}}\text{LR}_{ij}$.
-This step-down procedure avoids $\alpha$-inflation by using the same global test statistic, thus controlling the false discovery rate at the $\alpha$ level.
-A consistent estimate of $\omega_{j}$ is obtained by maximizing the profile likelihood $\ell_{j}(\omega_{j})$, with $\hat{\lambda}_{ij}=\max\{1,n/E_{ij}\}$ as the MLE of $\lambda_{ij}$.
-The inference on zero-inflation is based on the estimate $\hat{\eta}_{ij}(\lambda_{ij},\hat{\omega}_{j})=\frac{\hat{\omega}_{j}}{\hat{\omega}_{j}+(1-\hat{\omega}_{j})\exp(-\lambda_{ij}E_{ij})}$, representing the estimated probability of zero-inflation for cells with zero count.
-A prediction of the zero-inflation indicator $z_{ij}$ can be made by generating Bernoulli samples from $\hat{\eta}_{ij}$.
 Appendix S.1 ADDITIONAL DETAILS FOR THE DP LOCAL-GLOBAL FRAMEWORK
@@
 And we propose a simpler shrinkage model as an alternative for the “Other drugs” as described in Section 4.2.
-S.2.1 Complete hierarchical Bayesian representation of our models
+S.1.1 Complete hierarchical Bayesian representation of our models
 In Section 4.2 in the main text, we present the concept of the DP Poisson local-global model.
 The fully formulated hierarchical representations are provided in the following two sections for the DP Poisson and the DP ZIP local-global model.
-S.2.1.1 DP Poison local-global model
+S.1.1.1 DP Poison local-global model
 A hierarchical representation of the complete model, including vague priors for the hyperparameters, is presented as follows.
@@
 $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$
-[…]
+$\text{DP local-global prior for a non-baseline drug }j\in\{1,\dots,J-1\}\!:$
 $\lambda_{ij}\sim z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$
@@
 $\eta_{1}=v_{1},\ \eta_{h}=v_{h}\prod_{t=1}^{h-1}(1-v_{t}),h\geq 2$
-$v_{h}\sim\text{Beta}(1,\alpha)$
+$v_{h}\sim\text{Beta}(1,\alpha),h=1,\dots,K-1;\ v_{K}=1$
 Gamma shrinkage prior for a baseline drug-$J$:
@@
 $\tau\sim\text{Cauchy}^{+}(0,0.5)$
-S.2.1.2 DP ZIP local-global model
+S.1.1.2 DP ZIP local-global model
 A hierarchical representation of the complete model, including vague priors for the hyperparameters, is presented as follows.
@@
 $y\sim\text{Bernoulli}(\omega_{j})$
-[…]
+$\text{DP local-global prior for a non-baseline drug }j\in\{1,\dots,J-1\}\!:$
 $\lambda_{ij}\sim z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$
@@
 DP local prior for drug-$j$:
-$D_{j}\equiv DP(\alpha_{j},G(\beta_{j}))\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}}\approx\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$
-$\theta_{hj}\sim G_{0j}\equiv\text{Gamma}(\text{shape}=\beta_{j},\text{rate}=\beta_{j})$
+$D_{j}\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}}\approx\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$
+$\theta_{hj}\sim G(\beta_{j})\equiv\text{Gamma}(\text{shape}=\beta_{j},\text{rate}=\beta_{j})$
 $\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj}),h\geq 2$
-$v_{hj}\sim\text{Beta}(1,\alpha_{j});v_{Kj}=1$
+$v_{hj}\sim\text{Beta}(1,\alpha_{j}),h=1,\dots,K-1;\ v_{Kj}=1$
 DP global prior:
-$D\equiv DP(\alpha,G(\beta))\equiv\sum_{h=1}^{\infty}\eta_{h}\delta_{\{\theta_{h}\}}\approx\sum_{h=1}^{K}\eta_{h}\delta_{\{\theta_{h}\}}$
-$\theta_{h}\sim G_{0}\equiv\text{Gamma}(\beta,\beta)$
+$D\equiv\sum_{h=1}^{\infty}\eta_{h}\delta_{\{\theta_{h}\}}\approx\sum_{h=1}^{K}\eta_{h}\delta_{\{\theta_{h}\}}\sim DP(\alpha,G(\beta))$
+$\theta_{h}\sim G(\beta)\equiv\text{Gamma}(\text{shape}=\beta,\text{rate}=\beta)$
 $\eta_{1}=v_{1},\ \eta_{h}=v_{h}\prod_{t=1}^{h-1}(1-v_{t}),h\geq 2$
-$v_{h}\sim\text{Beta}(1,\alpha)$
+$v_{h}\sim\text{Beta}(1,\alpha),h=1,\dots,K-1;\ v_{K}=1$
 Gamma shrinkage prior for a baseline drug-$J$:
@@
 vague proper priors for hyperparameters:
-$\pi^{g}\sim\text{Uniform}(0,1)$
+$\pi^{l}\sim\text{Uniform}(0,1)$
 $\alpha_{j},\alpha\sim\text{Gamma}(\text{shape}=1,\text{rate}=3)$
@@
 $\tau\sim\text{Cauchy}^{+}(0,0.5)$
-S.2.2 Some notes on the “global-only” model
+S.1.2 Some notes on the “global-only” model
 A special case of our local-global models leads to a reduced “global-only” model when $\pi^{g}=0$, i.e., $\lambda_{ij}=\lambda_{i}^{g}$ for $j=1,\dots,J-1$.
@@
 In this section, we first provide MCMC algorithms implementing our models in Section S.2.1.
-The MCMC algorithms are derived using finite mixture approximation of the underlying DP processes (Ishwaran and Zarepour,, 2002; Frühwirth-Schnatter and Malsiner-Walli,, 2019), and combine Gibbs sampling updates with stepping-out slice sampling updates (Neal,, 2003) for various parameters.
+The MCMC algorithms are derived using finite mixture approximation of the underlying DP processes [23, 16], and combine Gibbs sampling updates with stepping-out slice sampling updates [31] for various parameters.
 Some additional notes on the MCMC implementation, including initial value generation for our MCMC algorithms are then provided in Section S.2.2.
 Next we provide detailed notes on the grid-search method identifying optimal quantile-level $p$ and threshold $k$ used for our hypothesis tests for signal detection (Section S.2.3).
-The JAGS code (Plummer et al.,, 2003) for Hu et al., (2015)’s models is shown as a reference in Section S.1.3.
+The JAGS code [32] for Hu et al.’s model [18] is shown as a reference in Section S.2.4.
 The example statin dataset is included in Section S.2.5.
 In our notation below, $\text{DPSB}_{K}(\bm{\eta}=(\eta_{1},\dots,\eta_{K}),\bm{\theta}=(\theta_{1},\dots,\theta_{K}),\alpha,\beta)$ denotes parameters associated with the model $\sum_{h=1}^{K}\eta_{h}1_{\{\theta_{h}\}}\sim\text{DP}(\alpha,G(\beta))$, comprising the stick-breaking weights $\bm{\eta}=(\eta_{1},\dots,\eta_{K})$ and atoms $\bm{\theta}=(\theta_{1},\dots,\theta_{K})$ and DP concentration hyperparamter $\alpha$ and base (gamma) distribution hyperparameter $\beta$.
-S.3.1 MCMC algorithms
+S.2.1 MCMC algorithms
 Algorithm S.1 One Iteration of the MCMC algorithm updating $\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$ parameters from its posterior given data $(n_{i},E_{i},z_{i})$ under the model: $n_{i}\sim\text{Pois}(\lambda_{i}E_{i})$, $\lambda_{i}\sim\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$, $\alpha\sim\text{Exponential}(\text{rate}=\psi_{\alpha})$, $\beta\sim\pi_{\beta}(\beta)$, and $z_{i}$ indicating if observation $i$ contributes to the model ($z_{i}=1$) or not ($z_{i}=0$); $i=1,\dots,N$.
@@
 Direct random sampling from this density is not feasible.
-Instead, we use stepping-out slice sampling (Neal,, 2003) to generate Markov chain draws for $\beta$.
+Instead, we use stepping-out slice sampling [31] to generate Markov chain draws for $\beta$.
   4. 4.
@@
     $\tilde{n}_{j}=\sum_{j=1}^{J-1}n_{ij}(1-z_{ij}^{l}),\quad\tilde{E}_{j}=\sum_{j=1}^{J-1}E_{ij}(1-z_{ij}^{l}).$
-    1. 1.1.
+    1. $1.$
       Set $\lambda_{i}^{\text{g}}=\lambda_{i}$ returned from Algorithm S.1.
@@
     where
-    $p_{ij}^{l}=\frac{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})}{f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})+(1-\pi^{l})f_{\text{Poisson}}(n_{ij}\mid\lambda_{i}^{g}E_{ij})}.$
+    $p_{ij}^{l}=\frac{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})}{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})+(1-\pi^{l})f_{\text{Poisson}}(n_{ij}\mid\lambda_{i}^{g}E_{ij})}.$
   4. 4.
@@
     Run Algorithm S.2 with all $(n_{ij},E_{ij})$ such that $y_{ij}=0$.
-S.3.2 Initial value generation for the MCMC algorithms
-Given the observed data $n_{ij}$ and $E_{ij}$ for $i=1,\dots,I$, $j=1,\dots,J$, and a prespecified number $K$ of DP mixture components for finite approximation, we initialize the number of nonempty components $K_{+}$ to the smallest integer greater than $\min\{K/2,I-1\}$.
+S.2.2 Initial value generation for the MCMC algorithms
+Given the observed data $n_{ij}$ and $E_{ij}$ for $i=1,\dots,I$, $j=1,\dots,J$, and a prespecified number $K$ of DP mixture components for finite approximation.
+The default value for $K$ in our software is set to $\max(\log(IJ),10)$, a setting that is generally sufficient for most cases.
+For the simulation study in Section 5, we used $K=10$, which satisfies the condition $K>\log(1492\times 7)$.
+In the data analysis presented in Section 6, we chose a more conservative value of $K=20$ to allow for a larger number of atoms in the DP, facilitating a more cautious interpretation of the results.
+We initialize the number of nonempty components $K_{+}$ to the smallest integer greater than $\min\{K/2,I-1\}$.
 The local-global indicators $z_{ij}^{l}$ are randomly initialized by independently sampling each cell $(i,j)$ from a $\text{Bernoulli}(0.5)$ distribution.
 The DP parameters for each local and global DP process are then initialized using the subset of data $(n_{ij},E_{ij})$ assigned to that process, as determined by the indicators $z_{ij}^{l}$, for $i=1,\dots,J-1$.
-For the initialization of parameters in the local DP prior for the $j$-th drug ($j=1,\dots,J-1$), we perform $k$-means clustering with the initialized $K_{+}$ number of clusters on the set ${(n_{ij}+0.5)/(E_{ij}+0.5):z_{ij}=1,\ i=1,\dots,I}$, and use the resulting cluster centers as initial values for the atoms $\theta_{hj}$, $h=1,\dots,K_{+}$, of the non-empty clusters.
+For the initialization of parameters in the local DP prior for the $j$-th drug ($j=1,\dots,J-1$), we perform $k$-means clustering with the initialized $K_{+}$ number of clusters on the set $\{(n_{ij}+0.5)/(E_{ij}+0.5):z_{ij}=1,\ i=1,\dots,I\}$, and use the resulting cluster centers as initial values for the atoms $\theta_{hj}$, $h=1,\dots,K_{+}$, of the non-empty clusters.
 The probability weights $\eta_{hj}$ for these non-empty clusters are initialized using the relative frequencies of the $k$-means clusters, and the clusters are then relabeled to ensure $\eta_{1j}\geq\eta_{2j}\geq\dots\geq\eta_{K_{+}j}$.
@@
 Finally, the reference drug $J$ shrinkage hyperparameter $\tau$ is initialized at $1$, while the zero-inflation parameters $\omega_{j}$ (if included in the model) are initialized each at $0.01$.
-S.3.3 Optimization method in pFDR calculation
+S.2.3 Optimization method in pFDR calculation
 A sequence of threshold values for $k$, ranging from 1.1 to 3 in small increments, is established.
@@
 For every quantile level in the $p$-grid, we obtain a matrix of test statistics $T_{ij}(p)$ based on the $p$-th posterior quantiles for $\lambda_{ij}$ computed from their MCMC draws.
-Subsequently, we compute FDR$(p,k)$ and FNR$(p,k)$ estimates using the formulas presented in the main text (Section 3.4) for all possible $p$ and $k$ defined by the grid.
-From the set of all $(p,k)$ pairs, we then retain only those pairs with computed FDR $\leq\alpha$ for a preset threshold $\alpha$ ($=0.05$ in our analyses).
+Subsequently, we compute FDR(p,k)(p,k) and FNR(p,k)(p,k) estimates using the formulas presented in the main text (Section 4.4) for all possible $p$ and $k$ defined by the grid.
+From the set of all $(p,k)$ pairs, we then retain only those pairs with computed FDR $\leq\alpha$ for a preset threshold $\alpha$ (=0.05=0.05 in our analyses).
 Finally, among the retained $(p,k)$, we identify the pair that minimizes FNR.
 If multiple $(p,k)$ yield the same minimum FNR, we select the pair with the smallest $p$ and the largest $k$.
-S.3.4 JAGS code for DP Hu et al. model
+S.2.4 JAGS code for DP Hu et al. model
 1 # Likelihood
@@
 Listing 1: JAGS exmpale
-S.3.5 Example statin dataset
-We present the reference SRS contingency table for statin drug in Table S.1. It is used for data generation in Section 4 and data analysis in Section 5.
-Table S.1: The SRS contingency table summarized from the FDA FEARS database (Q3 2014-Q4 2020) which catalogs a total of 63,976,610 reported cases.
+S.2.5 Example statin dataset
+We present the first 30 rows of the reference SRS contingency table for statin drug in Table S.1. It is used for data generation in Section 5 and data analysis in Section 6.
+Table S.1: The first 30 rows of the SRS contingency table summarized from the FDA FEARS database (Q3 2014-Q4 2020) which catalogs a total of 63,976,610 reported cases.
 Adverse Event Atorvastatin Fluvastatin Lovastatin Pravastatin Rosuvastatin Simvastatin Other
-Acute Kidney Injury 1353 42 7 154 689 823 355651
-Anuria 71 0 0 2 43 62 10403
-Blood Calcium Decreased 14 2 0 0 110 17 15918
-Blood Creatine Phosphokinase Abnormal 34 0 0 0 8 11 261
-Blood Creatine Phosphokinase Increased 1175 125 32 200 562 768 23805
-Blood Creatine Phosphokinase Mm Increased 2 0 0 0 9 0 14
-Blood Creatinine Abnormal 27 0 0 0 5 3 3385
-Blood Creatinine Increased 227 10 0 17 210 97 74742
-Chromaturia 340 10 6 33 174 114 19294
-Chronic Kidney Disease 152 16 2 19 177 37 339179
-Compartment Syndrome 53 0 0 1 12 12 2644
-Creatinine Renal Clearance Decreased 6 0 0 2 124 6 7768
-Diaphragm Muscle Weakness 14 0 0 0 8 1 94
-Electromyogram Abnormal 2 0 0 2 0 0 132
-End Stage Renal Disease 30 0 0 0 19 6 97553
-Glomerular Filtration Rate Abnormal 8 0 0 0 1 0 1069
-Glomerular Filtration Rate Decreased 59 1 0 8 39 29 13190
-Hypercreatininaemia 0 0 0 0 8 0 648
-Hypocalcaemia 36 0 0 16 8 18 23102
-Muscle Disorder 291 2 7 21 191 87 7329
-Muscle Enzyme Increased 48 1 0 0 13 9 410
-Muscle Fatigue 85 0 2 16 30 39 4257
-Muscle Haemorrhage 24 0 0 5 13 4 3806
-Muscle Necrosis 68 2 0 1 10 20 662
-Muscle Rupture 181 25 0 61 36 120 3219
-Muscular Weakness 1857 45 31 152 808 859 111003
-Musculoskeletal Discomfort 137 18 15 25 187 93 19931
-Musculoskeletal Disorder 56 3 0 9 65 73 25881
-Musculoskeletal Pain 420 3 2 38 324 228 82576
-Myalgia 5362 341 151 939 2757 3216 143819
-Myasthenic Syndrome 1 0 9 0 3 6 643
-Myoglobin Blood Increased 71 4 0 4 16 39 1003
-Myoglobin Blood Present 2 0 0 0 0 0 8
-Myoglobin Urine Present 0 0 0 0 1 2 70
-Myoglobinaemia 15 0 0 0 0 0 62
-Myoglobinuria 26 4 0 1 1 9 296
-Myopathy 849 64 45 145 377 544 6695
-Myopathy Toxic 31 0 0 1 2 21 457
-Myositis 219 8 10 28 62 141 7482
-Necrotising Myositis 279 0 0 2 10 52 278
-Oliguria 52 0 0 4 24 37 7590
-Renal Failure 534 26 11 69 225 195 250710
-Renal Impairment 390 52 11 37 161 169 103343
-Renal Tubular Necrosis 40 0 0 12 10 24 12762
-Rhabdomyolysis 2041 52 44 163 936 1376 31707
-Tendon Discomfort 9 0 0 3 22 10 794
-Other Pt 180699 4886 2845 20296 113960 76068 61724222
+Abasia 117 0 7 9 84 45 18046
+Abdominal Adhesions 10 0 0 0 3 2 3831
+Abdominal Discomfort 501 33 14 92 497 217 214807
+Abdominal Distension 267 7 5 57 251 262 101261
+Abdominal Pain 655 27 7 116 428 372 232708
+Abdominal Pain Lower 60 1 1 2 52 22 22527
+Abdominal Pain Upper 696 14 15 78 554 266 197925
+Abdominal Rigidity 12 0 0 1 1 12 2919
+Abdominal Symptom 7 0 0 0 2 2 718
+Abdominal Tenderness 20 0 0 0 5 10 5783
+Abnormal Behaviour 72 2 5 9 23 77 35802
+Abnormal Faeces 23 0 0 1 18 8 7982
+Abnormal Loss Of Weight 89 0 1 3 15 79 9042
+Abnormal Sensation In Eye 4 0 0 0 3 0 2571
+Abortion Induced 10 0 0 3 1 55 6668
+Abortion Spontaneous 32 2 0 5 12 4 28245
+Abscess Limb 4 0 0 0 4 0 5258
+Accident 24 0 0 9 28 6 12034
+Accident At Work 0 0 0 0 14 2 2087
+Accidental Exposure To Product 13 2 1 3 12 7 46642
+Accidental Overdose 56 0 0 11 7 36 42543
+Accidental Underdose 4 0 0 0 2 0 7423
+Acidosis 55 3 0 4 4 22 8117
+Acne 28 0 1 10 35 22 105646
+Activities Of Daily Living Impaired 135 1 3 15 82 200 14483
+Acute Coronary Syndrome 105 1 0 1 66 63 9259
+Acute Generalised Exanthematous Pustulosis 30 0 0 0 11 3 8161
+Acute Hepatic Failure 107 0 10 1 18 66 14260
 Appendix S.3 ADDITIONAL SIMULATION RESULTS
@@
 Frequentist and Bayesian MSE are compared to further evaluate the loss of the estimation.
-S.4.1 Simulation results for DP-based models
+S.3.1 Simulation results for DP-based models
 We present the four evaluation metrics in Tables S.2, S.3, S.4, and S.5 to provide a clearer numerical comparison beyond Figure 3.
@@
 Table S.2: Simulation of FDR for DP-based models
-Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
-Case 1a Weak correlation No ZI 1.1 0.0035 0.0022 0.0050 0.0102
-1.2 0.0072 0.0063 0.0056 0.0171
-1.3 0.0081 0.0088 0.0053 0.0172
-1.4 0.0074 0.0077 0.0061 0.0168
-1.5 0.0081 0.0079 0.0059 0.0149
-1.7 0.0057 0.0057 0.0050 0.0133
-1.9 0.0049 0.0047 0.0039 0.0108
-2.0 0.0053 0.0056 0.0046 0.0103
-2.5 0.0045 0.0046 0.0043 0.0106
-3.0 0.0031 0.0029 0.0034 0.0086
-Case 1b Weak correlation ZI 1.1 0.0010 0.0017 0.0060 0.0197
-1.2 0.0044 0.0034 0.0096 0.0214
-1.3 0.0067 0.0070 0.0070 0.0156
-1.4 0.0086 0.0085 0.0051 0.0157
-1.5 0.0089 0.0092 0.0059 0.0138
-1.7 0.0067 0.0065 0.0056 0.0136
-1.9 0.0058 0.0057 0.0048 0.0117
-2.0 0.0049 0.0050 0.0045 0.0118
-2.5 0.0032 0.0034 0.0032 0.0089
-3.0 0.0036 0.0037 0.0042 0.0095
-Case 2a Moderate correlation No ZI 1.1 0.0000 0.0010 0.0065 0.0193
-1.2 0.0107 0.0128 0.0063 0.0150
-1.3 0.0114 0.0132 0.0069 0.0160
-1.4 0.0120 0.0134 0.0060 0.0146
-1.5 0.0095 0.0098 0.0069 0.0157
-1.7 0.0072 0.0079 0.0059 0.0145
-1.9 0.0052 0.0056 0.0058 0.0128
-2.0 0.0042 0.0045 0.0048 0.0125
-2.5 0.0035 0.0035 0.0041 0.0102
-3.0 0.0033 0.0032 0.0032 0.0093
-Case 2b Moderate correlation ZI 1.1 0.0000 0.0000 0.0055 0.0122
-1.2 0.0093 0.0106 0.0065 0.0186
-1.3 0.0117 0.0130 0.0069 0.0168
-1.4 0.0102 0.0114 0.0061 0.0153
-1.5 0.0066 0.0076 0.0054 0.0149
-1.7 0.0053 0.0057 0.0058 0.0134
-1.9 0.0039 0.0040 0.0059 0.0126
-2.0 0.0047 0.0050 0.0045 0.0110
-2.5 0.0039 0.0037 0.0042 0.0103
-3.0 0.0030 0.0032 0.0041 0.0094
-Case 3a Strong correlation No ZI 1.1 0.0000 0.0000 0.0073 0.0133
-1.2 0.0084 0.0114 0.0074 0.0183
-1.3 0.0073 0.0084 0.0080 0.0184
-1.4 0.0086 0.0096 0.0069 0.0145
-1.5 0.0061 0.0069 0.0067 0.0150
-1.7 0.0045 0.0050 0.0066 0.0136
-1.9 0.0045 0.0050 0.0051 0.0129
-2.0 0.0039 0.0042 0.0049 0.0118
-2.5 0.0033 0.0035 0.0037 0.0096
-3.0 0.0030 0.0029 0.0031 0.0085
-Case 3b Strong correlation ZI 1.1 0.0000 0.0000 0.0040 0.0171
-1.2 0.0106 0.0120 0.0084 0.0162
-1.3 0.0077 0.0090 0.0063 0.0161
-1.4 0.0078 0.0083 0.0061 0.0153
-1.5 0.0068 0.0073 0.0051 0.0125
-1.7 0.0044 0.0049 0.0049 0.0119
-1.9 0.0030 0.0032 0.0050 0.0112
-2.0 0.0035 0.0038 0.0052 0.0113
-2.5 0.0033 0.0034 0.0040 0.0092
-3.0 0.0036 0.0037 0.0039 0.0094
+Case $\lambda$
+DP Poisson
+local-global
+DP ZIP
+local-global
+DP Hu et al.
+DP Hu et al.
+posterior quantile optim
+Case 1a
+Weak correlation
+No ZI
+1.1 0.0030 0.0030 0.0040 0.0080
+1.2 0.0084 0.0111 0.0075 0.0144
+1.3 0.0119 0.0117 0.0074 0.0152
+1.4 0.0084 0.0088 0.0040 0.0103
+1.5 0.0064 0.0063 0.0037 0.0101
+1.7 0.0068 0.0067 0.0040 0.0082
+1.9 0.0056 0.0057 0.0038 0.0077
+2.0 0.0055 0.0057 0.0030 0.0071
+2.5 0.0047 0.0048 0.0029 0.0067
+3.0 0.0042 0.0042 0.0026 0.0056
+Case 1b
+Weak correlation
+30% ZI
+1.1 0.0010 0.0000 0.0010 0.0053
+1.2 0.0079 0.0079 0.0060 0.0114
+1.3 0.0084 0.0088 0.0048 0.0102
+1.4 0.0074 0.0076 0.0044 0.0103
+1.5 0.0063 0.0065 0.0044 0.0098
+1.7 0.0058 0.0058 0.0031 0.0089
+1.9 0.0056 0.0057 0.0037 0.0080
+2.0 0.0055 0.0055 0.0037 0.0078
+2.5 0.0046 0.0045 0.0028 0.0065
+3.0 0.0037 0.0037 0.0022 0.0049
+Case 2a
+Moderate correlation
+No ZI
+1.1 0.0010 0.0000 0.0000 0.0026
+1.2 0.0111 0.0128 0.0061 0.0153
+1.3 0.0083 0.0082 0.0070 0.0148
+1.4 0.0064 0.0067 0.0052 0.0108
+1.5 0.0074 0.0072 0.0052 0.0101
+1.7 0.0046 0.0047 0.0035 0.0078
+1.9 0.0041 0.0041 0.0036 0.0079
+2.0 0.0032 0.0032 0.0031 0.0075
+2.5 0.0030 0.0030 0.0028 0.0063
+3.0 0.0024 0.0023 0.0026 0.0056
+Case 2b
+Moderate correlation
+30% ZI
+1.1 0.0033 0.0005 0.0005 0.0082
+1.2 0.0124 0.0142 0.0078 0.0159
+1.3 0.0105 0.0110 0.0059 0.0133
+1.4 0.0056 0.0057 0.0047 0.0121
+1.5 0.0056 0.0054 0.0039 0.0098
+1.7 0.0045 0.0045 0.0034 0.0088
+1.9 0.0033 0.0032 0.0031 0.0073
+2.0 0.0027 0.0029 0.0032 0.0073
+2.5 0.0025 0.0026 0.0030 0.0065
+3.0 0.0025 0.0025 0.0029 0.0062
+Case 3a
+Strong correlation
+No ZI
+1.1 0.0027 0.0017 0.0013 0.0032
+1.2 0.0129 0.0132 0.0052 0.0135
+1.3 0.0105 0.0106 0.0059 0.0144
+1.4 0.0077 0.0076 0.0054 0.0122
+1.5 0.0059 0.0061 0.0049 0.0110
+1.7 0.0040 0.0041 0.0045 0.0101
+1.9 0.0040 0.0038 0.0032 0.0081
+2.0 0.0030 0.0029 0.0029 0.0073
+2.5 0.0022 0.0021 0.0030 0.0064
+3.0 0.0015 0.0017 0.0026 0.0058
+Case 3b
+Strong correlation
+30% ZI
+1.1 0.0000 0.0000 0.0000 0.0074
+1.2 0.0102 0.0106 0.0061 0.0153
+1.3 0.0100 0.0100 0.0051 0.0120
+1.4 0.0076 0.0075 0.0044 0.0115
+1.5 0.0059 0.0058 0.0044 0.0109
+1.7 0.0043 0.0041 0.0039 0.0094
+1.9 0.0031 0.0032 0.0036 0.0082
+2.0 0.0031 0.0031 0.0033 0.0078
+2.5 0.0024 0.0023 0.0026 0.0063
+3.0 0.0015 0.0015 0.0026 0.0059
 Table S.3: Simulation of sensitivity for DP-based models
-Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
-Case 1a Weak correlation No ZI 1.1 0.0016 0.0014 0.0009 0.0021
-1.2 0.0427 0.0422 0.0346 0.0435
-1.3 0.1227 0.1243 0.0899 0.1041
-1.4 0.1923 0.1942 0.1432 0.1606
-1.5 0.2444 0.2472 0.1915 0.2115
-1.7 0.3153 0.3180 0.2685 0.2897
-1.9 0.3692 0.3707 0.3258 0.3475
-2.0 0.3926 0.3957 0.3495 0.3720
-2.5 0.4706 0.4729 0.4346 0.4554
-3.0 0.5251 0.5273 0.4920 0.5128
-Case 1b Weak correlation ZI 1.1 0.0012 0.0012 0.0011 0.0022
-1.2 0.0406 0.0399 0.0331 0.0416
-1.3 0.1216 0.1232 0.0848 0.1006
-1.4 0.1948 0.1968 0.1401 0.1581
-1.5 0.2426 0.2451 0.1888 0.2102
-1.7 0.3182 0.3196 0.2675 0.2898
-1.9 0.3709 0.3733 0.3261 0.3457
-2.0 0.3921 0.3939 0.3467 0.3678
-2.5 0.4717 0.4742 0.4327 0.4564
-3.0 0.5227 0.5253 0.4910 0.5124
-Case 2a Moderate correlation No ZI 1.1 0.0061 0.0060 0.0012 0.0026
-1.2 0.2096 0.2170 0.0526 0.0674
-1.3 0.3304 0.3377 0.1334 0.1530
-1.4 0.3991 0.4036 0.1998 0.2210
-1.5 0.4489 0.4537 0.2499 0.2751
-1.7 0.5172 0.5197 0.3359 0.3637
-1.9 0.5663 0.5681 0.4043 0.4323
-2.0 0.5837 0.5847 0.4310 0.4589
-2.5 0.6396 0.6411 0.5363 0.5608
-3.0 0.6734 0.6746 0.5984 0.6214
-Case 2b Moderate correlation ZI 1.1 0.0067 0.0072 0.0011 0.0025
-1.2 0.2067 0.2133 0.0523 0.0659
-1.3 0.3309 0.3367 0.1348 0.1535
-1.4 0.3968 0.4021 0.1973 0.2186
-1.5 0.4438 0.4477 0.2475 0.2723
-1.7 0.5138 0.5157 0.3340 0.3599
-1.9 0.5601 0.5620 0.4000 0.4274
-2.0 0.5772 0.5790 0.4297 0.4570
-2.5 0.6329 0.6339 0.5313 0.5561
-3.0 0.6696 0.6708 0.5967 0.6205
-Case 3a Strong correlation No ZI 1.1 0.0065 0.0076 0.0009 0.0020
-1.2 0.2680 0.2784 0.0329 0.0455
-1.3 0.4595 0.4682 0.1088 0.1313
-1.4 0.5675 0.5746 0.1807 0.2100
-1.5 0.6272 0.6330 0.2480 0.2791
-1.7 0.7015 0.7049 0.3546 0.3868
-1.9 0.7460 0.7482 0.4300 0.4612
-2.0 0.7629 0.7644 0.4608 0.4940
-2.5 0.8096 0.8106 0.5793 0.6094
-3.0 0.8330 0.8336 0.6530 0.6799
-Case 3b Strong correlation ZI 1.1 0.0083 0.0094 0.0010 0.0021
-1.2 0.2627 0.2769 0.0338 0.0471
-1.3 0.4588 0.4688 0.1098 0.1326
-1.4 0.5638 0.5717 0.1826 0.2115
-1.5 0.6275 0.6333 0.2492 0.2801
-1.7 0.7044 0.7080 0.3534 0.3848
-1.9 0.7485 0.7507 0.4276 0.4599
-2.0 0.7630 0.7651 0.4574 0.4893
-2.5 0.8100 0.8109 0.5787 0.6097
-3.0 0.8341 0.8348 0.6556 0.6825
+Case $\lambda$
+DP Poisson
+local-global
+DP ZIP
+local-global
+DP Hu et al.
+DP Hu et al.
+posterior quantile optim
+Case 1a
+Weak correlation
+No ZI
+1.1 0.0005 0.0005 0.0001 0.0003
+1.2 0.0296 0.0292 0.0166 0.0201
+1.3 0.1080 0.1080 0.0548 0.0629
+1.4 0.1678 0.1678 0.0979 0.1087
+1.5 0.2143 0.2147 0.1409 0.1533
+1.7 0.2854 0.2851 0.2130 0.2274
+1.9 0.3404 0.3407 0.2738 0.2895
+2.0 0.3647 0.3648 0.3002 0.3164
+2.5 0.4584 0.4588 0.4061 0.4231
+3.0 0.5245 0.5252 0.4818 0.4988
+Case 1b
+Weak correlation
+30% ZI
+1.1 0.0006 0.0004 0.0001 0.0003
+1.2 0.0296 0.0291 0.0167 0.0203
+1.3 0.1090 0.1086 0.0545 0.0625
+1.4 0.1694 0.1700 0.0995 0.1102
+1.5 0.2170 0.2165 0.1423 0.1546
+1.7 0.2847 0.2849 0.2129 0.2269
+1.9 0.3399 0.3396 0.2728 0.2883
+2.0 0.3639 0.3639 0.2999 0.3152
+2.5 0.4571 0.4575 0.4048 0.4211
+3.0 0.5256 0.5261 0.4813 0.4998
+Case 2a
+Moderate correlation
+No ZI
+1.1 0.0027 0.0029 0.0003 0.0006
+1.2 0.2210 0.2219 0.0455 0.0538
+1.3 0.3322 0.3325 0.1104 0.1221
+1.4 0.3989 0.3991 0.1677 0.1821
+1.5 0.4410 0.4410 0.2193 0.2354
+1.7 0.5011 0.5014 0.3000 0.3158
+1.9 0.5415 0.5417 0.3638 0.3810
+2.0 0.5572 0.5571 0.3905 0.4078
+2.5 0.6167 0.6172 0.4976 0.5147
+3.0 0.6597 0.6599 0.5667 0.5826
+Case 2b
+Moderate correlation
+30% ZI
+1.1 0.0028 0.0025 0.0004 0.0007
+1.2 0.2187 0.2202 0.0454 0.0543
+1.3 0.3323 0.3329 0.1098 0.1215
+1.4 0.3977 0.3977 0.1666 0.1809
+1.5 0.4411 0.4412 0.2202 0.2350
+1.7 0.5002 0.5002 0.2999 0.3158
+1.9 0.5408 0.5408 0.3632 0.3801
+2.0 0.5571 0.5572 0.3914 0.4085
+2.5 0.6177 0.6180 0.4974 0.5141
+3.0 0.6611 0.6613 0.5674 0.5835
+Case 3a
+Strong correlation
+No ZI
+1.1 0.0053 0.0055 0.0004 0.0009
+1.2 0.3080 0.3092 0.0480 0.0583
+1.3 0.4558 0.4565 0.1219 0.1360
+1.4 0.5480 0.5486 0.1875 0.2042
+1.5 0.6090 0.6094 0.2466 0.2647
+1.7 0.6805 0.6806 0.3412 0.3601
+1.9 0.7177 0.7176 0.4124 0.4316
+2.0 0.7290 0.7290 0.4449 0.4636
+2.5 0.7675 0.7676 0.5590 0.5768
+3.0 0.7959 0.7961 0.6307 0.6470
+Case 3b
+Strong correlation
+30% ZI
+1.1 0.0039 0.0044 0.0004 0.0007
+1.2 0.3061 0.3076 0.0492 0.0593
+1.3 0.4517 0.4530 0.1213 0.1356
+1.4 0.5458 0.5462 0.1876 0.2041
+1.5 0.6058 0.6061 0.2461 0.2644
+1.7 0.6804 0.6805 0.3397 0.3588
+1.9 0.7149 0.7150 0.4131 0.4330
+2.0 0.7275 0.7276 0.4441 0.4636
+2.5 0.7669 0.7670 0.5584 0.5761
+3.0 0.7960 0.7962 0.6299 0.6462
 Table S.4: Simulation of average type I error for DP-based models
-Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
-Case 1a Weak correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0000
-1.2 0.0001 0.0001 0.0001 0.0002
-1.3 0.0002 0.0003 0.0001 0.0003
-1.4 0.0002 0.0002 0.0001 0.0004
-1.5 0.0003 0.0003 0.0002 0.0005
-1.7 0.0002 0.0003 0.0002 0.0005
-1.9 0.0002 0.0002 0.0002 0.0005
-2.0 0.0003 0.0003 0.0002 0.0005
-2.5 0.0003 0.0003 0.0002 0.0006
-3.0 0.0002 0.0002 0.0002 0.0006
-Case 1b Weak correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
-1.2 0.0001 0.0001 0.0001 0.0002
-1.3 0.0002 0.0002 0.0001 0.0002
-1.4 0.0003 0.0003 0.0001 0.0004
-1.5 0.0003 0.0004 0.0002 0.0004
-1.7 0.0003 0.0003 0.0002 0.0005
-1.9 0.0003 0.0003 0.0002 0.0005
-2.0 0.0002 0.0002 0.0002 0.0006
-2.5 0.0002 0.0002 0.0002 0.0005
-3.0 0.0003 0.0003 0.0003 0.0006
-Case 2a Moderate correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0001
-1.2 0.0005 0.0007 0.0001 0.0002
-1.3 0.0006 0.0008 0.0002 0.0004
-1.4 0.0008 0.0009 0.0002 0.0005
-1.5 0.0006 0.0007 0.0002 0.0006
-1.7 0.0005 0.0006 0.0003 0.0007
-1.9 0.0004 0.0004 0.0003 0.0007
-2.0 0.0003 0.0004 0.0003 0.0008
-2.5 0.0003 0.0003 0.0003 0.0007
-3.0 0.0003 0.0003 0.0002 0.0007
-Case 2b Moderate correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
-1.2 0.0005 0.0006 0.0001 0.0002
-1.3 0.0007 0.0007 0.0001 0.0004
-1.4 0.0006 0.0007 0.0002 0.0005
-1.5 0.0004 0.0005 0.0002 0.0006
-1.7 0.0004 0.0004 0.0003 0.0006
-1.9 0.0003 0.0003 0.0003 0.0007
-2.0 0.0004 0.0004 0.0003 0.0007
-2.5 0.0003 0.0003 0.0003 0.0007
-3.0 0.0003 0.0003 0.0003 0.0008
-Case 3a Strong correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0001
-1.2 0.0005 0.0007 0.0001 0.0002
-1.3 0.0006 0.0007 0.0001 0.0004
-1.4 0.0009 0.0010 0.0002 0.0005
-1.5 0.0007 0.0007 0.0003 0.0007
-1.7 0.0005 0.0006 0.0004 0.0008
-1.9 0.0005 0.0006 0.0003 0.0009
-2.0 0.0005 0.0005 0.0003 0.0009
-2.5 0.0004 0.0004 0.0003 0.0009
-3.0 0.0004 0.0004 0.0003 0.0009
-Case 3b Strong correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
-1.2 0.0006 0.0007 0.0001 0.0002
-1.3 0.0007 0.0008 0.0001 0.0004
-1.4 0.0008 0.0009 0.0002 0.0005
-1.5 0.0007 0.0008 0.0002 0.0006
-1.7 0.0005 0.0006 0.0003 0.0007
-1.9 0.0004 0.0004 0.0003 0.0008
-2.0 0.0004 0.0005 0.0004 0.0009
-2.5 0.0004 0.0004 0.0004 0.0009
-3.0 0.0005 0.0005 0.0004 0.0010
+Case $\lambda$
+DP Poisson
+local-global
+DP ZIP
+local-global
+DP Hu et al.
+DP Hu et al.
+posterior quantile optim
+Case 1a
+Weak correlation
+No ZI
+1.1 0.0000 0.0000 0.0000 0.0000
+1.2 0.0000 0.0000 0.0000 0.0000
+1.3 0.0000 0.0000 0.0000 0.0000
+1.4 0.0000 0.0000 0.0000 0.0000
+1.5 0.0000 0.0000 0.0000 0.0000
+1.7 0.0000 0.0000 0.0000 0.0000
+1.9 0.0000 0.0000 0.0000 0.0000
+2.0 0.0000 0.0000 0.0000 0.0000
+2.5 0.0000 0.0000 0.0000 0.0000
+3.0 0.0000 0.0000 0.0000 0.0000
+Case 1b
+Weak correlation
+30% ZI
+1.1 0.0000 0.0000 0.0000 0.0000
+1.2 0.0000 0.0000 0.0000 0.0000
+1.3 0.0000 0.0000 0.0000 0.0000
+1.4 0.0000 0.0000 0.0000 0.0000
+1.5 0.0000 0.0000 0.0000 0.0000
+1.7 0.0000 0.0000 0.0000 0.0000
+1.9 0.0000 0.0000 0.0000 0.0000
+2.0 0.0000 0.0000 0.0000 0.0000
+2.5 0.0000 0.0000 0.0000 0.0000
+3.0 0.0000 0.0000 0.0000 0.0000
+Case 2a
+Moderate correlation
+No ZI
+1.1 0.0000 0.0000 0.0000 0.0000
+1.2 0.0000 0.0000 0.0000 0.0000
+1.3 0.0000 0.0000 0.0000 0.0000
+1.4 0.0000 0.0000 0.0000 0.0000
+1.5 0.0000 0.0000 0.0000 0.0000
+1.7 0.0000 0.0000 0.0000 0.0000
+1.9 0.0000 0.0000 0.0000 0.0000
+2.0 0.0000 0.0000 0.0000 0.0000
+2.5 0.0000 0.0000 0.0000 0.0000
+3.0 0.0000 0.0000 0.0000 0.0000
+Case 2b
+Moderate correlation
+30% ZI
+1.1 0.0000 0.0000 0.0000 0.0000
+1.2 0.0000 0.0000 0.0000 0.0000
+1.3 0.0000 0.0000 0.0000 0.0000
+1.4 0.0000 0.0000 0.0000 0.0000
+1.5 0.0000 0.0000 0.0000 0.0000
+1.7 0.0000 0.0000 0.0000 0.0000
+1.9 0.0000 0.0000 0.0000 0.0000
+2.0 0.0000 0.0000 0.0000 0.0000
+2.5 0.0000 0.0000 0.0000 0.0000
+3.0 0.0000 0.0000 0.0000 0.0000
+Case 3a
+Strong correlation
+No ZI
+1.1 0.0000 0.0000 0.0000 0.0000
+1.2 0.0001 0.0001 0.0000 0.0000
+1.3 0.0001 0.0001 0.0000 0.0000
+1.4 0.0001 0.0001 0.0000 0.0000
+1.5 0.0000 0.0000 0.0000 0.0000
+1.7 0.0000 0.0000 0.0000 0.0000
+1.9 0.0000 0.0000 0.0000 0.0000
+2.0 0.0000 0.0000 0.0000 0.0000
+2.5 0.0000 0.0000 0.0000 0.0000
+3.0 0.0000 0.0000 0.0000 0.0000
+Case 3b
+Strong correlation
+30% ZI
+1.1 0.0000 0.0000 0.0000 0.0000
+1.2 0.0000 0.0000 0.0000 0.0000
+1.3 0.0001 0.0001 0.0000 0.0000
+1.4 0.0001 0.0001 0.0000 0.0000
+1.5 0.0000 0.0000 0.0000 0.0000
+1.7 0.0000 0.0000 0.0000 0.0000
+1.9 0.0000 0.0000 0.0000 0.0000
+2.0 0.0000 0.0000 0.0000 0.0000
+2.5 0.0000 0.0000 0.0000 0.0000
+3.0 0.0000 0.0000 0.0000 0.0000
 Table S.5: Simulation of F-score for DP-based models
-Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
-Case 1a Weak correlation No ZI 1.1 0.0030 0.0025 0.0018 0.0040
-1.2 0.0783 0.0772 0.0650 0.0810
-1.3 0.2129 0.2152 0.1615 0.1844
-1.4 0.3180 0.3204 0.2466 0.2722
-1.5 0.3887 0.3924 0.3175 0.3444
-1.7 0.4761 0.4791 0.4200 0.4446
-1.9 0.5358 0.5375 0.4883 0.5115
-2.0 0.5602 0.5634 0.5148 0.5380
-2.5 0.6369 0.6390 0.6027 0.6213
-3.0 0.6859 0.6879 0.6566 0.6738
-Case 1b Weak correlation ZI 1.1 0.0023 0.0022 0.0022 0.0042
-1.2 0.0748 0.0733 0.0622 0.0774
-1.3 0.2109 0.2130 0.1531 0.1787
-1.4 0.3212 0.3240 0.2415 0.2682
-1.5 0.3863 0.3894 0.3138 0.3429
-1.7 0.4793 0.4808 0.4187 0.4449
-1.9 0.5378 0.5403 0.4886 0.5094
-2.0 0.5602 0.5621 0.5118 0.5333
-2.5 0.6382 0.6406 0.6011 0.6226
-3.0 0.6837 0.6859 0.6556 0.6731
-Case 2a Moderate correlation No ZI 1.1 0.0104 0.0102 0.0024 0.0050
-1.2 0.3333 0.3429 0.0973 0.1231
-1.3 0.4894 0.4973 0.2322 0.2616
-1.4 0.5642 0.5684 0.3299 0.3578
-1.5 0.6141 0.6186 0.3966 0.4269
-1.7 0.6775 0.6796 0.4997 0.5287
-1.9 0.7198 0.7212 0.5724 0.5988
-2.0 0.7344 0.7351 0.5992 0.6242
-2.5 0.7777 0.7789 0.6954 0.7141
-3.0 0.8025 0.8034 0.7463 0.7622
-Case 2b Moderate correlation ZI 1.1 0.0114 0.0122 0.0021 0.0049
-1.2 0.3292 0.3375 0.0969 0.1204
-1.3 0.4900 0.4964 0.2345 0.2626
-1.4 0.5624 0.5677 0.3265 0.3548
-1.5 0.6101 0.6136 0.3938 0.4237
-1.7 0.6750 0.6766 0.4975 0.5248
-1.9 0.7150 0.7166 0.5681 0.5941
-2.0 0.7289 0.7303 0.5980 0.6226
-2.5 0.7727 0.7735 0.6911 0.7102
-3.0 0.7999 0.8007 0.7447 0.7614
-Case 3a Strong correlation No ZI 1.1 0.0114 0.0131 0.0018 0.0039
-1.2 0.4044 0.4167 0.0623 0.0852
-1.3 0.6199 0.6280 0.1940 0.2290
-1.4 0.7176 0.7231 0.3035 0.3437
-1.5 0.7660 0.7701 0.3947 0.4325
-1.7 0.8212 0.8234 0.5209 0.5537
-1.9 0.8515 0.8527 0.5988 0.6270
-2.0 0.8629 0.8638 0.6284 0.6571
-2.5 0.8929 0.8934 0.7314 0.7531
-3.0 0.9072 0.9076 0.7881 0.8055
-Case 3b Strong correlation ZI 1.1 0.0145 0.0162 0.0020 0.0040
-1.2 0.3987 0.4159 0.0639 0.0881
-1.3 0.6192 0.6287 0.1955 0.2312
-1.4 0.7147 0.7211 0.3062 0.3455
-1.5 0.7658 0.7702 0.3964 0.4342
-1.7 0.8231 0.8255 0.5197 0.5520
-1.9 0.8537 0.8551 0.5966 0.6261
-2.0 0.8633 0.8645 0.6251 0.6529
-2.5 0.8931 0.8936 0.7308 0.7536
-3.0 0.9076 0.9080 0.7897 0.8071
+Case $\lambda$
+DP Poisson
+local-global
+DP ZIP
+local-global
+DP Hu et al.
+DP Hu et al.
+posterior quantile optim
+Case 1a
+Weak correlation
+No ZI
+1.1 0.0009 0.0009 0.0003 0.0006
+1.2 0.0564 0.0556 0.0323 0.0390
+1.3 0.1927 0.1927 0.1032 0.1173
+1.4 0.2857 0.2858 0.1773 0.1948
+1.5 0.3513 0.3519 0.2459 0.2644
+1.7 0.4423 0.4420 0.3498 0.3688
+1.9 0.5062 0.5065 0.4286 0.4472
+2.0 0.5328 0.5329 0.4605 0.4788
+2.5 0.6269 0.6272 0.5762 0.5925
+3.0 0.6864 0.6869 0.6489 0.6635
+Case 1b
+Weak correlation
+30% ZI
+1.1 0.0011 0.0009 0.0003 0.0007
+1.2 0.0564 0.0554 0.0326 0.0394
+1.3 0.1946 0.1939 0.1025 0.1168
+1.4 0.2881 0.2890 0.1799 0.1974
+1.5 0.3551 0.3544 0.2481 0.2665
+1.7 0.4416 0.4418 0.3498 0.3681
+1.9 0.5056 0.5053 0.4273 0.4456
+2.0 0.5319 0.5319 0.4600 0.4773
+2.5 0.6257 0.6261 0.5749 0.5907
+3.0 0.6875 0.6879 0.6486 0.6646
+Case 2a
+Moderate correlation
+No ZI
+1.1 0.0049 0.0052 0.0006 0.0011
+1.2 0.3585 0.3596 0.0864 0.1014
+1.3 0.4960 0.4964 0.1981 0.2166
+1.4 0.5683 0.5684 0.2863 0.3067
+1.5 0.6099 0.6099 0.3586 0.3795
+1.7 0.6661 0.6662 0.4605 0.4784
+1.9 0.7010 0.7012 0.5324 0.5498
+2.0 0.7144 0.7143 0.5606 0.5774
+2.5 0.7616 0.7620 0.6633 0.6775
+3.0 0.7938 0.7939 0.7223 0.7342
+Case 2b
+Moderate correlation
+30% ZI
+1.1 0.0052 0.0047 0.0008 0.0014
+1.2 0.3552 0.3571 0.0863 0.1024
+1.3 0.4959 0.4964 0.1972 0.2157
+1.4 0.5672 0.5672 0.2848 0.3050
+1.5 0.6103 0.6104 0.3600 0.3791
+1.7 0.6652 0.6652 0.4604 0.4783
+1.9 0.7006 0.7007 0.5317 0.5490
+2.0 0.7144 0.7144 0.5614 0.5781
+2.5 0.7626 0.7627 0.6631 0.6770
+3.0 0.7948 0.7949 0.7227 0.7347
+Case 3a
+Strong correlation
+No ZI
+1.1 0.0094 0.0097 0.0009 0.0017
+1.2 0.4665 0.4678 0.0911 0.1095
+1.3 0.6224 0.6230 0.2167 0.2386
+1.4 0.7050 0.7055 0.3150 0.3378
+1.5 0.7545 0.7547 0.3947 0.4171
+1.7 0.8081 0.8081 0.5078 0.5276
+1.9 0.8339 0.8339 0.5830 0.6010
+2.0 0.8420 0.8420 0.6149 0.6316
+2.5 0.8674 0.8675 0.7160 0.7295
+3.0 0.8856 0.8856 0.7725 0.7836
+Case 3b
+Strong correlation
+30% ZI
+1.1 0.0068 0.0077 0.0008 0.0014
+1.2 0.4644 0.4661 0.0934 0.1113
+1.3 0.6186 0.6198 0.2158 0.2380
+1.4 0.7031 0.7034 0.3151 0.3378
+1.5 0.7519 0.7521 0.3942 0.4167
+1.7 0.8079 0.8080 0.5062 0.5262
+1.9 0.8324 0.8324 0.5836 0.6023
+2.0 0.8409 0.8410 0.6140 0.6315
+2.5 0.8670 0.8671 0.7156 0.7290
+3.0 0.8857 0.8858 0.7718 0.7829

### v3

@@
 While clinical trials remain the gold standard for assessing medical product safety before approval, they have limitations due to factors such as small sample sizes, ethical considerations, restricted timelines, etc., and the critical need for postmarket monitoring is increasingly recognized.
-For example, severe AEs such as congenital malformations from Thalidomide and liver injury from the Ketoconazole lavertu2021new were identified only after these drugs were approved for the U.S.
+For example, severe AEs such as liver injury from Ketoconazole Lavertu et al.
+(2021) were identified only after these drugs were approved for the U.S.
 market.
+Outside the U.S., congenital malformations caused by thalidomide were identified only after their introduction in European and other markets.
 To complement clinical trial findings with real-world evidence (RWE), continuous post-marketing surveillance—a cornerstone of contemporary pharmacovigilance—has proven essential.
@@
 For example, the FAERS database has compiled over $29$ million reports by 2024.
-These databases are inherently observational, and the data curated in them suffer from several statistical challenges, including selection bias, measurement error, confounding, missing data, and absence of information on non-events, non-users, and general health characteristics of users, among other factors markatou2014pattern , precluding formal causal inference.
+These databases are inherently observational, and the data curated in them suffer from several statistical challenges, including selection bias, measurement error, confounding, missing data, and absence of information on non-events, non-users, and general health characteristics of users, among other factors Markatou and Ball (2014), precluding formal causal inference.
 In SRS data mining, the focus instead is on detecting signals— identifying drug-AE pairs whose observed counts are disproportionately higher than their null expected counts, the theoretical baseline counts for the pairs under the assumption that the drug and the AE are independent.
@@
 Instead, SRS data mining methods rely on statistical comparators (all other reports in the database not involving the drug or AE in question) to approximate marginal total report counts for each drug and AE.
-These approximated margins are then used to estimate the null expected counts for the drug-AE pairs gravel2023considerations ; kontsioti2023exploring ; zhang2018three ; markatou2014pattern .
-Despite these limitations, meaningful safety signals can be detected from these databases when appropriate statistical methods are applied, as demonstrated by a growing body of research.
+These approximated margins are then used to estimate the null expected counts for the drug-AE pairs.
+Despite these limitations, meaningful safety signals can be detected from these databases when appropriate statistical methods are applied, as demonstrated by a growing body of research Gravel and Douros (2023); Kontsioti et al.
+(2023); Zhang et al.
+(2018); Markatou and Ball (2014).
 For principled and accurate signal detection from SRS databases, numerous statistical data mining methods have been developed over the past decades.
-Examples include the proportional reporting ratio (PRR) evans2001use , the reporting odds ratio (ROR) rothman2004reporting , formal frequentist methods such as the likelihood ratio test (LRT) and its variants huang2011likelihood ; chakraborty2022use ; ding2020evaluation , and Bayesian methods such as the Gamma Poisson Shrinker (GPS) dumouchel1999bayesian , Bayesian confidence propagation neural network (BCPNN) bate1998bayesian , and Dirichlet process (DP) Poisson mixture model Hu et al.hu2015signal .
+Examples include the proportional reporting ratio (PRR) Evans et al.
+(2001), the reporting odds ratio (ROR) Rothman et al.
+(2004), formal frequentist methods such as the likelihood ratio test (LRT) and its variants Huang et al.
+(2011); Chakraborty et al.
+(2022); Ding et al.
+(2020), and Bayesian methods such as the Gamma Poisson Shrinker (GPS) DuMouchel (1999), Bayesian confidence propagation neural network (BCPNN) Bate et al.
+(1998), and Dirichlet process (DP) Poisson mixture model Hu et al.Hu et al.
+(2015).
 Some approaches—e.g., PRR and ROR—employ ad-hoc thresholds on empirical signal strength estimates, while the more formal methods—e.g., LRT and Bayesian approaches—employ rigorous probabilistic tools for statistically guaranteed signal detection, including controlled type I errors and false discovery rates (FDR).
 Despite these advances, signal detection for rare AEs and newer drugs remains challenging due to their small (marginal total) sample sizes, which leads to small observed and expected report counts in individual drugs, reducing the statistical power to detect the underlying signals.
-Yet, a large and growing body of researchpoluzzi2012data ; sardella2018pharmacovigilance ; raschi2021value has documented the critical need for reliable early signal detection precisely under these settings involving rare AEs and newer drugs.
+Yet, a large and growing body of researchPoluzzi et al.
+(2012); Sardella and Belcher (2018); Raschi et al.
+(2021) has documented the critical need for reliable early signal detection precisely under these settings involving rare AEs and newer drugs.
 In this paper, we argue that an appropriately formulated Bayesian framework can yield important insights for pharmacovigilance under these challenging settings.
@@
 Prior specification, however, is crucial for optimal performance.
-An overly restrictive prior—e.g., a single, common gamma prior for all drug-AE signal strength parameters—may cause excessive shrinkage, pulling estimates toward a single common valuehu2015signal .
+An overly restrictive prior—e.g., a single, common gamma prior for all drug-AE signal strength parameters—may cause excessive shrinkage, pulling estimates toward a single common valueHu et al.
+(2015).
 Conversely, an overly flexible prior—e.g., independent flat priors for all parameters—may inadequately regularize their estimation, resulting in overfitting and noisy estimates.
-State-of-the-art Bayesian data mining approaches balance flexibility and regularization using semi/nonparametric shrinkage priors, such as a two-component gamma mixture prior (the GPS modeldumouchel1999bayesian ) or a DP mixture prior with a base gamma distributionhu2015signal , for AE signal strengths separately for each drug.
+State-of-the-art Bayesian data mining approaches balance flexibility and regularization using semi/nonparametric shrinkage priors, such as a two-component gamma mixture prior (the GPS modelDuMouchel (1999)) or a DP mixture prior with a base gamma distributionHu et al.
+(2015), for AE signal strengths separately for each drug.
 These approaches stabilize signal detection; however, they still impose restrictive assumptions, treating drugs in question either entirely independently or as a single homogeneous group.
 Real-world SRS databases, however, record drugs with nuanced similarities in their AEs.
-For instance, various statin drugs, used to lower blood cholesterol and cardiovascular risk, share chemical compositions and function, leading to common AEs like rhabdomyolysis montastruc2023rhabdomyolysis .
-Yet, due to small differences in drug composition and function they may cause distinct AEs or have variations in AE severity; e.g., myoglobinuria is typically associated with Atorvastatin but not Lovastatin seneff2016statins .
+For instance, various statin drugs, used to lower blood cholesterol and cardiovascular risk, share chemical compositions and function, leading to common AEs like rhabdomyolysis Montastruc (2023).
+Yet, due to small differences in drug composition and function they may cause distinct AEs or have variations in AE severity; e.g., myoglobinuria is typically associated with Atorvastatin but not Lovastatin Seneff (2016).
 Figure 1 shows the pairwise Kendall’s $\tau$ correlation coefficients between six statin drugs with respect to the occurrence counts of 1491 AEs as cataloged in the FAERS database (Q3 2014-Q4 2024; see Section 5 for details on the dataset).
-These correlations reveal strong yet nuanced between-drug associations: they are consistently above $0.40$ with some as high as $0.68$, ruling out independence, yet many are noticeably around or below $0.50$, suggesting against a single grouping of all drugs.
+These correlations reveal strong yet nuanced between-drug associations: they are consistently above $0.40$ with some as high as $0.68$, ruling out independence, yet many are noticeably around or below $0.50$, arguing against treating all statins as a single homogeneous group.
+These empirical findings illustrate a key challenge in pharmacovigilance: drugs often exhibit correlated but non-identical AE profiles, making assumptions of either independence or identicalness inappropriate and motivating flexible modeling of between-drug dependencies in SRS data mining.
 Figure 1: Kendall’s $\tau$ correlation matrix for occurrence counts of 1491 commonly occurring AEs (PTs) among six statin drugs in the FDA FAERS data (2014 Q1-2020 Q4).
-Existing SRS data mining approaches, both Bayesian and frequentist, rarely account for such nuanced between-drug associations.
-Instead, they often assume either independence of drugs (the original LRT, BCPNN, GPS, and DP Poisson-Gamma) or a single group for all drugs (extended LRT of Huang et al.huang2011likelihood ).
+There are several prominent case studies that reinforce this challenge.
+For example, the statin cerivastatin was withdrawn from the U.S.
+market Staffa et al.
+(2002) within four years of its introduction due to a heightened risk of fatal rhabdomyolysis; this early detection was facilitated by prior knowledge that rhabdomyolysis was a recognized AE of atorvastatin and simvastatin, which had been marketed earlier.
+Conversely, increased cardiovascular risk first became evident with the COX-2 inhibitor rofecoxib, which was consequently removed from the marketTopol (2004).
+Similar concerns subsequently emerged for the related COX-2 inhibitor celecoxib after the withdrawal of rofecoxibSolomon et al.
+(2005), until a large 10-year randomized trial (PRECISION) enrolling approximately 24,000 patients ultimately established the cardiovascular risk of celecoxib to be comparable to that of traditional nonsteroidal anti-inflammatory drugs (NSAIDs) such as ibuprofen and naproxenNissen et al.
+(2016).
+Together, these examples illustrate that safety signals frequently arise at the drug-class level owing to between-drug similarities, while also highlighting that these associations are typically partial rather than identical across drug-specific AE profiles.
+Flexible acknowledgment of these partial associations is therefore necessary for effective signal detection.
+Yet, existing SRS data mining approaches, both Bayesian and frequentist, rarely account for such nuanced between-drug associations.
+Instead, they often assume either independence of drugs (the original LRT, BCPNN, GPS, and DP Poisson-Gamma) or a single group for all drugs (extended LRT of Huang et al.Huang et al.
+(2011)).
 These restrictive assumptions limit the statistical power and sensitivity for detecting signals, particularly for rare AEs and newer drugs with small observed and expected counts.
@@
 First, we propose local-global mixture DP models for SRS signal detection (Section 4.2).
-Local-global priors are widely used in contemporary statistics (see Bhadra et al.bhadra2019lasso and references cited therein) and have been employed in DP modeling of grouped data, most prominently through hierarchical DP models teh2004sharing .
-Such models require handling complex combinatorial partition structures—often represented via Chinese restaurant franchise-style distributions teh2004sharing —making implementation challenging, particularly for sizable datasets.
-In contrast, our framework adopts a more transparent two-component mixture of DPs antoniak1974mixtures for local-global pattern decomposition.
+Local-global priors are widely used in contemporary statistics (see Bhadra et al.Bhadra et al.
+(2019) and references cited therein) and have been employed in DP modeling of grouped data, most prominently through hierarchical DP models Teh et al.
+(2004).
+Such models require handling complex combinatorial partition structures—often represented via Chinese restaurant franchise-style distributions Teh et al.
+(2004)—making implementation challenging, particularly for sizable datasets.
+In contrast, our framework adopts a more transparent two-component mixture of DPs Antoniak (1974) for local-global pattern decomposition.
 This enables our framework to achieve efficient MCMC implementation using standard distributions while retaining the ability to capture both shared global structure and drug-specific local variations.
@@
 Third, through extensive simulations (Section 5), we show that our methods achieve comparable or superior sensitivity to existing approaches (reviewed in Section 3) when only a few strong signals are present across common AEs but yield substantial sensitivity gains—often twofold or more—when many weak signals exist among rare AEs and newer drugs, all while maintaining strong FDR and type I error control.
-Finally, our analysis of a real-world statin dataset (Section 6) from the FDA FAERS database demonstrates our methods’ ability to effectively detect AE signals in practice from large-scale SRS databases.
+Finally, our analysis of a real-world statin dataset (Section 6) from the FDA FAERS database demonstrates our method’s ability to effectively detect AE signals in practice from large-scale SRS databases, particularly for rarer AEs and newer drugs, that are often missed by existing approaches.
 2 SRS DATA, OUR NOTATION, AND PROBABILITY MODELS
@@
 This category of “Other drugs” is conceptually too heterogeneous to have any meaningful, unique AEs, but is still included to aid better comparison for disproportionality analysis, and a greater AE (row) total ($n_{i\bullet}$) counts for more accurate $E_{ij}$ estimation.
-Analogously, a heterogeneous group of ’Other AEs’, obtained by collapsing all AEs outside those of interest (or with very small sample sizes), may be constructed ding2020evaluation .
+Analogously, a heterogeneous group of ’Other AEs’, obtained by collapsing all AEs outside those of interest (or with very small sample sizes), may be constructed Ding et al.
+(2020).
 Signal detection in SRS contingency tables typically relies on the observed relative reporting rates, $n_{ij}/E_{ij}$, or a related quantity, as a measure of signal strength.
@@
 3.1 The Bayesian Propagation Neural Network (BCPNN) method
-The BCPNN approach (Bate et al.bate1998bayesian ) assumes that conditional on the grand total count $n_{\bullet\bullet}$, the cell-specific counts $n_{ij}$, the marginal row totals $n_{i\bullet}$, and the marginal column totals $n_{\bullet j}$ each distributed as binomial-beta: (a) $n_{ij}\mid p_{ij}\sim\text{Binomial}(n_{\bullet\bullet},p_{ij}),\ p_{ij}\sim\text{Beta}(\alpha_{ij},\beta_{ij})$, (b) $n_{i\bullet}\mid p_{i\bullet}\sim\text{Binomial}(n_{\bullet\bullet},p_{i\bullet}),\allowbreak\ p_{i\bullet}\sim\text{Beta}(\alpha_{i\bullet},\beta_{i\bullet})$, (c) $n_{\bullet j}\mid p_{\bullet j}\sim\text{Binomial}(n_{\bullet\bullet},p_{\bullet j}),\allowbreak\ p_{\bullet j}\sim\text{Beta}(\alpha_{\bullet j},\beta_{\bullet j})$, where $p_{ij}$ is the cell-specific reporting rate, and $p_{i\bullet}$ and $p_{\bullet j}$ are the marginal row and column reporting rates.
+The BCPNN approach (Bate et al.Bate et al.
+(1998)) assumes that conditional on the grand total count $n_{\bullet\bullet}$, the cell-specific counts $n_{ij}$, the marginal row totals $n_{i\bullet}$, and the marginal column totals $n_{\bullet j}$ each distributed as binomial-beta: (a) $n_{ij}\mid p_{ij}\sim\text{Binomial}(n_{\bullet\bullet},p_{ij}),\ p_{ij}\sim\text{Beta}(\alpha_{ij},\beta_{ij})$, (b) $n_{i\bullet}\mid p_{i\bullet}\sim\text{Binomial}(n_{\bullet\bullet},p_{i\bullet}),\allowbreak\ p_{i\bullet}\sim\text{Beta}(\alpha_{i\bullet},\beta_{i\bullet})$, (c) $n_{\bullet j}\mid p_{\bullet j}\sim\text{Binomial}(n_{\bullet\bullet},p_{\bullet j}),\allowbreak\ p_{\bullet j}\sim\text{Beta}(\alpha_{\bullet j},\beta_{\bullet j})$, where $p_{ij}$ is the cell-specific reporting rate, and $p_{i\bullet}$ and $p_{\bullet j}$ are the marginal row and column reporting rates.
 We note that the binomial model assumptions for $n_{ij}$, $n_{i\bullet}$ and $n_{\bullet j}$ can be understood as consequences of an independent Poisson model for $n_{ij}$ with certain choices of $p_{ij}$, $p_{i\bullet}$, and $p_{\bullet j}$.
-Bate et al.bate1998bayesian consider the information content $\text{IC}_{ij}=\log_{2}\left[p_{ij}/(p_{i\bullet}p_{\bullet j})\right]$ as a measure of signal strength for the drug-$j$ AE-$i$ pair, with a higher IC indicating a stronger association.
+Bate et al.Bate et al.
+(1998) consider the information content $\text{IC}_{ij}=\log_{2}\left[p_{ij}/(p_{i\bullet}p_{\bullet j})\right]$ as a measure of signal strength for the drug-$j$ AE-$i$ pair, with a higher IC indicating a stronger association.
 The authors suggest an empirical Bayes estimation of these quantities using $\hat{\alpha}_{ij}=\hat{\alpha}_{i\bullet}=\hat{\beta}_{i\bullet}=\hat{\alpha}_{\bullet j}=\hat{\beta}_{\bullet j}=1$ and $\hat{\beta}_{ij}=1/(E(p_{i\bullet}\mid n_{i\bullet})E(p_{\bullet j}\mid n_{\bullet j}))-1$, with the denominator of $\hat{\beta}_{ij}$ being computed using the posterior distributions $p_{i\bullet}\mid n_{i\bullet}\sim\text{Beta}(\hat{\alpha}_{i\bullet}+n_{i\bullet},\hat{\beta}_{i\bullet}+n_{\bullet\bullet}-n_{i\bullet})$ and $p_{\bullet j}\mid n_{\bullet j}\sim\text{Beta}(\hat{\alpha}_{\bullet j}+n_{\bullet j},\hat{\beta}_{\bullet j}+n_{\bullet\bullet}-n_{\bullet j})$.
@@
 A signal is detected at the drug-$j$ and AE-$i$ pair if the 2.5th posterior percentile for $\text{IC}_{ij}$ exceeds 0.
-While this fixed threshold of 0 is not necessarily guaranteed to control false discoveries, methods to generate data-driven thresholds ensuring a predefined FDR level (e.g., $0.05$) have been proposed.ahmed2010false
+While this fixed threshold of 0 is not necessarily guaranteed to control false discoveries, methods to generate data-driven thresholds ensuring a predefined FDR level (e.g., $0.05$) have been proposed.Ahmed et al.
+(2010)
 3.2 The Gamma Poisson Shrinker (GPS) model
@@
 The resulting posterior of $\lambda_{ij}$ conditional on these estimated hyperparameters is used for inference and signal detection.
-Specifically, dumouchel2001empirical define the empirical Bayesian score $E(\log_{2}(\lambda_{ij})\mid n_{ij},E_{ij},\hat{\kappa},\hat{\alpha}_{1},\hat{\beta}_{1},\hat{\alpha}_{2},\hat{\beta}_{2})$ as a point estimate of signal strength for the drug-$j$ and AE-$i$ pair, and also suggest determining the pair to be a signal if the 5th percentile (EB05) of the above conditional posterior for $\lambda_{ij}$ exceeds a specific threshold, commonly 2.
-Similar to the BCPNN method, the deterministic threshold of $2$ may not necessarily ensure strong FDR control; however, a data-driven threshold controlling FDR to a pre-specified level (e.g., 0.05) may be obtained.ahmed2010false
-3.3 DP Poisson method of Hu et al.hu2015signal
-Several extensions and generalizations of the two-component mixture prior used in the GPS method have been suggested in the literaturehu2015signal ; zhang2018three ; tan2025flexible .
-Of these, we focus on one of the most flexible generalizations as proposed in Hu et al.hu2015signal , which employs a Dirichlet process mixture as the prior for $\lambda_{ij}$.
-Hu et al.hu2015signal consider the same Poisson likelihood $n_{ij}\sim$ Poisson($\lambda_{ij}E_{ij}$) as in GPS, but replace the two-gamma mixture prior for $\lambda_{ij}$ by a DP (infinite) mixture prior:
+Specifically, DuMouchel and Pregibon (2001) define the empirical Bayesian score $E(\log_{2}(\lambda_{ij})\mid n_{ij},E_{ij},\hat{\kappa},\hat{\alpha}_{1},\hat{\beta}_{1},\hat{\alpha}_{2},\hat{\beta}_{2})$ as a point estimate of signal strength for the drug-$j$ and AE-$i$ pair, and also suggest determining the pair to be a signal if the 5th percentile (EB05) of the above conditional posterior for $\lambda_{ij}$ exceeds a specific threshold, commonly 2.
+Similar to the BCPNN method, the deterministic threshold of $2$ may not necessarily ensure strong FDR control; however, a data-driven threshold controlling FDR to a pre-specified level (e.g., 0.05) may be obtained.Ahmed et al.
+(2010)
+3.3 DP Poisson method of Hu et al.Hu et al. (2015)
+Several extensions and generalizations of the two-component mixture prior used in the GPS method have been suggested in the literatureHu et al.
+(2015); Zhang et al.
+(2018); Tan et al.
+(2025).
+Of these, we focus on one of the most flexible generalizations as proposed in Hu et al.Hu et al.
+(2015), which employs a Dirichlet process mixture as the prior for $\lambda_{ij}$.
+Hu et al.Hu et al.
+(2015) consider the same Poisson likelihood $n_{ij}\sim$ Poisson($\lambda_{ij}E_{ij}$) as in GPS, but replace the two-gamma mixture prior for $\lambda_{ij}$ by a DP (infinite) mixture prior:
 $\lambda_{ij}\sim D_{j}\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}};\ D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j})).$ (1)
@@
 Meanwhile, the atoms $\theta_{hj}$ are generated independently from the DP base gamma distribution: $\theta_{hj}\sim G(\beta_{j})$; $h\geq 1$.
-In practice, the infinite mixture (1) defining $D_{j}$ can be well-approximated ishwaran2002exact by a finite mixture $\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$ with a sufficiently large number of components $K$; this finite mixture approximation provides a computationally tractable approach to implementing this model.
-Hu et al.hu2015signal consider full Bayesian inference on $\lambda_{ij}$ by first eliciting independent vague prior distributions on the hyperparameters $\alpha_{j}$ (specifically a Uniform(0.2, 10) prior) and $\beta_{j}$ (a Uniform(0, 1) prior) and then utilizing MCMC sampling from the resulting joint posterior.
+In practice, the infinite mixture (1) defining $D_{j}$ can be well-approximated Ishwaran and Zarepour (2002) by a finite mixture $\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$ with a sufficiently large number of components $K$; this finite mixture approximation provides a computationally tractable approach to implementing this model.
+Hu et al.Hu et al.
+(2015) consider full Bayesian inference on $\lambda_{ij}$ by first eliciting independent vague prior distributions on the hyperparameters $\alpha_{j}$ (specifically a Uniform(0.2, 10) prior) and $\beta_{j}$ (a Uniform(0, 1) prior) and then utilizing MCMC sampling from the resulting joint posterior.
 Subsequently, the drug-$j$ and AE-$i$ pair is identified as a signal if the 5th posterior percentile of $\lambda_{ij}$ computed using its posterior MCMC draws exceeds a certain threshold.
-The authors suggest using the deterministic threshold of $2$ for this purpose; a more data-driven threshold controlling FDR at a prespecified level (e.g., $0.05$) may be obtained.ahmed2010false
+The authors suggest using the deterministic threshold of $2$ for this purpose; a more data-driven threshold controlling FDR at a prespecified level (e.g., $0.05$) may be obtained.Ahmed et al.
+(2010)
 Our proposed approach is motivated by this method, and we revisit this approach elucidating the key roles played by the model hyperparameters $\alpha_{j}$ and $\beta_{j}$, and guiding their prior specification for full Bayesian inference in Section 4.1.
 3.4 ZIP model based Pseudo LRT (frequentist) method
-In parallel to Bayesian developments, major advances on SRS data mining for AE signal detection have also been made using frequentist methods, notably using the machinery of the likelihood ratio test huang2011likelihood ; huang2013likelihood ; huang2014likelihood .
+In parallel to Bayesian developments, major advances on SRS data mining for AE signal detection have also been made using frequentist methods, notably using the machinery of the likelihood ratio test Huang et al.
+(2011, 2013, 2014).
 These approaches often have the advantage of built-in FDR control while ensuring high sensitivity.
 We consider the ZIP model for the reporting count $n_{ij}\sim\text{ZIP}(n=n_{ij}\mid\lambda_{ij}\times E_{ij},\omega_{j})$.
-The pseudo LRT, as proposed by Chakraborty et al.chakraborty2022use , is a recent method designed to accommodate the zero-inflated nature of SRS data while easing computational demands.
+The pseudo LRT, as proposed by Chakraborty et al.Chakraborty et al.
+(2022), is a recent method designed to accommodate the zero-inflated nature of SRS data while easing computational demands.
 Similar to the standard LRT method, this approach first tests a global alternative hypothesis $H_{aj}:\lambda_{ij}>1$ for at least one $i\in\{1,\dots,I\}$ against the null hypothesis $H_{0j}:\lambda_{ij}=1$ for all $i\in\{1,\dots,I\}$ for a given drug-$j$.
@@
 In the global test, a parametric bootstrap is used to approximate the p-value $P_{j}=\Pr(\text{MLR}_{j}\geq\text{MLR}_{j}^{\text{obs}})$ using (parametric bootstrap) draws for $\text{LR}_{ij}$ under $H_{0j}$ and hence $\text{MLR}_{j}$, where $\text{MLR}_{j}^{\text{obs}}$ is the observed test statistic.
-If the global test is rejected, step-down p-values chakraborty2022use ; huang2011likelihood are computed for the individual tests using the same parametric bootstrap samples for $\text{MLR}_{j}$ and using the individual observed likelihood ratio test statistic values, such that $P_{ij}=\Pr(\text{LR}_{ij}\geq\text{MLR}_{ij}^{\text{obs}})$.
+If the global test is rejected, step-down p-values Chakraborty et al.
+(2022); Huang et al.
+(2011) are computed for the individual tests using the same parametric bootstrap samples for $\text{MLR}_{j}$ and using the individual observed likelihood ratio test statistic values, such that $P_{ij}=\Pr(\text{LR}_{ij}\geq\text{MLR}_{ij}^{\text{obs}})$.
 For multiple drugs, the same approach can be applied, defining an overall maximum likelihood ratio as $\text{MLR}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\},j\in\{1,\dots,J\}\end{subarray}}\text{LR}_{ij}$.
-This step-down procedure employing the same global test statistic strictly controls the false discovery rate at the same level $\alpha$ of the global test performed using the MLR huang2011likelihood ; chakraborty2022use .
+This step-down procedure employing the same global test statistic strictly controls the false discovery rate at the same level $\alpha$ of the global test performed using the MLR Huang et al.
+(2011); Chakraborty et al.
+(2022).
 When multiple drugs $j$ are tested simultaneously, one considers the extended version with global maximum likelihood ratio $\max_{j\in\{1,\dots,J\}}\text{MLR}_{j}$.
@@
 We begin this section by revisiting the DP Poisson model of Hu et al.
-hu2015signal (Section 4.1), highlighting the role of its hyperparameters in SRS signal discovery with the aid of a schematic plot (Figure 2a) and providing guidance on prior specification for these hyperparameters.
+Hu et al.
+(2015) (Section 4.1), highlighting the role of its hyperparameters in SRS signal discovery with the aid of a schematic plot (Figure 2a) and providing guidance on prior specification for these hyperparameters.
 We then introduce our novel local-global DP framework (Section 4.2), which accommodates between-drug associations and enables effective information sharing across drugs, thereby enhancing SRS data mining.
@@
 When $K_{+}<I$, multiple $\lambda_{ij}$ share the same cluster $h$ admitting the common value $\theta_{hj}$, facilitating their pooling for information sharing.
-It follows that a smaller $\alpha$ encourages a smaller $K_{+}$, resulting in fewer clusters and thus greater information sharing among the $\lambda_{ij}$, while a larger $\alpha$ induces a larger $K_{+}$, resulting in more clusters that collectively capture more heterogeneity among the $\lambda_{i}j$ values.
+It follows that a smaller $\alpha$ encourages a smaller $K_{+}$, resulting in fewer clusters and thus greater information sharing among the $\lambda_{ij}$, while a larger $\alpha$ induces a larger $K_{+}$, resulting in more clusters that collectively capture more hetergeneity among the $\lambda_{i}j$ values.
 Importantly, when supported by the data and induced through a (moderately) large $\alpha$, values of $K_{+}>2$ are permitted in the posterior of this method.
-This contrasts with the GPS method of dumouchel1999bayesian (see Section 3 for a review) which strictly categorizes $\lambda_{ij}$ into two groups, intuitively, signal ($\lambda_{ij}\leq 1$) and non-signal ($\lambda_{ij}\gg 1$).
+This contrasts with the GPS method of DuMouchel (1999) (see Section 3 for a review) which strictly categorizes $\lambda_{ij}$ into two groups, intuitively, signal ($\lambda_{ij}\leq 1$) and non-signal ($\lambda_{ij}\gg 1$).
 The DP framework, by allowing $K_{+}>2$, enhances heterogeneity acknowledgment within both signal and non-signal groups.
@@
 Because a smaller $\alpha_{j}$ promotes greater information sharing, a prior assigning sufficient mass to small positive values is recommended for stable estimation of $\lambda_{ij}$.
-Hu et al.hu2015signal used a Uniform(0.2, 10) prior for $\alpha_{j}$.
-We suggest a more aggressive weakly informative default prior, viz., Exponential(rate = $\psi_{\alpha}$) with a moderately large $\psi_{\alpha}$, e.g., $\psi_{\alpha}=3$ (placing $\approx 95\%$ mass in $(0,1)$).
-This prior favors small $\alpha_{j}$ while allowing larger values if needed.
-This choice also ensures conjugacy for efficient Gibbs sampling updates for MCMC implementation.
+Hu et al.Hu et al.
+(2015) used a Uniform(0.2, 10) prior for $\alpha_{j}$.
+We suggest a more aggressive weakly informative default prior, viz., Exponential(rate = $\psi_{\alpha}$) with a moderately large $\psi_{\alpha}>1$.
+This prior favors small $\alpha_{j}$ while allowing larger values if needed and ensures conjugacy for efficient Gibbs sampling updates in the MCMC implementation.
+As a default choice, we suggest $\psi_{\alpha}=3$, which places $\approx 95\%$ prior mass in $(0,1)$.
+However, our simulation-based sensitivity analyses using typically sized SRS datasets indicate that model performance is largely robust to the choice of $\psi_{\alpha}$ when specified within a weakly informative range; see Remark Remark.
 Figure 2: Schematic diagrams visualizing the mechanisms of a “local only” DP prior with one drug (Panel A) and local-global mixture DP prior with $J=3$ drugs (Panel B).
@@
 Panel B: The noisy observed data $(n_{ij},E_{ij})$ leading to noisy observed relative reporting rates $n_{ij}/E_{ij}$ (Panel B.1; different shades of green) are used to produce “local” signal strengths $\lambda_{ij}^{l}$ for $I=6$ AEs and $J=3$ drugs using a separate, independent local DP prior for each drug (Panel B.2.1), and global AE signal strengths $\lambda_{i}^{g}$ for shared by all $J=3$ drugs based on a common global DP prior (Column B.2.2).
-Probabilistic two-component mixtures of $\lambda_{ij}^{l}$ and $\lambda_{i}$ (Panel B.3) based on the local distribution indicators $z_{ij}^{l}$ produce the final realized signal strengths $\lambda_{ij}$ (Panel B.4).
+Probabilistic two-component mixtures of $\lambda_{ij}^{l}$ and $\lambda_{i}$ (Panel B.3) based on the local distribution indicators $z_{ij}$ produce the final realized signal strengths $\lambda_{ij}$ (Panel B.4).
 Role and prior specification for the hyperparameter $\beta_{j}$
@@
 Under the same likelihood $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$, we consider the following local-global mixture of DP prior for $\lambda_{ij}$:
-$\lambda_{ij}=z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g};\ \lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D;\ z_{ij}^{l}\sim\text{Bernoulli}(\pi^{l})$
+$\lambda_{ij}=z_{ij}\lambda_{ij}^{l}+(1-z_{ij})\lambda_{i}^{g};\ \lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D;\ z_{ij}\sim\text{Bernoulli}(\pi)$
 $D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}));\ D\sim\text{DP}(\alpha,G(\beta)).$ (3)
 Here, the realized signal strengths $\{\lambda_{1j},\dots,\lambda_{Ij}\}$ for a drug $j$ is generated by combining local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$ drawn from a drug-$j$-specific local distribution $D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$ as in (1) and global strengths $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ generated from a global distribution $D\sim\text{DP}(\alpha,G(\beta))$ shared by all drugs.
-The binary local distribution indicators $z_{ij}^{l}$ governed by the common local proportion $\pi^{l}$, determine the probabilistic mixing of $\lambda_{ij}^{l}$ and $\lambda_{i}^{g}$.
+The binary local distribution indicators $z_{ij}$ governed by the common local proportion $\pi$, determine the probabilistic mixing of $\lambda_{ij}^{l}$ and $\lambda_{i}^{g}$.
 We note that the proposed local-global mixture DP prior (3) enables information sharing in signal strengths across both AEs and drugs.
@@
 Figure 2B illustrates this mechanism using $I=6$ AEs and $J=3$ drugs.
-This bi-level clustering contrasts with the single-level (AE-only) clustering induced by the DP prior of Hu et al.hu2015signal .
+This bi-level clustering contrasts with the single-level (AE-only) clustering induced by the DP prior of Hu et al.Hu et al.
+(2015).
 As demonstrated in extensive simulations (Section 5), the bi-level clustering and information sharing play a critical role in stabilizing the posterior and enhancing signal detection for $\lambda_{ij}$, particularly for drug-AE pairs where $n_{i\bullet}$ is small (rare AEs) and/or $n_{\bullet j}$ is small (newer drugs), leading to small $E_{ij}$.
@@
 Remark.
-The Hu et al.hu2015signal model can be viewed as a “local-only” special case of our local-global model, with only drug-specific local distributions $D_{j}$ for $\lambda_{ij}$ and no global distribution (i.e., $\pi^{l}=1$).
-At the other extreme, a “global-only” model ($\pi^{l}=0$) yields common $\lambda_{i}$ across drugs.
-This is a restrictive model that cannot differentiate between drugs and only permits “global” signal detection—assessing whether an AE-$i$ is a signal for any drug under consideration (see Supplement S.1.2 for further discussion on this model).
+As a special case of our local-global model, the “local-only” DP model of Hu et al.Hu et al.
+(2015) can be recovered by setting the common local proportion $\pi=1$, which leads to $z_{ij}=1$ and hence $\lambda_{ij}=\lambda_{ij}^{l}$ for all $i=1,\dots,I$ and $j=1,\dots,J-1$.
+Conversely, a “global-only” model can be obtained by setting $\pi=0$, leading to $\lambda_{ij}=\lambda_{i}^{g}$ for all $i$ and all $j<J$.
+We note, however, that this model is considerably more restrictive, as it does not distinguish among $\{\lambda_{i1},\dots,\lambda_{i,J-1}\}$ for any $i$.
+This yields the simplified likelihood $n_{ij}\sim\text{Poisson}(\lambda_{i}E_{ij})$, which produces the sufficient statistic $\sum_{j=1}^{J-1}n_{ij}\sim\text{Poisson}(\lambda_{i}\sum_{j=1}^{J-1}E_{ij})$ for $\lambda_{i}$.
+The model reduction also impacts signal detection.
+Under both the general local-global and local-only frameworks, hypothesis testing is conducted for each individual $\lambda_{ij}$.
+In contrast, under the global-only model, inference reduces to testing whether $\lambda_{i}^{g}>1$.
+Testing $\lambda_{i}^{g}>1$ amounts to assessing whether the aggregated count $\sum_{j=1}^{J-1}n_{ij}$ is large relative to the expected total count $\sum_{j=1}^{J-1}E_{ij}$ under the reduced model.
+This represents a type of “global test” that collapses all drugs into one homogeneous group, whereas the local-global (and local-only) models allow hypothesis testing for each individual drug-AE pair.
+Consequently, direct comparison between the global-only and local-global models is usually not practically meaningful, as they target different inferential objectives.
 Remark.
 Extensive literature exists on local-global DP models for sharing information across groups while preserving DP-based clustering within each group.
-A common approach employs hierarchical DPs teh2004sharing , where local DPs share a global DP as a base distribution.
-In contrast, our approach leverages a two-component DP mixture framework antoniak1974mixtures , separately modeling local and global AE signal strengths and thus offering a more transparent bi-level clustering over AEs and drugs, enabling more flexible estimation of the local and global parameters for greater flexibility in SRS applications.
+A common approach employs hierarchical DPs Teh et al.
+(2004), where local DPs share a global DP as a base distribution.
+In contrast, our approach leverages a two-component DP mixture framework Antoniak (1974), separately modeling local and global AE signal strengths and thus offering a more transparent bi-level clustering over AEs and drugs, enabling more flexible estimation of the local and global parameters for greater flexibility in SRS applications.
 Remark.
 In this paper, we focus on drug-based pharmacovigilance, identifying AE (row) signals given one or more drugs (columns in SRS contingency tables).
-An alternative AE-based approach can be performed to discover drugs strongly associated with a given set of AEs ding2020evaluation ; liu2024pattern .
+An alternative AE-based approach can be performed to discover drugs strongly associated with a given set of AEs Ding et al.
+(2020); Liu et al.
+(2024).
 Our proposed approach can be readily adapted for such analyses: a row-based version of our column-based prior (3), obtained by switching the roles of $i$ and $j$, can be used for this purpose.
@@
 For full Bayesian inference on our models, we place independent weakly informative proper priors on all model hyperparameters.
-Along with the exponential prior for $\alpha_{j}$ and $\alpha$, and half-Cauchy prior for $\beta_{j}$ and $\beta$ as suggested in Section 4.1, we consider a similar Cauchy+(0, $\psi_{\tau}$) prior for $\tau$ with $\psi_{\tau}=1/2$, and independent Uniform(0,1)(0,1) priors for the probability weight parameters, viz., $\pi_{l}$ and $\omega_{j}$ (if included in the model).
-We develop efficient MCMC algorithms for rigorous posterior sampling, approximating the DPs using finite mixtures with $K=O(\log(IJ))$ components ishwaran2001gibbs .
+Along with the exponential prior for $\alpha_{j}$ and $\alpha$, and half-Cauchy prior for $\beta_{j}$ and $\beta$ as suggested in Section 4.1, we consider a similar Cauchy+(0, $\psi_{\tau}$) prior for $\tau$ with $\psi_{\tau}=1/2$, and independent Uniform(0,1)(0,1) priors for the probability weight parameters, viz., $\pi$ and $\omega_{j}$ (if included in the model).
+We develop efficient MCMC algorithms for rigorous posterior sampling, approximating the DPs using finite mixtures with $K=O(\log(IJ))$ components Ishwaran and James (2001).
 Each MCMC iteration involves data augmentation to generate latent DP mixture allocations, followed by Gibbs sampling updates for the stick-breaking weights and atoms and the concentration hyperparameters for each local and global DP.
-The local distribution indicators $z_{ij}^{l}$, local distribution proportion $\pi^{l}$, AE signal strengths $\lambda_{iJ}$ for the reference drug $J$, and zero-inflation indicators $y_{ij}$ and probabilities $\omega_{j}$ (if included in the model) also admit standard Gibbs updates.
-However, the full conditionals for the DP base Gamma hyperparameters ($\beta_{j}$ and $\beta$) and the reference category shrinkage hyperparameter $\tau$ lack standard forms; for these we employ univariate stepping-out slice sampling neal2003slice steps.
+The local distribution indicators $z_{ij}$, local distribution proportion $\pi$, AE signal strengths $\lambda_{iJ}$ for the reference drug $J$, and zero-inflation indicators $y_{ij}$ and probabilities $\omega_{j}$ (if included in the model) also admit standard Gibbs updates.
+However, the full conditionals for the DP base Gamma hyperparameters ($\beta_{j}$ and $\beta$) and the reference category shrinkage hyperparameter $\tau$ lack standard forms; for these we employ univariate stepping-out slice sampling Neal (2003) steps.
 Detailed MCMC steps are provided in Supplementary Algorithms S.1, S.2 and S.3.
@@
 In this section, we conduct extensive simulations to assess the signal detection performance of our local-global DP Poisson and ZIP models under a wide array of data-generating settings.
-We follow the SRS contingency table generation approach previously used in the literature chakraborty2022use ; liu2024pattern .
+We follow the SRS contingency table generation approach previously used in the literature Chakraborty et al.
+(2022); Liu et al.
+(2024).
 Specifically, to generate random SRS contingency tables of dimension $I\times J$, we first define a prespecified true signal strength matrix (described below) $\Lambda^{0}=((\lambda_{ij}^{0}))$ encoding signals, structural zeros, non-signals, and between-drug dependency patterns.
@@
 Reference SRS dataset to compute row and column totals
-We use the FDA FAERS database (Q3 2014-Q4 2020), which catalogs a total of 63,976,61063{,}976{,}610 reported cases chakraborty2023likelihood , to construct our reference SRS contingency table.
+We use the FDA FAERS database (Q3 2014-Q4 2020), which catalogs a total of 63,976,61063{,}976{,}610 reported cases Chakraborty et al.
+(2023), to construct our reference SRS contingency table.
 We focus on six statin medications: Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, collapsing all other drugs into a comparator “Other drugs” category.
-For expository purposes, and to create a sufficiently large SRS contingency table, we focused on 1491 preferred terms (PTs) as AEs that are common (appear at least once) for these statin drugs as well for a separate group of drugs unrelated to statin, namely, Gadolinium-based contrast agents (GBCA) drugs chakraborty2022use .
+For expository purposes, and to create a sufficiently large SRS contingency table, we focused on $I=1491$ preferred terms (PTs) as AEs that are common (appear at least once) for these statin drugs as well for a separate group of drugs unrelated to statin, namely, Gadolinium-based contrast agents (GBCA) drugs Chakraborty et al.
+(2022).
 All other PTs recorded in the database within this period were collapsed into a heterogeneous category labeled “Other AE”, yielding a $1492\times 7$ contingency table summarizing the report counts for each drug-AE pair.
+Alongside this larger setting, we also consider a smaller setting involving a subset of this dataset with $I=46$ relevant AEs as considered in prior workChakraborty et al.
+(2022); Huang et al.
+(2011) and a collapsed Other AE cateogory; see Supplement S.2.5.
 Constructing the true signal strength matrix $\Lambda^{0}$
@@
 This approach yields a controlled $\Lambda^{0}$ to generate random SRS contingency tables.
-Within this setup, we designed three simulation scenarios, each with and without zero-inflation (see Table 1).
+Within this setup, we consider four simulation scenarios with varied range of between-column correlations in $\Lambda^{0}$, each with and without zero-inflation (see Table 1).
 Table 1: Simulation scenarios describing the construction of the true signal strength matrix $\Lambda^{0}$ encoding signals, non-signals, zero-inflation, and between-drug associations.
@@
 Finally, the fourth row in the table shows the average Kendall’s $\tau$ between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ across all statin drug pairs $j$ and $j^{\prime}$ (excluding the “Other drugs” category) as a measure of pairwise association of their AE signal strengths.
-Simulation Scenario 1a 1b 2a 2b 3a 3b
-No. of fixed row signals 5 5 10 10 20 20
-No. of random signals per column 17 17 10 10 3 3
-Zero-inflation rate per column 0 0.3 0 0.3 0 0.3
-Average Kendall’s $\tau$ between columns 0.12 0.12 0.49 0.49 0.77 0.77
-  • A common signal strength value $\lambda^{0,\text{signal}}$ is assigned to all signal cells in a given simulation setting, and we vary $\lambda^{0,\text{signal}}\in\{1.1,1.2,1.3,1.4,1.5,1.7,1.9,2,2.5,3\}$.
+Simulation Scenario 0a 0b 1a 1b 2a 2b 3a 3b
+No. of fixed row signals 1 1 3 3 10 10 20 20
+No. of random signals per column 30 30 20 20 10 10 3 3
+Zero-inflation rate per column 0 0.3 0 0.3 0 0.3 0 0.3
+Average Kendall’s $\tau$ between column pairs $\leq$0.01 $\leq$0.01 0.12 0.12 0.49 0.49 0.77 0.77
+  • A common signal strength value $\lambda^{0,\text{signal}}$ is assigned to all signal cells in a given simulation setting, and we vary $\lambda^{0,\text{signal}}\in\{1.01,1.1,1.2,1.3,1.4,1.5,1.7,1.9,2,2.5,3\}$.
 Replicated Data Generation, Model Fitting, and signal detection
@@
 From the computed posteriors, we conduct signal detection using hypothesis testing (Section 4.4).
-We compare our signal detection performance against existing Bayesian SRS data mining methods, viz., BCPNN, GPS, and the local-only DP model of Hu et al.hu2015signal (henceforth DP Hu et al).
+We compare our signal detection performance against existing Bayesian SRS data mining methods, viz., BCPNN, GPS, and the local-only DP model of Hu et al.Hu et al.
+(2015) (henceforth DP Hu et al).
 For our models and DP Hu et al., signal detection uses an optimized posterior quantile level $\hat{p}$ (fixed at 0.05 for DP Hu et al.) and threshold $\hat{k}$ ensuring FDR $\leq 0.05$ while minimizing FNR (Section 4.4), with the additional constraint $n_{ij}>1$ for our methods.
-To ensure fair FDR control ($\leq 0.05$) in BCPNN and GPS, we leverage the independence of signal strengths in the (empirical) Bayes posteriors for these two models and compute Benjamini-Hochberg-adjusted muller2006fdr posterior probabilities of being a non-signal; a drug-$j$ and AE-$i$ pair is then classified as a signal if the adjusted probability is $\leq 0.05$.
-Additionally, we evaluate the state-of-the-art frequentist ZIP-based extended (pseudo) LRT chakraborty2022use ; huang2011likelihood ; ding2020evaluation , which ensures FDR control ($\leq 0.05$) with high sensitivity while addressing zero inflation in the counts when present.
+To ensure fair FDR control ($\leq 0.05$) in BCPNN and GPS, we leverage the independence of signal strengths in the (empirical) Bayes posteriors for these two models and compute Benjamini-Hochberg-adjusted Müller et al.
+(2006) posterior probabilities of being a non-signal; a drug-$j$ and AE-$i$ pair is then classified as a signal if the adjusted probability is $\leq 0.05$.
+Additionally, we evaluate the state-of-the-art frequentist ZIP-based extended (pseudo) LRT Chakraborty et al.
+(2022); Huang et al.
+(2011); Ding et al.
+(2020), which ensures FDR control ($\leq 0.05$) with high sensitivity while addressing zero inflation in the counts when present.
 We assess signal detection for all methods using replication-based FDR, sensitivity, average Type I error, and F-score (Figure 3).
@@
 Each column presents one metric and each row presents one simulation scenario in Table 1.
-The first two rows of Figure 3 show comparable (when $\lambda^{0,\text{signal}}<1.5$) or slightly better (when $\lambda^{0,\text{signal}}\geq 1.5$) sensitivity for our local-global methods compared to all other methods when the between-drug association is very small ($\tau=0.12$; Cases 1a,b).
+Our results show that across all settings, the FDR and average type I error are strictly controlled by all methods except BCPNN, which appears to entirely lose control of FDR under the near-null setting of $\lambda^{0,\text{signal}}=1.01$, despite the use of a Benjamini-Hochberg-based FDR adjustment during signal detection.
+Focusing on sensitivity, the first four rows of Figure 3 show comparable (when $\lambda^{0,\text{signal}}<1.5$) or slightly better (when $\lambda^{0,\text{signal}}\geq 1.5$) sensitivity for our local-global methods compared to all other methods when the there is very weak ($\tau\leq 0.01$; Cases 0a, b) or weak ($\tau=0.12$; Cases 1a,b) between-drug associations.
 Under moderate between-drug association ($\tau=0.49$; Cases 2a,b), our methods perform noticeably better in detecting more true signals, achieving up to twice the sensitivity, particularly for weak or moderate true signal strengths ($1.1\leq\lambda^{0,\text{signal}}\leq 1.7$).
@@
 These results highlight the substantial boost in signal detection from effectively leveraging between-drug similarities as facilitated by our models.
-In all settings, the FDR and average type I error is strictly controlled for all methods.
 The Poisson and ZIP-based DP local-global models perform similarly across all scenarios regardless of the presence of zero-inflated observed counts (sub-cases a and b), due to the flexibility of the DP-based nonparametric framework (see Section 4.1).
+The results from the smaller setting involving $I=46$ AEs are essentially similar to those obtained in the larger setting; see Supplementary Figure S.2 in Supplement S.3.4 for further details.
 All computations were performed on the cluster computing facility at the Center for Computational Research at the University at Buffalo, with each process running on 16 GB RAM and a clock speed of 1.6-3 GHz.
-The median computation time per replication for our DP local-global methods (both Poisson and ZIP) is 48 minutes.
+The median computation time per replication for our DP local-global methods (both Poisson and ZIP) is approximately 48 minutes for the larger setting with 1491 AEs, compared with 0.66, 0.02, 0.42, and 26 minutes for BCPNN, GPS, LRT, and DP Hu et al., respectively.
+For the smaller 46-AE dataset, our methods require about 2 minutes per replication, versus 0.02, 0.002, 0.07, and 0.7 minutes for the same competing approaches.
+The higher computational cost reflects our substantially richer modeling assumptions: except for DP Hu et al., the competing methods achieve highly efficient computation through partial analytical tractability but do not provide flexible, regularized estimation of signal strength parameters with full uncertainty quantification, nor do they capture complex between-drug associations in AE signal patterns.
+Relative to DP Hu et al., our methods introduce only one additional global DP component and the local-global mixture parameters, incurring modest extra cost while delivering substantial gains in sensitivity and signal detection, as demonstrated in our simulations.
+Remark.
+Leveraging the simulation setup described as above, we performed additional sensitivity analyses to examine the robustness of the our methods with respect to the choice of hyperparameters, including the prior hyperparameters $\psi_{\alpha}$ and $\psi_{\beta}$, as well as the grid specification for the hypothesis testing threshold $k$.
+For each of $\psi_{\alpha}$ and $\psi_{\beta}$, we considered two alternative choices representing values toward the lower and upper extremes of the weakly informative range along with the default $\psi_{\alpha}=3$ and $\psi_{\beta}=1/2$.
+Specifically, for $\psi_{\alpha}$ we considered $\psi_{\alpha}=1$ (placing $\approx 64\%$ mass in $(0,1)$) and $\psi_{\alpha}=5$ ($\approx 99\%$ mass in $(0,1)$), and for $\psi_{\beta}$ we considered $\psi_{\beta}=1/7$ ($\approx 90\%$ mass in $(0,1)$) and $\psi_{\beta}=6$ (only $10\%$ mass in $(0,1)$).
+For the $k$-grid construction, we considered a more adaptive approach monitoring the computed maximum of all $\lambda_{ij}$ posterior draws instead of the default grid on $[1.1,3]$.
+Across all settings, the resulting performance curves were virtually identical to those obtained under the default specifications, indicating that the proposed method is largely stable across a range of reasonable prior choices and grid settings.
+A detailed description of the sensitivity analysis results is provided in the Supplement S.3.2.
 6 REAL DATA ANALYSIS
+Figure 4: Heatmaps of 15 selected PTs.
+Cells not detected as signals by our method are colored gray.
+White-to-blue shades represent the median of the posterior of signal strength parameter $\lambda_{ij}$, with darker (lighter) blue indicating a stronger (weaker) signal strength.
+Panel A contains 5 common PTs among all statin drugs.
+Panel B contains 5 semi-rare PTs with new discovered signals by our proposed method.
+Panel C contains 5 rare PTs with new discovered signals by our proposed method.
+Count $n$, expected count $E$, and the 90% credible interval of the posterior distribution of $\lambda_{ij}$ are shown in each cell.
+The letters “H”, “B”, “G”, and “L” indicate signal detection by the DP Hu et al., BCPNN, GPS, and likelihood ratio test, respectively.
 We use the statin data from Section 5 to demonstrate our method’s utility in real-world SRS data mining.
@@
 Panel A displays the five most common AEs with the largest frequencies (total counts for the AEs across all six statin drugs being greater than $2000$) and the strongest estimated signal strengths (having the largest posterior medians of $\lambda_{ij}$ across all statin drugs $j$) that are detected as a signal for at least one statin drug by our method.
-The AEs displayed in this panel are all well-documented muscle-related conditions, such as rhabdomyolysis antons2006clinical ; tournadre2020statins and myopathy ballantyne2003risk ; abd2011statin , and their detection by our method is consistent with the performance of all competing methods.
+The AEs displayed in this panel are all well-documented muscle-related conditions, such as rhabdomyolysis Antons et al.
+(2006); Tournadre (2020) and myopathy Ballantyne et al.
+(2003); Abd and Jacobson (2011), and their detection by our method is consistent with the performance of all competing methods.
 Panel B contains five semi-rare AEs (with observed counts across statin drugs ranging between 500 and 2000) that exhibit the strongest estimated signal strengths.
 This category primarily includes some specific types of disorders, such as nervous system disorders, and interestingly, several of these detected signals are not identified by any competing method for the newer or less frequently used statin drugs.
-For example, unlike other methods, our model detects lovastatin-induced drug intolerance as a signal, which has been flagged as a potential side effect of lovastatinahmad2014statin .
-Additionally, our method detects signals for paraesthesia and hypoaesthesia, symptoms of peripheral neuropathy, associated with lovastatin use; this finding is also consistent with previous research chong2004statin ; wang2021association .
+For example, unlike other methods, our model detects lovastatin-induced drug intolerance as a signal, which has been flagged as a potential side effect of lovastatinAhmad (2014).
+Additionally, our method detects signals for paraesthesia and hypoaesthesia, symptoms of peripheral neuropathy, associated with lovastatin use; this finding is also consistent with previous research Chong et al.
+(2004); Wang et al.
+(2021).
 By appropriately borrowing strengths from the more commonly used statins, our method is successfully able to identify potential lovastatin-induced AE signals in these specific examples.
@@
 Many of these AEs were not originally discovered in clinical trials, but later were flagged as potential signals in specific, targeted pharmacovigilance investigations.
-For example, fluvastatin-induced swollen tongue, also known as angioedema, has been recently recorded kumar2024fluvastatin .
+For example, fluvastatin-induced swollen tongue, also known as angioedema, has been recently recorded Kumar et al.
+(2024).
 Similar to the semi-rare AEs displayed in Panel B, our method is able to identify these rare AE signals for the less frequently used drug fluvastatin by appropriately sharing AE signal strength information across all statin drugs.
-Figure 4: Heatmaps of 15 selected PTs.
-Cells not detected as signals by our method are colored gray.
-White-to-blue shades represent the median of the posterior of signal strength parameter $\lambda_{ij}$, with darker (lighter) blue indicating a stronger (weaker) signal strength.
-Panel A contains 5 common PTs among all statin drugs.
-Panel B contains 5 semi-rare PTs with new discovered signals by our proposed method.
-Panel C contains 5 rare PTs with new discovered signals by our proposed method.
-Count $n$, expected count $E$, and the 90% credible interval of the posterior distribution of $\lambda_{ij}$ are shown in each cell.
-The letters “H”, “B”, “G”, and “L” indicate signal detection by the DP Hu et al., BCPNN, GPS, and likelihood ratio test, respectively.
 7 DISCUSSION
@@
 A reformulation via mixed membership models could further enhance interpretability by allowing the same AE to appear in multiple clusters/groups.
-Second, while we consider data-driven characterization of similarities in AE signal strengths between drugs, a more flexible approach acknowledging known Drug-Drug Interactions (DDI poleksic2019database ), possibly via appropriate prior distributions, may further enhance signal detection.
+Second, while we consider data-driven characterization of similarities in AE signal strengths between drugs, a more flexible approach acknowledging known Drug-Drug Interactions (DDI Poleksic and Xie (2019)), possibly via appropriate prior distributions, may further enhance signal detection.
 Finally, our current framework aggregates drug-AE data across time points for a cross-sectional analysis.
@@
 $\text{DP local-global prior for a non-baseline drug }j\in\{1,\dots,J-1\}\!:$
-$\lambda_{ij}\sim z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$
+$\lambda_{ij}\sim z_{ij}\lambda_{ij}^{l}+(1-z_{ij})\lambda_{i}^{g}$
 $\lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D$
-$z_{ij}^{l}\sim\text{Bernoulli}(\pi^{l})$
+$z_{ij}\sim\text{Bernoulli}(\pi)$
 DP local prior for drug-$j$:
@@
 $\text{DP local-global prior for a non-baseline drug }j\in\{1,\dots,J-1\}\!:$
-$\lambda_{ij}\sim z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$
+$\lambda_{ij}\sim z_{ij}\lambda_{ij}^{l}+(1-z_{ij})\lambda_{i}^{g}$
 $\lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D$
-$z_{ij}^{l}\sim\text{Bernoulli}(\pi^{l})$
+$z_{ij}\sim\text{Bernoulli}(\pi)$
 DP local prior for drug-$j$:
@@
 vague proper priors for hyperparameters:
-$\pi^{l}\sim\text{Uniform}(0,1)$
+$\pi\sim\text{Uniform}(0,1)$
 $\alpha_{j},\alpha\sim\text{Gamma}(\text{shape}=1,\text{rate}=3)$
@@
 $\tau\sim\text{Cauchy}^{+}(0,0.5)$
-S.1.2 Some notes on the “global-only” model
-A special case of our local-global models leads to a reduced “global-only” model when $\pi^{g}=0$, i.e., $\lambda_{ij}=\lambda_{i}^{g}$ for $j=1,\dots,J-1$.
-Under this model there is no provision for different $\lambda_{ij}$ for any $j\neq J$.
-The global model has a simplified likelihood $n_{ij}\sim\text{Poisson}(\lambda_{i}E_{ij})$ here $\lambda_{i}^{g}\sim D$ if $j\neq J$, which produces $\sum_{j=1}^{J-1}n_{ij}\sim\text{Poisson}(\lambda_{i}\sum_{j=1}^{J-1}E_{ij})$.
-The model reduction also impacts the signal detection.
-For the local-global framework, hypothesis testing is conducted for each individual $\lambda_{ij}$.
-On the other hand, under the global model, the test is simplified to evaluating whether $\lambda_{i}^{g}>1$.
-Testing $\lambda_{i}^{g}>1$ entails assessing if the summarized Poisson count $\sum_{i}n_{ij}$ is large compared to the expected count $\sum_{i}E_{ij}$, owing to the reduced model.
-This approach represents a “global test”, whereas our proposed local-global model allows conducting hypothesis testing for each individual drug-AE pair.
-As a result, directly comparing the global model with the local-global model is not feasible, as they target different types of assessment.
 Appendix S.2 COMPUTATIONAL DETAILS
@@
 Next we provide detailed notes on the grid-search method identifying optimal quantile-level $p$ and threshold $k$ used for our hypothesis tests for signal detection (Section S.2.3).
-The JAGS code [32] for Hu et al.’s model [18] is shown as a reference in Section S.2.4.
+The JAGS code [33] for Hu et al.
+[18]’s models is shown as a reference in Section S.2.4.
 The example statin dataset is included in Section S.2.5.
@@
     $\pi(\beta\mid\bm{\theta})\propto\prod_{h=1}^{K}f_{\text{Gamma}}(\theta_{h}\mid\text{shape}=\beta,\text{rate}=\beta)\pi_{\beta}(\beta)$
-    Here $f_{\text{Gamma}}(\theta\mid\text{shape}=\beta,\text{rate}=\beta)$ denotes the probability density function of the Gamma($\beta,\beta$) distribution at $\theta$.
+    Here $f_{\text{Gamma}}(\theta\mid\text{shape}=\beta,\text{rate}=\beta)$ denotes the probability density function of the Gamma($\beta,\beta$) distribution evaluated at $\theta$.
 $\pi(\beta)$ is the prior distribution of $\beta$.
@@
     $\alpha\mid v_{1},\dots,v_{K}\sim\text{Gamma}\left(\text{shape}=K,\text{rate}=-\sum_{h=1}^{K-1}\log(1-v_{h})+\psi_{\alpha}\right).$
-Algorithm S.2 One Step Iteration of DP Poisson local-global column model (Given data $\{(n_{ij},E_{ij},z_{ij}^{l}):i=1,\dots,I;j=1,\dots,J\}$ )
+Algorithm S.2 One Step Iteration of DP Poisson local-global column model (Given data $\{(n_{ij},E_{ij},z_{ij}):i=1,\dots,I;j=1,\dots,J\}$ )
   1. 1.
-    Generate global $\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$ parameters using Algorithm S.1, with $\{(\tilde{n}_{i},\tilde{E}_{i},1(\sum_{j=1}^{J-1}z_{ij}^{l}<J-1)):i=1,\dots,I\}$ as data where
-    $\tilde{n}_{j}=\sum_{j=1}^{J-1}n_{ij}(1-z_{ij}^{l}),\quad\tilde{E}_{j}=\sum_{j=1}^{J-1}E_{ij}(1-z_{ij}^{l}).$
+    Generate global $\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$ parameters using Algorithm S.1, with $\{(\tilde{n}_{i},\tilde{E}_{i},1(\sum_{j=1}^{J-1}z_{ij}<J-1)):i=1,\dots,I\}$ as data where
+    $\tilde{n}_{i}=\sum_{j=1}^{J-1}n_{ij}(1-z_{ij}),\quad\tilde{E}_{i}=\sum_{j=1}^{J-1}E_{ij}(1-z_{ij}).$
     1. $1.$
@@
     Generate “local” allocations:
-    $z_{ij}^{l}\mid n_{ij},E_{ij},z_{ij}^{l},\lambda_{ij}^{l},\lambda_{i}^{g}\sim\text{Bernoulli}(p_{ij}^{l}),$
+    $z_{ij}\mid n_{ij},E_{ij},z_{ij},\lambda_{ij}^{l},\lambda_{i}^{g}\sim\text{Bernoulli}(p_{ij}),$
     where
-    $p_{ij}^{l}=\frac{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})}{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})+(1-\pi^{l})f_{\text{Poisson}}(n_{ij}\mid\lambda_{i}^{g}E_{ij})}.$
+    $p_{ij}=\frac{\pi f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})}{\pi f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})+(1-\pi)f_{\text{Poisson}}(n_{ij}\mid\lambda_{i}^{g}E_{ij})}.$
   4. 4.
-    Define $\lambda_{ij}=z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$ for all $i=1,\dots,I$ and $j=1,\dots,(J-1)$.
+    Define $\lambda_{ij}=z_{ij}\lambda_{ij}^{l}+(1-z_{ij})\lambda_{i}^{g}$ for all $i=1,\dots,I$ and $j=1,\dots,(J-1)$.
   5. 5.
-    Generate $\pi^{l}$ using:
-    $\pi^{l}\mid z_{ij}^{l}\sim\text{Beta}\left(a^{l}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}z_{ij},b^{l}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}(1-z_{ij})\right),$
-    Notes: $a^{l}$ = 1, $b^{l}$ = 1
+    Generate $\pi$ using:
+    $\pi\mid z_{ij}\sim\text{Beta}\left(a_{\pi}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}z_{ij},b_{\pi}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}(1-z_{ij})\right),$
+    Notes: $a_{\pi}$ = 1, $b_{\pi}$ = 1.
   6. 6.
@@
     $\pi(\beta_{J}\mid\lambda_{1J},\dots,\lambda_{IJ})\propto\prod_{i=1}^{I}f_{\text{Gamma}}(\lambda_{ij}\mid\text{shape}=\tau,\text{rate}=\tau)\pi_{\tau}(\tau).$
-Algorithm S.3 One Iteration of DP ZIP local-global Model Given data $(n_{ij},E_{ij},z_{ij}^{l})$
+Algorithm S.3 One Iteration of DP ZIP local-global Model Given data $(n_{ij},E_{ij},z_{ij})$
   1. 1.
@@
 We initialize the number of nonempty components $K_{+}$ to the smallest integer greater than $\min\{K/2,I-1\}$.
-The local-global indicators $z_{ij}^{l}$ are randomly initialized by independently sampling each cell $(i,j)$ from a $\text{Bernoulli}(0.5)$ distribution.
-The DP parameters for each local and global DP process are then initialized using the subset of data $(n_{ij},E_{ij})$ assigned to that process, as determined by the indicators $z_{ij}^{l}$, for $i=1,\dots,J-1$.
+The local-global indicators $z_{ij}$ are randomly initialized by independently sampling each cell $(i,j)$ from a $\text{Bernoulli}(0.5)$ distribution.
+The DP parameters for each local and global DP process are then initialized using the subset of data $(n_{ij},E_{ij})$ assigned to that process, as determined by the indicators $z_{ij}$, for $i=1,\dots,J-1$.
 For the initialization of parameters in the local DP prior for the $j$-th drug ($j=1,\dots,J-1$), we perform $k$-means clustering with the initialized $K_{+}$ number of clusters on the set $\{(n_{ij}+0.5)/(E_{ij}+0.5):z_{ij}=1,\ i=1,\dots,I\}$, and use the resulting cluster centers as initial values for the atoms $\theta_{hj}$, $h=1,\dots,K_{+}$, of the non-empty clusters.
@@
 Table S.1: The first 30 rows of the SRS contingency table summarized from the FDA FEARS database (Q3 2014-Q4 2020) which catalogs a total of 63,976,610 reported cases.
+There are 1490 preferred terms in the full dataset.
 Adverse Event Atorvastatin Fluvastatin Lovastatin Pravastatin Rosuvastatin Simvastatin Other
@@
 Acute Hepatic Failure 107 0 10 1 18 66 14260
+We also present the full dataset with a smaller setting involving a subset of Table S.1 with only $I=46$ relevant AEs as considered in prior work[9, 19].
+It is used for data generation in the simulation of Section S.3.4.
+Table S.2: The SRS contingency table summarized from the FDA FEARS database (Q3 2014-Q4 2020) which catalogs a total of 63,976,610 reported cases. There are 46 preferred terms in the full dataset.
+Adverse Event Atorvastatin Fluvastatin Lovastatin Pravastatin Rosuvastatin Simvastatin Other
+Acute Kidney Injury 1353 42 7 154 689 823 355651
+Anuria 71 0 0 2 43 62 10403
+Blood Calcium Decreased 14 2 0 0 110 17 15918
+Blood Creatine Phosphokinase Abnormal 34 0 0 0 8 11 261
+Blood Creatine Phosphokinase Increased 1175 125 32 200 562 768 23805
+Blood Creatine Phosphokinase Mm Increased 2 0 0 0 9 0 14
+Blood Creatinine Abnormal 27 0 0 0 5 3 3385
+Blood Creatinine Increased 227 10 0 17 210 97 74742
+Chromaturia 340 10 6 33 174 114 19294
+Chronic Kidney Disease 152 16 2 19 177 37 339179
+Compartment Syndrome 53 0 0 1 12 12 2644
+Creatinine Renal Clearance Decreased 6 0 0 2 124 6 7768
+Diaphragm Muscle Weakness 14 0 0 0 8 1 94
+Electromyogram Abnormal 2 0 0 2 0 0 132
+End Stage Renal Disease 30 0 0 0 19 6 97553
+Glomerular Filtration Rate Abnormal 8 0 0 0 1 0 1069
+Glomerular Filtration Rate Decreased 59 1 0 8 39 29 13190
+Hypercreatininaemia 0 0 0 0 8 0 648
+Hypocalcaemia 36 0 0 16 8 18 23102
+Muscle Disorder 291 2 7 21 191 87 7329
+Muscle Enzyme Increased 48 1 0 0 13 9 410
+Muscle Fatigue 85 0 2 16 30 39 4257
+Muscle Haemorrhage 24 0 0 5 13 4 3806
+Muscle Necrosis 68 2 0 1 10 20 662
+Muscle Rupture 181 25 0 61 36 120 3219
+Muscular Weakness 1857 45 31 152 808 859 111003
+Musculoskeletal Discomfort 137 18 15 25 187 93 19931
+Musculoskeletal Disorder 56 3 0 9 65 73 25881
+Musculoskeletal Pain 420 3 2 38 324 228 82576
+Myalgia 5362 341 151 939 2757 3216 143819
+Myasthenic Syndrome 1 0 9 0 3 6 643
+Myoglobin Blood Increased 71 4 0 4 16 39 1003
+Myoglobin Blood Present 2 0 0 0 0 0 8
+Myoglobin Urine Present 0 0 0 0 1 2 70
+Myoglobinaemia 15 0 0 0 0 0 62
+Myoglobinuria 26 4 0 1 1 9 296
+Myopathy 849 64 45 145 377 544 6695
+Myopathy Toxic 31 0 0 1 2 21 457
+Myositis 219 8 10 28 62 141 7482
+Necrotising Myositis 279 0 0 2 10 52 278
+Oliguria 52 0 0 4 24 37 7590
+Renal Failure 534 26 11 69 225 195 250710
+Renal Impairment 390 52 11 37 161 169 103343
+Renal Tubular Necrosis 40 0 0 12 10 24 12762
+Rhabdomyolysis 2041 52 44 163 936 1376 31707
+Tendon Discomfort 9 0 0 3 22 10 794
+Other Pt 180699 4886 2845 20296 113960 76068 61724222
 Appendix S.3 ADDITIONAL SIMULATION RESULTS
-We present detailed values of the simulation to compare among our drug (column)-based models (Poisson/ZIP), DP Hu et al., and posterior quantile optimized DP Hu et al.
-method.
-Four evaluation metrics (FDR, sensitivity, average type I error, and F-score) are evaluated as well as the main text with same simulation settings in Table 1.
+We present detailed results from our simulation experiments to compare among our local-global models (Poisson/ZIP), DP Hu et al.
+model, and a posterior quantile-based, FNR-optimized DP Hu et al.
+test method.
+Four evaluation metrics (FDR, sensitivity, average type I error, and F-score) are reported under the same simulation settings in Table 1 of the main test.
 Frequentist and Bayesian MSE are compared to further evaluate the loss of the estimation.
 S.3.1 Simulation results for DP-based models
-We present the four evaluation metrics in Tables S.2, S.3, S.4, and S.5 to provide a clearer numerical comparison beyond Figure 3.
+We present the four evaluation metrics in Tables S.3, S.4, S.5, and S.6 to provide a clearer numerical comparison beyond Figure 3.
 Additionally, we further explored the effect of posterior quantile optimization on the DP Hu et al.
@@
 posterior quantile optim
-Case 1a
-Weak correlation
-No ZI
+Case 0a No correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
+1.1 0.0000 0.0000 0.0000 0.0000
+1.2 0.0136 0.0142 0.0047 0.0165
+1.3 0.0152 0.0148 0.0088 0.0161
+1.4 0.0119 0.0106 0.0049 0.0117
+1.5 0.0088 0.0090 0.0046 0.0120
+1.7 0.0097 0.0097 0.0044 0.0112
+1.9 0.0066 0.0068 0.0037 0.0072
+2 0.0073 0.0074 0.0043 0.0090
+2.5 0.0060 0.0060 0.0032 0.0062
+3 0.0048 0.0050 0.0023 0.0056
+Case 0b No correlation 30% ZI 1.01 0.0050 0.0000 0.0050 0.0150
+1.1 0.0052 0.0052 0.0103 0.0155
+1.2 0.0105 0.0093 0.0044 0.0155
+1.3 0.0101 0.0105 0.0077 0.0146
+1.4 0.0102 0.0092 0.0056 0.0125
+1.5 0.0113 0.0102 0.0053 0.0142
+1.7 0.0079 0.0078 0.0043 0.0098
+1.9 0.0075 0.0075 0.0033 0.0088
+2 0.0070 0.0071 0.0032 0.0074
+2.5 0.0068 0.0067 0.0035 0.0076
+3 0.0056 0.0055 0.0025 0.0063
+Case 1a Weak correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0155
 1.1 0.0030 0.0030 0.0040 0.0080
@@
 1.9 0.0056 0.0057 0.0038 0.0077
-2.0 0.0055 0.0057 0.0030 0.0071
+2 0.0055 0.0057 0.0030 0.0071
 2.5 0.0047 0.0048 0.0029 0.0067
-3.0 0.0042 0.0042 0.0026 0.0056
-Case 1b
-Weak correlation
-30% ZI
+3 0.0042 0.0042 0.0026 0.0056
+Case 1b Weak correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0100
 1.1 0.0010 0.0000 0.0010 0.0053
@@
 1.9 0.0056 0.0057 0.0037 0.0080
-2.0 0.0055 0.0055 0.0037 0.0078
+2 0.0055 0.0055 0.0037 0.0078
 2.5 0.0046 0.0045 0.0028 0.0065
-3.0 0.0037 0.0037 0.0022 0.0049
-Case 2a
-Moderate correlation
-No ZI
+3 0.0037 0.0037 0.0022 0.0049
+Case 2a Moderate correlation No ZI 1.01 0.0050 0.0050 0.0000 0.0000
 1.1 0.0010 0.0000 0.0000 0.0026
@@
 1.9 0.0041 0.0041 0.0036 0.0079
-2.0 0.0032 0.0032 0.0031 0.0075
+2 0.0032 0.0032 0.0031 0.0075
 2.5 0.0030 0.0030 0.0028 0.0063
-3.0 0.0024 0.0023 0.0026 0.0056
-Case 2b
-Moderate correlation
-30% ZI
+3 0.0024 0.0023 0.0026 0.0056
+Case 2b Moderate correlation 30% ZI 1.01 0.0100 0.0050 0.0100 0.0100
 1.1 0.0033 0.0005 0.0005 0.0082
@@
 1.9 0.0033 0.0032 0.0031 0.0073
-2.0 0.0027 0.0029 0.0032 0.0073
+2 0.0027 0.0029 0.0032 0.0073
 2.5 0.0025 0.0026 0.0030 0.0065
-3.0 0.0025 0.0025 0.0029 0.0062
-Case 3a
-Strong correlation
-No ZI
+3 0.0025 0.0025 0.0029 0.0062
+Case 3a Strong correlation No ZI 1.01 0.0051 0.0051 0.0000 0.0152
 1.1 0.0027 0.0017 0.0013 0.0032
@@
 1.9 0.0040 0.0038 0.0032 0.0081
-2.0 0.0030 0.0029 0.0029 0.0073
+2 0.0030 0.0029 0.0029 0.0073
 2.5 0.0022 0.0021 0.0030 0.0064
-3.0 0.0015 0.0017 0.0026 0.0058
-Case 3b
-Strong correlation
-30% ZI
+3 0.0015 0.0017 0.0026 0.0058
+Case 3b Strong correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0050
 1.1 0.0000 0.0000 0.0000 0.0074
@@
 1.9 0.0031 0.0032 0.0036 0.0082
-2.0 0.0031 0.0031 0.0033 0.0078
+2 0.0031 0.0031 0.0033 0.0078
 2.5 0.0024 0.0023 0.0026 0.0063
-3.0 0.0015 0.0015 0.0026 0.0059
+3 0.0015 0.0015 0.0026 0.0059
 Table S.4: Simulation of sensitivity for DP-based models
@@
 posterior quantile optim
-Case 1a
-Weak correlation
-No ZI
+Case 0a No correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
+1.1 0.0005 0.0004 0.0001 0.0003
+1.2 0.0269 0.0262 0.0202 0.0248
+1.3 0.0734 0.0727 0.0582 0.0656
+1.4 0.1199 0.1201 0.0985 0.1082
+1.5 0.1605 0.1606 0.1366 0.1483
+1.7 0.2269 0.2267 0.2012 0.2163
+1.9 0.2848 0.2850 0.2596 0.2748
+2 0.3098 0.3104 0.2849 0.3018
+2.5 0.4169 0.4177 0.3939 0.4120
+3 0.4921 0.4923 0.4711 0.4902
+Case 0b No correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
+1.1 0.0004 0.0004 0.0001 0.0003
+1.2 0.0269 0.0273 0.0212 0.0252
+1.3 0.0731 0.0732 0.0579 0.0662
+1.4 0.1208 0.1203 0.0985 0.1099
+1.5 0.1601 0.1601 0.1369 0.1492
+1.7 0.2302 0.2297 0.2048 0.2192
+1.9 0.2865 0.2868 0.2611 0.2769
+2 0.3117 0.3115 0.2873 0.3038
+2.5 0.4171 0.4164 0.3930 0.4115
+3 0.4902 0.4903 0.4695 0.4875
+Case 1a Weak correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0005 0.0005 0.0001 0.0003
@@
 1.9 0.3404 0.3407 0.2738 0.2895
-2.0 0.3647 0.3648 0.3002 0.3164
+2 0.3647 0.3648 0.3002 0.3164
 2.5 0.4584 0.4588 0.4061 0.4231
-3.0 0.5245 0.5252 0.4818 0.4988
-Case 1b
-Weak correlation
-30% ZI
+3 0.5245 0.5252 0.4818 0.4988
+Case 1b Weak correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0006 0.0004 0.0001 0.0003
@@
 1.9 0.3399 0.3396 0.2728 0.2883
-2.0 0.3639 0.3639 0.2999 0.3152
+2 0.3639 0.3639 0.2999 0.3152
 2.5 0.4571 0.4575 0.4048 0.4211
-3.0 0.5256 0.5261 0.4813 0.4998
-Case 2a
-Moderate correlation
-No ZI
+3 0.5256 0.5261 0.4813 0.4998
+Case 2a Moderate correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0027 0.0029 0.0003 0.0006
@@
 1.9 0.5415 0.5417 0.3638 0.3810
-2.0 0.5572 0.5571 0.3905 0.4078
+2 0.5572 0.5571 0.3905 0.4078
 2.5 0.6167 0.6172 0.4976 0.5147
-3.0 0.6597 0.6599 0.5667 0.5826
-Case 2b
-Moderate correlation
-30% ZI
+3 0.6597 0.6599 0.5667 0.5826
+Case 2b Moderate correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0028 0.0025 0.0004 0.0007
@@
 1.9 0.5408 0.5408 0.3632 0.3801
-2.0 0.5571 0.5572 0.3914 0.4085
+2 0.5571 0.5572 0.3914 0.4085
 2.5 0.6177 0.6180 0.4974 0.5141
-3.0 0.6611 0.6613 0.5674 0.5835
-Case 3a
-Strong correlation
-No ZI
+3 0.6611 0.6613 0.5674 0.5835
+Case 3a Strong correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0053 0.0055 0.0004 0.0009
@@
 1.9 0.7177 0.7176 0.4124 0.4316
-2.0 0.7290 0.7290 0.4449 0.4636
+2 0.7290 0.7290 0.4449 0.4636
 2.5 0.7675 0.7676 0.5590 0.5768
-3.0 0.7959 0.7961 0.6307 0.6470
-Case 3b
-Strong correlation
-30% ZI
+3 0.7959 0.7961 0.6307 0.6470
+Case 3b Strong correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0039 0.0044 0.0004 0.0007
@@
 1.9 0.7149 0.7150 0.4131 0.4330
-2.0 0.7275 0.7276 0.4441 0.4636
+2 0.7275 0.7276 0.4441 0.4636
 2.5 0.7669 0.7670 0.5584 0.5761
-3.0 0.7960 0.7962 0.6299 0.6462
+3 0.7960 0.7962 0.6299 0.6462
 Table S.5: Simulation of average type I error for DP-based models
@@
 posterior quantile optim
-Case 1a
-Weak correlation
-No ZI
+Case 0a No correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0000 0.0000 0.0000 0.0000
@@
 1.9 0.0000 0.0000 0.0000 0.0000
+2 0.0000 0.0000 0.0000 0.0001
 2.5 0.0000 0.0000 0.0000 0.0000
-2.5 0.0000 0.0000 0.0000 0.0000
-3.0 0.0000 0.0000 0.0000 0.0000
-Case 1b
-Weak correlation
-30% ZI
+3 0.0000 0.0000 0.0000 0.0001
+Case 0b No correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0000 0.0000 0.0000 0.0000
@@
 1.9 0.0000 0.0000 0.0000 0.0000
-2.0 0.0000 0.0000 0.0000 0.0000
-2.5 0.0000 0.0000 0.0000 0.0000
-3.0 0.0000 0.0000 0.0000 0.0000
-Case 2a
-Moderate correlation
-No ZI
+2 0.0000 0.0000 0.0000 0.0000
+2.5 0.0001 0.0001 0.0000 0.0001
+3 0.0001 0.0001 0.0000 0.0001
+Case 1a Weak correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0000 0.0000 0.0000 0.0000
@@
 1.9 0.0000 0.0000 0.0000 0.0000
+2 0.0000 0.0000 0.0000 0.0000
 2.5 0.0000 0.0000 0.0000 0.0000
-2.5 0.0000 0.0000 0.0000 0.0000
-3.0 0.0000 0.0000 0.0000 0.0000
-Case 2b
-Moderate correlation
-30% ZI
+3 0.0000 0.0000 0.0000 0.0000
+Case 1b Weak correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0000 0.0000 0.0000 0.0000
@@
 1.9 0.0000 0.0000 0.0000 0.0000
+2 0.0000 0.0000 0.0000 0.0000
+2.5 0.0000 0.0000 0.0000 0.0000
+3 0.0000 0.0000 0.0000 0.0000
+Case 2a Moderate correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0000 0.0000 0.0000 0.0000
@@
 1.3 0.0000 0.0000 0.0000 0.0000
-Case 3a
-Strong correlation
-No ZI
+1.4 0.0000 0.0000 0.0000 0.0000
+1.5 0.0000 0.0000 0.0000 0.0000
+1.7 0.0000 0.0000 0.0000 0.0000
+1.9 0.0000 0.0000 0.0000 0.0000
+2 0.0000 0.0000 0.0000 0.0000
+2.5 0.0000 0.0000 0.0000 0.0000
+3 0.0000 0.0000 0.0000 0.0000
+Case 2b Moderate correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
+1.1 0.0000 0.0000 0.0000 0.0000
+1.2 0.0000 0.0000 0.0000 0.0000
+1.3 0.0000 0.0000 0.0000 0.0000
+1.4 0.0000 0.0000 0.0000 0.0000
+1.5 0.0000 0.0000 0.0000 0.0000
+1.7 0.0000 0.0000 0.0000 0.0000
+1.9 0.0000 0.0000 0.0000 0.0000
+2 0.0000 0.0000 0.0000 0.0000
+2.5 0.0000 0.0000 0.0000 0.0000
+3 0.0000 0.0000 0.0000 0.0000
+Case 3a Strong correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0000 0.0000 0.0000 0.0000
@@
 1.9 0.0000 0.0000 0.0000 0.0000
+2 0.0000 0.0000 0.0000 0.0000
 2.5 0.0000 0.0000 0.0000 0.0000
-2.5 0.0000 0.0000 0.0000 0.0000
-3.0 0.0000 0.0000 0.0000 0.0000
-Case 3b
-Strong correlation
-30% ZI
+3 0.0000 0.0000 0.0000 0.0000
+Case 3b Strong correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0000 0.0000 0.0000 0.0000
@@
 1.9 0.0000 0.0000 0.0000 0.0000
+2 0.0000 0.0000 0.0000 0.0000
 2.5 0.0000 0.0000 0.0000 0.0000
-2.5 0.0000 0.0000 0.0000 0.0000
-3.0 0.0000 0.0000 0.0000 0.0000
+3 0.0000 0.0000 0.0000 0.0000
 Table S.6: Simulation of F-score for DP-based models
@@
 posterior quantile optim
-Case 1a
-Weak correlation
-No ZI
+Case 0a No correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
+1.1 0.0010 0.0007 0.0002 0.0007
+1.2 0.0520 0.0506 0.0394 0.0480
+1.3 0.1359 0.1348 0.1094 0.1224
+1.4 0.2132 0.2135 0.1787 0.1944
+1.5 0.2755 0.2756 0.2395 0.2571
+1.7 0.3686 0.3683 0.3340 0.3542
+1.9 0.4419 0.4421 0.4111 0.4296
+2 0.4715 0.4722 0.4423 0.4620
+2.5 0.5868 0.5876 0.5640 0.5819
+3 0.6580 0.6581 0.6394 0.6561
+Case 0b No correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
+1.1 0.0009 0.0008 0.0002 0.0007
+1.2 0.0520 0.0527 0.0412 0.0488
+1.3 0.1355 0.1356 0.1089 0.1235
+1.4 0.2147 0.2139 0.1786 0.1971
+1.5 0.2749 0.2749 0.2401 0.2584
+1.7 0.3729 0.3722 0.3390 0.3582
+1.9 0.4439 0.4443 0.4131 0.4321
+2 0.4738 0.4735 0.4453 0.4645
+2.5 0.5868 0.5861 0.5631 0.5811
+3 0.6561 0.6562 0.6380 0.6535
+Case 1a Weak correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0009 0.0009 0.0003 0.0006
@@
 1.9 0.5062 0.5065 0.4286 0.4472
-2.0 0.5328 0.5329 0.4605 0.4788
+2 0.5328 0.5329 0.4605 0.4788
 2.5 0.6269 0.6272 0.5762 0.5925
-3.0 0.6864 0.6869 0.6489 0.6635
-Case 1b
-Weak correlation
-30% ZI
+3 0.6864 0.6869 0.6489 0.6635
+Case 1b Weak correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0011 0.0009 0.0003 0.0007
@@
 1.9 0.5056 0.5053 0.4273 0.4456
-2.0 0.5319 0.5319 0.4600 0.4773
+2 0.5319 0.5319 0.4600 0.4773
 2.5 0.6257 0.6261 0.5749 0.5907
-3.0 0.6875 0.6879 0.6486 0.6646
-Case 2a
-Moderate correlation
-No ZI
+3 0.6875 0.6879 0.6486 0.6646
+Case 2a Moderate correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0049 0.0052 0.0006 0.0011
@@
 1.9 0.7010 0.7012 0.5324 0.5498
-2.0 0.7144 0.7143 0.5606 0.5774
+2 0.7144 0.7143 0.5606 0.5774
 2.5 0.7616 0.7620 0.6633 0.6775
-3.0 0.7938 0.7939 0.7223 0.7342
-Case 2b
-Moderate correlation
-30% ZI
+3 0.7938 0.7939 0.7223 0.7342
+Case 2b Moderate correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0052 0.0047 0.0008 0.0014
@@
 1.9 0.7006 0.7007 0.5317 0.5490
-2.0 0.7144 0.7144 0.5614 0.5781
+2 0.7144 0.7144 0.5614 0.5781
 2.5 0.7626 0.7627 0.6631 0.6770
-3.0 0.7948 0.7949 0.7227 0.7347
-Case 3a
-Strong correlation
-No ZI
+3 0.7948 0.7949 0.7227 0.7347
+Case 3a Strong correlation No ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0094 0.0097 0.0009 0.0017
@@
 1.9 0.8339 0.8339 0.5830 0.6010
-2.0 0.8420 0.8420 0.6149 0.6316
+2 0.8420 0.8420 0.6149 0.6316
 2.5 0.8674 0.8675 0.7160 0.7295
-3.0 0.8856 0.8856 0.7725 0.7836
-Case 3b
-Strong correlation
-30% ZI
+3 0.8856 0.8856 0.7725 0.7836
+Case 3b Strong correlation 30% ZI 1.01 0.0000 0.0000 0.0000 0.0000
 1.1 0.0068 0.0077 0.0008 0.0014
@@
 1.9 0.8324 0.8324 0.5836 0.6023
-2.0 0.8409 0.8410 0.6140 0.6315
+2 0.8409 0.8410 0.6140 0.6315
 2.5 0.8670 0.8671 0.7156 0.7290
-3.0 0.8857 0.8858 0.7718 0.7829
+3 0.8857 0.8858 0.7718 0.7829
+S.3.2 Simulation results for sensitivity analysis for grid configuration of $k$ in hypothesis testing
+In our signal detection procedure, a sequence of threshold values for $k$ is constructed on a predefined grid.
+By default, we consider a grid ranging from $k=1.1$ to $k=3$, with small increments.
+This range is primarily motivated by the simulation design, where the maximum signal strength is set to be no larger than $3$, rendering this grid sufficient to capture all relevant signal levels under the data-generating mechanism.
+To assess whether the default range for $k$ is restrictive, we conducted an additional sensitivity analysis under simulation Cases 2a.
+In this analysis, we adopted a highly conservative strategy by extending the upper bound of the grid to the maximum value observed among all MCMC draws of $\lambda$, effectively performing a global search over a substantially wider range of $k$.
+The resulting performance curves under this expanded grid are indistinguishable from those obtained using the default grid, across all evaluation metrics.
+These results in Figure S.3.2 indicate that, for the considered simulation setting under Case 2a, the default grid range $(1.1,3)$ is adequate and does not lead to any loss in detection performance.
+More importantly, the consistency between the two strategies suggests that the proposed method is robust to the choice of the search range for $k$.
+Consequently, in applications involving datasets with potentially stronger signals, adopting a more conservative, data-driven upper bound for $k$ remains a safe option and does not adversely affect performance.
+S.3.3 Simulation results for sensitivity analysis for hyperparameters
+We conducted a simulation study under Case 2a to evaluate the robustness of the proposed DP Poisson local-global model with respect to the choice of hyperparameters.
+The model structure was kept fixed throughout the simulations, and sensitivity analysis was performed by varying the hyperparameters governing the prior distributions of $\alpha$ and $\beta$, which control the strength and concentration of shrinkage in the local-global framework.
+Specifically, the baseline model adopts $\psi_{\alpha}=3$ and $\psi_{\beta}=0.5$ as our default model.
+To assess sensitivity, we considered alternative hyperprior specifications that represent substantially different prior beliefs while remaining within reasonable ranges.
+For $\alpha$, we examined a more informative prior with $\psi_{\alpha}=5$, placing approximately $99\%$ prior mass within $(0,1)$, as well as a weaker prior with $\psi_{\alpha}=1$, placing approximately $64\%$ mass in $(0,1)$.
+For $\beta$, we considered $\psi_{\beta}=1/7$, corresponding to a prior favoring values in $(0,1)$ with approximately $90\%$ prior mass in this interval, and $\psi_{\beta}=6$, which places only $10\%$ prior mass in $(0,1)$ and thus favors more extreme shrinkage behavior.
+These hyperprior configurations are summarized in Table S.7.
+The comparison across these settings allows us to evaluate the stability of model performance under varying degrees of prior informativeness and shrinkage strength.
+Table S.7: Hyperprior configurations used in the sensitivity analysis. The baseline setting corresponds to the default hyperprior choice in the proposed DP Poisson local–global model.
+Model setting $\psi_{\alpha}$ $\psi_{\beta}$ Interpretation
+Baseline 3 0.5 Default hyperprior specification in the proposed model
+$\alpha$-strong 5 0.5 More informative prior on $\alpha$, placing $\approx 99\%$ mass in $(0,1)$
+$\alpha$-weak 1 0.5 Weaker prior on $\alpha$, placing $\approx 64\%$ mass in $(0,1)$
+$\beta$-favoring 3 $1/7$ Prior favoring $\beta\in(0,1)$ with $\approx 90\%$ mass in $(0,1)$
+$\beta$-sparse 3 6 Prior discouraging $\beta\in(0,1)$, placing only $10\%$ mass in $(0,1)$
+Figure S.1: Simulation results under different hyperprior configurations for the DP Poisson local-global model.
+Curves corresponding to all hyperprior settings largely overlap across performance metrics, indicating robust model performance with respect to reasonable variations in hyperparameter choices.
+As shown in Figure S.1, across all performance metrics, the curves corresponding to different hyperprior configurations largely overlap in the simulation results.
+This indicates that the proposed DP Poisson local-global model exhibits stable performance under a wide range of hyperparameter choices.
+In particular, varying the hyperpriors for $\alpha$ and $\beta$, from relatively informative to weakly informative specifications, does not materially affect model behavior at the considered data scale.
+These findings suggest that the proposed model is reasonably robust to reasonable variations in hyperprior assumptions.
+Moreover, the sensitivity analysis covers several representative hyperprior scenarios, which together are expected to reflect a broad range of practical data-generating settings.
+S.3.4 Simulation results under smaller setting
+We conduct the simulation based on Table S.2 under similar correlation strength setting in Table 1.
+The result shown in Figure S.2 has the same pattern as we observed in Section 5 based on larger setting.
+Figure S.2: Simulation results: the horizontal axis shows the true signal strength $\lambda_{0}$ for signal cells, the vertical axis shows the value of evaluation metrics.
+Each column presents one metric and each row presents one simulation scenario.